### v2

# Any-Time Regret-Guaranteed Algorithm for Control of Linear Quadratic Systems

Abstract
In this paper, we propose a computationally efficient algorithm that achieves anytime regret of order $\mathcal{O}(\sqrt{t})$, with explicit dependence on the system dimensions and on the solution of the Discrete Algebraic Riccati Equation (DARE). Our approach builds on the SDP-based framework of cohen2019learning, using an appropriately tuned regularization and a sufficiently accurate initial estimate to construct confidence ellipsoids for control design. A carefully designed input-perturbation mechanism is incorporated to ensure anytime performance. We develop two variants of the algorithm. The first enforces strong sequential stability, requiring each policy to be stabilizing and successive policies to remain close. This sequential condition helps prevent state explosion at policy update times and enables high-probability guarantees on the state trajectory; however, it results in a conservative design and suboptimal regret scaling with respect to the DARE solution. Motivated by this limitation, we introduce a second class of algorithms that removes the sequential-stability requirement and instead requires only that each generated policy be stabilizing. Closed-loop stability is then preserved through a dwell-time–inspired policy-update rule, adapting ideas from switched-systems control to carefully balance exploration and exploitation. This class of algorithms also addresses key shortcomings of most existing approaches—including certainty-equivalence–based methods—which typically guarantee stability only in the Lyapunov sense and lack explicit uniform high-probability bounds on the state trajectory expressed in system-theoretic terms. In particular, such methods fail to prevent transient state spikes caused by transitions between stabilizing policies, a critical issue in safety-sensitive systems. Our analysis explicitly characterizes the trade-off between state amplification and regret, and shows that partially relaxing the sequential-stability requirement yields optimal regret. Finally, our method eliminates the need for any a priori bound on the norm of the DARE solution—an assumption required by all existing computationally efficient optimism in the face of uncertainty (OFU) based algorithms—and thereby removes the reliance of regret guarantees on such external inputs.
Keywords: Reinforcement Learning, Controls, LQR, Regret-Bound
1 Introduction
The problem of designing learning-based control algorithms for Linear Quadratic Regulator (LQR) systems with unknown dynamics—particularly under non-asymptotic performance guarantees—has received significant attention over the past decade. The LQR framework is a particularly important paradigm in the controls community, primarily due to its closed-form solution. In LQR setting the state of the system evolves according the linear dynamics
$x_{t+1}$ $=A_{*}x_{t}+B_{*}u_{t}+\omega_{t+1},$ (1)
where matrices $A_{*}\in\mathbb{R}^{n\times n}$ and $B_{*}\in\mathbb{R}^{n\times m}$ and the process noise $\omega_{t+1}$ usually is assumed to ben drawn i.i.d from a zero-mean gaussian distribution.
The incurred quadratic cost $c_{t}$, a weighted sum of the state $x_{t}$ and control input $u_{t}$ norms, is defined as
$c_{t}=x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t},$ (2)
where the cost matrices satisfy $Q\succeq 0$ and $R\succ 0$.
With full knowledge of system parameters $A_{*}$, $B_{*}$, and assuming controllability of $(A_{*},B_{*})$, it is known that the optimal sequence of control inputs $u^{*}=\{u^{*}_{t}\}_{t=1}^{T-1}$ minimizing average expected cost
$J_{\mathcal{A}}=$ $\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}[\sum_{t=0}^{T-1}c_{t}]$ (3)
subject to dynamics (1) has a close-form solution. This sequence indeed is in the form of static state feedback $u^{*}_{t}=K_{*}x_{t}$ where $K_{*}$ is computed using the solution of the Discrete Algebraic Riccati Equation (DARE), $P_{*}$. The optimal average expected cost is $J_{*}=P_{*}\bullet W$, where $W$ is the covariance matrix of the Gaussian noise from which the process noise $\omega_{t}$ is drawn.
In the absence of knowledge of the parameters $A_{*}$ and $B_{*}$, the core challenge is to design an algorithm that stabilizes the system while minimizing the accumulated quadratic cost, quantified by a finite-time performance measure. In this study, we adopt regret as the performance measure, defined as follows. For an arbitrary algorithm $\mathcal{A}$, which generates the input sequence $u=\{u_{t}\}_{t=0}^{T-1}$, regret is given by
$R_{\mathcal{A}}(T)=\sum_{t=0}^{T-1}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}-J_{*}\right),$ (4)
which captures the difference between the accumulated cost incurred by the algorithm and the best achievable cost, i.e., the optimal cost.
The overarching goal of this paper, as in the related literature, is to design an algorithm that minimizes the regret (4) subject to the system dynamics (1), when the parameters $A_{*}$ and $B_{*}$ are unknown.
1.1 Contributions and informal statements of results
The first algorithm proposed for the control of LQR with a worst-case regret bound of $\mathcal{O}(\sqrt{T})$ is the Optimism in the Face of Uncertainty (OFU)-based algorithm introduced by abbasi2011regret. This algorithm performs system identification by constructing confidence ellipsoids and designs the control policy by playing optimistically with respect to this set. Although the algorithm enjoys strong regret guarantees, its nonconvexity and the resulting computational burden has spurred subsequent research into more efficient formulations. with guaranteed $\mathcal{O}(\sqrt{T})$.
Despite success in addressing the computational complexity gap, most existing approaches in the literature are horizon-dependent, as they are designed under the assumption that the time horizon $T$ is known a priori for an episode. These methods typically provide a regret guarantee of order $\mathcal{O}(\sqrt{T})$ over the specified horizon, usually evaluated at the end of the episode. To extend such performance to longer horizons, the commonly adopted remedy is the doubling trick. However, even with this technique, there remains a reliance on prior knowledge of the epochs. This is particularly problematic in control applications, where the horizon is often unknown or subject to change, making horizon-dependent algorithms less suitable for real-world deployment. Therefore, algorithms with anytime guarantees, which do not require such prior knowledge, are generally more practical. In this work, we propose a new anytime regret-minimizing algorithm that does not require knowledge of the horizon in advance. Our approach builds on the OFU-based method introduced by cohen2019learning.
Another property that may be important for making learning-based algorithms transferable across control domains is the provision of system-theoretic guarantees on the state-norm trajectory. Establishing such bounds, however, requires a thorough analysis of the closed-loop stability induced by the proposed algorithms. Generating individually stabilizing policies is not sufficient to ensure closed-loop stability, as instability (or state explosion) may still occur due to policy updates, a well known fact in literature on switched-systems control. Previous works have attempted to address this challenge through various approaches. For instance, certainty-equivalence (CE) methods simchowitz2020naive; jedra2022minimal employ a common Lyapunov function to guarantee Lyapunov stability of the closed-loop system. However, it is important to note that Lyapunov decay alone does not necessarily provide explicit high-probability, uniform control of the state norm in the presence of process noise. Under zero-mean Gaussian disturbances, Lyapunov decay guarantees only that the expected state norm, $\mathbb{E}\left[\|x_{t}\|^{2}\right]$, remains bounded, and deriving a high-probability bound requires additional effort. Although a high-probability state-norm bound is provided by simchowitz2020naive (see their Lemma 5.3 and its proof), this bound is conservative and does not accurately capture the true scale of possible state-norm excursions. Therefore, while Lyapunov stability is sufficient for controlling accumulated regret—since a finite number of such spikes can typically be absorbed in regret bounds—it is insufficient for obtaining high-probability guarantees on the state norm, which are essential for safety-critical systems. Another work that provides a high-probability state-norm bound is abeille2020efficient, but the bound is not fully system-theoretic and depends on a priori bounds, which may be loose. In another work, cohen2019learning introduced the notion of strong sequential stability, which enables an explicit high-probability guarantee on the state norm. This notion imposes a stricter requirement than standard stability: the algorithm must not only output stabilizing policies, but also ensure that any two consecutively generated policies remain close in a suitable sense, preventing the state norm from blowing up due to policy updates. While this approach provides a high-probability bound on the state norm, in this work we show that the strong sequential-stability requirement is conservative from the perspective of regret guarantees, leading to an $\mathcal{O}(|P_{*}|^{8})$ dependence on the operator norm of the DARE solution. This issue motivates us to propose a method that incorporates a mechanism inspired by dwell-time design in switched-systems control. Using this mechanism, we carefully regulate the exploration–exploitation trade-off to improve regret bounds while explicitly restricting state amplification and ensuring uniform control of the closed-loop trajectories. A natural first idea would be to require the algorithm to generate only stabilizing policies and then rely on a dwell-time condition to control both the state norm and the regret. However, sequential stability is not a binary property; rather, it forms a tunable continuum. Our analysis shows that fully enforcing sequential stability leads to suboptimal regret, while discarding it entirely and relying solely on a dwell-time–type design for exploration–exploitation tuning also degrades regret. The optimal performance arises at an intermediate relaxation level, where stability and exploration are balanced. We identify a specific tuning that guarantees an $\mathcal{O}(\|P_{*}\|^{6})$ dependence in the regret bound, while maintaining control of the state norm with a dependence of order $\mathcal{O}(\|P_{*}\|^{2.5})$. Our analysis reveals a clear trade-off between the state-norm bound and the regret bound in terms of their dependence on $\|P_{*}\|$, and it provides the flexibility to tune this trade-off as needed.
Another challenge in designing regret-guaranteed algorithms for LQ control is the reliance on a priori information as an input. In this paper, we eliminate the need for prior access to the optimal average cost $J_{*}$, which is commonly required in existing computationally efficient OFU-based algorithms. While this issue has been resolved for certainty-equivalence methods, as studied in simchowitz2020naive; jedra2022minimal, it has remained unaddressed in current computationally efficient OFU-based algorithms, including abeille2020efficient; cohen2019learning. Such a requirement makes the performance of these algorithms dependent on an external bound that may be loose, which is reflected in their regret upper bounds. We address this limitation by introducing a warm-up phase that runs a dual SDP problem in the loop to detect when the parameter estimate enters the desired ball, whose radius depends on $\|P_{*}\|$.
1.1.1 Contributions.
Our core contributions are briefly threefold: (i) We introduce new computationally efficient anytime-regret algorithms for LQ control, motivated by the need for methods that remain effective when the implementation horizon is not known a priori. (ii) We emphasize the importance of providing a high-probability, system-theoretic bound on the state norm. We offer a new theoretical understanding of strong sequential stability, highlight its limitations, and demonstrate how relaxing this notion can improve algorithm performance. A key component of this development is a novel mechanism, inspired by dwell-time design in switched-systems control, which enables principled tuning of the exploration–exploitation trade-off while ensuring both favorable regret bounds and closed-loop stability. (iii) We eliminate the need for a priori knowledge of the optimal average cost $J_{*}$, an assumption required by existing computationally efficient OFU-based algorithms.
1.1.2 Informal statements of our main results.
$(1)$ Any-time regret algorithm with guaranteed system theoretic regret bound We propose an any-time SDP-based LQ cOntrol (ARSLO) built upon the SDP-based algorithm proposed by cohen2019learning, OSLO which is a horizon dependant algorithm. The OSLO algorithm when running for $T$ time steps requires $T$-dependant regularization parameters used for least square estimation and confidence bound construction as well as an initial parameter estimate in $\mathcal{O}(1/T^{1/4})$ neighborhood of true but unknown system dynamic parameters. The latter requires running a warm-up algorithm for $\mathcal{O}(\sqrt{T})$ time steps. To extend to a longer horizon using the doubling trick, such a warm-up phase must be implemented for the extended version in each epoch. Demonstrating this issue, we propose an any-time regret guaranteed algorithm by choosing an appropriate regularization parameter for the least square estimation and applying additive input perturbation $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, where $\|\Gamma_{t}\|_{op}=\mathcal{O}(1/\sqrt{t})$. Therefore, our ARSLO algorithm, provides anytime regret guarantees while also ensuring the strong sequential stability property of the generated policies (see Definition 1). Specifically, ARSLO achieves the regret bound
${R_{{ARSLO}}(t)\leq\mathcal{O}\big(\sqrt{{n^{2}(n+m)^{3}}{\|P_{*}\|_{op}^{16}}\;{t}\log^{4}\frac{t}{\delta}}\big)}.$
$(2)$ Doing away with strong sequential stability notion and improving regret bound
[…]
${R_{{ARSLO^{+}(\bar{\rho})}}(t)\leq\mathcal{O}\big(\sqrt{{n^{2}(n+m)^{3}}{\|P_{*}\|_{op}^{12}}\;{t}\log^{4}\frac{t}{\delta}}\big)}.$
for a specific value of $\bar{\rho}=2$.
Related relaxations of the sequential-stability requirement have been explored in prior work. For example, CE–based methods simchowitz2020naive; jedra2022minimal rely on a common Lyapunov function approach. However, these methods do not provide precise, explicit, system-theoretic high-probability bounds on the state-norm trajectory, as they guarantee stability only in the Lyapunov sense. Consequently, while such approaches optimize overall regret and achieve favorable regret guarantees, they offer limited control over transient state spikes at policy update times, which may be problematic in safety-critical settings. For OFU-based algorithms, the high-probability state-norm bound derived in abeille2020efficient is also not fully system-theoretic, as it depends on an a priori bound on the optimal average expected cost.
$(3)$ Contributing to the line of OFU-based algorithms for LQR control Compared to the only two existing convex and computationally efficient OFU-based algorithms by cohen2019learning and abeille2020efficient, our OFU-based algorithm provides anytime regret guarantees. To our knowledge, this makes it the first convex (i.e., computationally efficient) OFU-based algorithm with an anytime regret guarantee.
Leaving aside the anytime regret guarantee of our proposed algorithms, their system-theoretic regret upper bound challenges that of the Lag-LQR approach proposed by abeille2020efficient. Lag-LQR is developed by applying techniques from extended value iteration (EVI) to the Optimistic LQR with relaxed constraints, followed by a Lagrangian relaxation to design a computationally efficient algorithm. This algorithm comes with the following regret upper bound:
$R_{{LAG-LQ}}\leq\mathcal{O}({D^{\frac{3}{2}}}\sqrt{{n(n+m)^{2}}{\|P_{*}\|_{op}^{4}}{T}\log^{3}\frac{T}{\delta}})$
where $\operatorname{Tr}(P_{*})\leq{D}$ is an a priori known bound used by the algorithm. A similar issue also appears in cohen2019learning. In contrast to LAG-LQR and OSLO, our algorithm does not require such an a priori bound, and its regret upper bound is purely system-theoretic. We address this by jointly utilizing the solutions of the primal and dual SDP formulations to design the input perturbation noise and to tune the exploration–exploitation trade-off parameter. Furthermore, the dimension dependence of our regret bound is $\mathcal{O}\big(\sqrt{n^{2}(n+m)^{3}}\big)$, whereas for LAG-LQR it is $\mathcal{O}\big(\sqrt{n^{4}(n+m)^{2}}\big)$, when accounting for the hidden dependencies of $D$ on $n$ and $\|P_{*}\|_{\mathrm{op}}$.
Finally, leveraging both the primal and dual solutions allows us to design input perturbations that enable non-isotropic exploration. The covariance matrix of the Gaussian input perturbation is constructed from the relaxed primal SDP solution itself (not its norm), augmented with a bias term that depends on the norm of the dual solution, thereby facilitating non-isotropic exploration.
1.2 Related works
In recent years, the machine learning community has shown increasing interest in addressing control design in the context of LQR, under various scenarios of what is known and what is unknown. Notably, agarwal2019online studied the setting with (nonstochastic) adversarial perturbations and established regret guarantees. cohen2018online investigated the control of linear systems with a known model but adversarially chosen quadratic costs, also providing regret guarantees. When the system model matrices are unknown, which is the main focus of this paper, existing efforts can be categorized into model-free and model-based settings. Model-free approaches design controllers without explicit system identification; instead, they minimize the cost function directly, relying solely on available data (see dean2020sample; fazel2018global; abbasi2019model). On the other hand, model-based approaches perform control design through explicit model identification or confidence set construction and typically provide regret guarantees. Below, we briefly review several relevant papers in the literature, presented in chronological order.
The first algorithm proposed for the control of LQR with a worst-case regret bound of $\mathcal{O}(\sqrt{T})$ is the OFU-based algorithm introduced by abbasi2011regret. This algorithm performs system identification by constructing confidence ellipsoids and designs the control policy by playing optimistically with respect to this set. Despite its success, the algorithm’s nonconvexity became a key motivation for subsequent work in the literature to develop fully convex algorithms with guaranteed $\mathcal{O}(\sqrt{T})$ worst-case regret.
In the OFU line of research, faradonbeh2020optimism proposed an algorithm for general stabilizable systems, while also accounting for a fairly general heavy-tailed noise distribution. Despite this extension, the proposed algorithm remained computationally burdensome. dean2018regret introduced a computationally efficient robust adaptive algorithm with polynomial time complexity that attains only regret of $\mathcal{O}(T^{2/3})$ throughout the learning process. However, the question of whether a computationally efficient (convex) algorithm can achieve an $\mathcal{O}(\sqrt{T})$ regret remained unanswered until the works proposed by cohen2019learning, faradonbeh2020input, and mania2019certainty. faradonbeh2020input introduced an $\epsilon$-greedy-like algorithm that is computationally efficient algorithm and achieves $\mathcal{O}(\sqrt{T})$ regret, but without explicit bounds for parameter estimate errors. Similarly mania2019certainty applied certainity equivalence principle and $\epsilon$-greedy exploration noise and proposed a computationally efficient $\mathcal{O}({\sqrt{T}})$ regret. simchowitz2020naive proposed a certainty equivalence-based algorithm and derived a system-theoretic regret upper and lower bound with $\mathcal{O}(\sqrt{T})$. abeille2020efficient advanced the original OFU-based algorithm by applying ideas from extended value iteration in optimistic algorithms and proposing an $\epsilon$-optimistic control, which is computationally efficient while still guaranteeing $\mathcal{O}(\sqrt{T})$ regret. The problem is also addressed by the Thomson sampling approach. In this line of work, abeille2018improved introduced the first algorithm tailored for scalar systems, achieving a regret of $\mathcal{O}(\sqrt{T})$. Subsequently, kargin2022thompson addressed the issue for multidimensional systems.
A commonly used categorization of regret-guaranteed algorithms for LQ control—also adopted in jedra2022minimal—classifies existing works into (i) self-tuning regulator–based approaches and (ii) OFU–based approaches. Self-tuning regulator algorithms estimate the unknown system matrices $(A_{*},B_{*})$, treat these estimates as the true parameters, solve the corresponding DARE, and design the control input accordingly. To ensure sufficient excitation of the system—and hence enable consistent estimation of $(A_{*},B_{*})$—the control inputs are typically perturbed with carefully tuned noise. OFU-based approaches, on the other hand, construct a confidence set for the unknown system parameters and design the control policy by acting optimistically with respect to this set. Our algorithm, while OFU-based, also employs input perturbations for exploration. As a result, our approach can be viewed as lying at the intersection of these two categories. A more detailed discussion and quantitative comparisons are provided in Appendix A.
[…]
2 Problem Statement and Preliminaries
Notation. For matrix $M$, $\|M\|_{*}=\operatorname{tr}(\sqrt{M^{\top}M})$ is the trace norm, $\|M\|_{F}=\sqrt{\operatorname{tr}(M^{\top}M)}$ is the Frobenius norm and $\|M\|$ is the spectral norm. $\lambda_{\min}(.)$ represents the minimum eigenvalue. Finally, we use $A\bullet B$ to denote the entry-wise dot product between matrices, namely $A\bullet B=\operatorname{Tr}(A^{\top}B)$. We write $f\lesssim g$ to denote that $f(x)\leq C\,g(x)$ for some universal constant $C>0$. The notation $\mathcal{O}$ indicates dependencies only on the time $t$, the confidence parameter $\delta$, $\|P_{*}\|$, and the dimensions $n$ and $m$.
2.1 Problem Statement
Consider a linear time-invariant system (1), equivalently rewritten as
$x_{t+1}=\Theta_{*}^{\top}z_{t}+\omega_{t+1},\quad z_{t}=\begin{pmatrix}x_{t}\\ u_{t}\end{pmatrix}$ (5)
where ${\Theta_{*}}=(A_{*},B_{*})^{\top}\in\mathbb{R}^{(n+m)\times n}$ initially unknown. We assume that the process noise $\omega_{t+1}$ fulfills the following assumption:
Assumption 1
There exists a filtration $\mathcal{F}_{t}$ such that
$(1.1)$ $\omega_{t+1}$ is a martingale difference, i.e., $\mathbb{E}[\omega_{t+1}|\;\mathcal{F}_{t}]=0$
$(1.2)$ $\mathbb{E}[\omega_{t+1}\omega_{t+1}^{\top}|\;\mathcal{F}_{t}]={\sigma}_{\omega}^{2}I_{n}=:W$ for some ${\sigma}_{\omega}^{2}>0$;
Our goal is to design an algorithm that minimizes regret (4) subject to dynamics (5), i.e.,
$\textrm{minimize}_{\{u_{t}\}_{t=0}^{T-1}}\;\;\;$ $R_{\mathcal{A}}(T)=\sum_{t=0}^{T-1}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}-J_{*}\right)$
$\textrm{S.t.}\;\;\;x_{t+1}=\Theta_{*}^{\top}z_{t}+\omega_{t+1}.$
We need the following assumptions for algorithm design purposes.
Assumption 2
$(1)$ There are known constants $\alpha_{0},\alpha_{1},\vartheta$ such that,
$\alpha_{0}I\preceq Q,R\preceq\alpha_{1}I,\quad\|\Theta_{*}\|\leq\vartheta.$
$(2)$ There is an initial stabilizing policy $K_{0}$ to start the strategy with.
It is worth mentioning that, since $Q$ and $R$ are known, determining $\alpha_{0}$ and $\alpha_{1}$ is straightforward. Furthermore, unlike cohen2018online, which requires an a priori bound $J_{*}\leq\nu$, our proposed strategy eliminates the need for such an assumption. This is achieved by carefully leveraging both the primal and dual solutions in the design of input perturbation noise and a smart warm-up phase that continuously runs the dual SDP problem in the loop. It is also important to highlight that a similar a priori bound is required in the other OFU-based algorithm proposed in abeille2020efficient, where it explicitly appears in the regret upper bound. That being said, our algorithm is the first computationally efficient OFU-based method that does not require such prior knowledge.
We review the notion of strong sequential stability introduced in cohen2019learning, which will be used in our analysis.
Definition 1
Consider the linear time invariant plant (1). The closed-loop system matrix $A_{*}+B_{*}K_{t}$ is $(\kappa,\gamma)$- strongly stable for $\kappa>0$ and $0<\gamma<1$ if there exists $H_{t}\succ I$ and $L_{t}$ such that $A_{*}+B_{*}K_{t}=H_{t}L_{t}{H}^{-1}_{t}$ and
  1. 1.
    $\|L_{t}\|\leq 1-\gamma$ and $\|K_{t}\|\leq\kappa$
  2. 2.
    $\|H_{t}\|\leq B_{0}$ and $\|{H}_{t}^{-1}\|\leq 1/b_{0}$ with $\kappa=B_{0}/b_{0}$.
    In that case, we say that $K_{t}$ strongly stabilizes the system.
  3. 3.
    A sequence $A_{*}+B_{*}K_{1},A_{*}+B_{*}K_{2},$... of closed-loop system matrices is $(\kappa,\gamma)$-strongly sequentially stable if each $A_{*}+B_{*}K_{t}$ is $(\kappa,\gamma)$- strongly stable and $\|H^{-1}_{t+1}H_{t}\|\leq 1+\gamma/2$.
Furthermore, we say a sequence $K_{1},K_{2},$... of control gains is $(\kappa,\gamma)$-strongly sequentially stabilizing for a plant $(A_{*},B_{*})$ if the sequence $A_{*}+B_{*}K_{1},A_{*}+B_{*}K_{2},$... is $(\kappa,\gamma)$-strongly sequentially stable.
In Appendix B.2, it is shown that any stabilizing policy $K$ is indeed $(\kappa,\gamma)$-strongly stabilizing for some values of $\kappa$ and $\gamma$. The only nontrivial component of this definition is the sequentiality condition, which informally requires that any two consecutively generated policies remain close to each other in a precise sense, formalized by the condition $\|H_{t+1}^{-1}H_{t}\|\leq 1+\gamma/2$. This requirement ensures that switching from one stabilizing policy to another at any time does not cause an immediate blow-up of the state norm.
While Definition 1 may appear purely definitional, ensuring that the closed-loop system remains stable under this definition requires specific measures and mechanisms in the algorithm design process. In this paper, we discuss how enforcing this requirement can lead to inefficiencies in the regret upper bound and how relaxing the sequential-stability condition can yield improvements.
2.2 Preliminaries
Our algorithm has two main building blocks, namely confidence ellipsoid construction and control design. In this section, we provide a brief overview of each.
Confidence Ellipsoid Construction To begin, we outline how a confidence ellipsoid around the true but unknown parameters $\Theta_{*}$ is constructed using the available measurements, leaving further details to the supplementary material. Let $X_{t}$ and $Z_{t}$ denote the matrices whose rows are $x_{1}^{\top},\ldots,x_{t}^{\top}$ and $z_{0}^{\top},\ldots,z_{t-1}^{\top}$, respectively. By applying least-square estimation, regularized with $\operatorname*{Tr}((\Theta-\Theta_{0})^{\top}(\Theta-\Theta_{0}))$ for some $\Theta_{0}$ and regularization parameter $\lambda$, the least-square estimate of $\Theta_{*}$ is given by:
$\hat{\Theta}_{t}=V_{t}^{-1}\left(Z_{t}^{\top}X_{t}+\lambda\Theta_{0}\right)$ (6)
where the covariance matrix is defined as $V_{t}=\lambda I+\sum_{s=0}^{t-1}z_{s}z_{s}^{\top}$.
Subsequently, for some given $\epsilon^{2}$ such that $\operatorname*{Tr}((\Theta_{*}-\Theta_{0})^{\top}(\Theta_{*}-\Theta_{0}))\leq\epsilon^{2}$ the confidence ellipsoid ${\mathcal{C}}_{t}(\delta)$ that contains $\Theta_{*}$ with probability at least $1-\delta$ is built as follows:
${\mathcal{C}}_{t}(\delta):=\big\{\Theta\in\mathbb{R}^{(n+m)\times n}|\operatorname{Tr}\big((\Theta-\hat{\Theta}_{t})^{\top}V_{t}(\Theta-\hat{\Theta}_{t})\big)\leq r_{t}\big\}$ (7)
$r_{t}=\bigg(\sigma_{\omega}\sqrt{2n\log\frac{n\det(V_{t})}{\delta\det(\lambda I)}}+\sqrt{\lambda}\epsilon\bigg)^{2}.$ (8)
cohen2019learning uses an additional parameter $\beta$ when constructing the confidence set with the goal of normalizing the ellipsoid, which is required for their analysis. Our algorithm follows the standard framework provided by abbasi2011regret and does not need such complication.
Control Design Consider the Primal and Dual SDP formulation of the LQR control problem revisited in Appendix B.1. By accounting for the confidence ellipsoid (7), we can relax those SDPs, to enable control design and analysis solely relying on data. The procedure of relaxing SDP formulations given such an ellipsoid is carried out by applying some perturbation technique, which has been illustrated in Lemma 8. The lemma generalizes the one in cohen2019learning, which is for a normalized confidence ellipsoid.
Given the parameter estimate in the form of confidence ellipsoid $\mathcal{C}_{t}(\delta)$ (or simply $\mathcal{C}_{t}$) defined by (7) the relaxed primal SDP is formulated as follows:
[…]
where $\mu_{t}=r_{t}+\sqrt{r_{t}}\vartheta\|V_{t}\|^{1/2}$. The rationale for this specific form of $\mu_{t}$ is explained in Lemma 7. Denoting optimal solution of program (12) by $\Sigma(\mathcal{C}_{t})$, the control signal extracted from solving relaxed primal SDP (12), $u=K(\mathcal{C}_{t})x$ is deterministic and linear in state where
$K(\mathcal{C}_{t})=\Sigma_{ux}(\mathcal{C}_{t}){\Sigma_{xx}^{-1}(\mathcal{C}_{t})}.$ (13)
Furthermore, the relaxed dual SDP formulation is given as follows:
[…]
where we denote its optimal solution by $P(\mathcal{C}_{t})$. Note that in cohen2018online, the term $\mu_{t}\|P\|_{*}V_{t}^{-1}$ on the right-hand side of their relaxed dual SDP introduces unnecessary looseness in the analysis and it can be replaced by $\mu_{t}\|P\|V_{t}^{-1}$. Our perturbation lemma (Lemma 7 in Appendix E), together with the corresponding lemma in their analysis, shows that such an SDP relaxation is not necessary.
Now we proceed to propose our algorithms followed by stability and regret bound guarantees, and highlighting the contributions.
3 Any-time Regret SDP-based LQ cOntrol (ARSLO) algorithm
[…]
In this section, we propose the anytime–regret–guaranteed algorithm ARSLO, which applies the OFU principle to generate control policies that ensure closed-loop stability of system (5) in the sense of the strongly sequentially stabilizing notion defined in Definition 1, while minimizing the regret (4). ARSLO requires an initial estimate $\Theta_{0}$ within an appropriate $\bar{\epsilon}(\bar{\kappa}_{1})$-neighborhood of $\Theta_{*}$, which is provided by the warm-up phase. Using the initial estimate, ARSLO algorithm builds and updates a confidence ellipsoid $\mathcal{C}_{t}$ around the true but unknown parameters of the system $\Theta_{*}$, applying an appropriate regularization parameter $\lambda$. The process for construction of $\mathcal{C}_{t}$ has already been illustrated in Section 2.2. Using this confidence set, the algorithm solves the relaxed primal SDP problem (12) and computes the policy $K(\mathcal{C}_{t})$ by (13). When applying the policy, the algorithm adds a zero-mean gaussian random perturbation noise $\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big)$ to the input, i.e., $u_{t}=K(\mathcal{C}_{t})x_{t}+\eta_{t}$. It is worth noting that the covariance matrix $\Gamma_{t}$ of the Gaussian perturbation noise computed by Theorem 1 is constructed leveraging the solution of the relaxed primal and dual SDPs, with a feature of non-isotropic exploration.
It can be shown that excessive updates in the policy can cause inefficiency, so there is a need for a criterion to decide on policy updates. By choosing $\det(V_{t})>(1+\beta)\det(V_{\tau})$ where $V_{\tau}$ is the covariance matrix of the most recently updated policy and arbitrarily setting $\beta=1$, we trigger a policy update when the ellipsoid volume is roughly halved. The parameters $\lambda$, $\Gamma_{t}$, and $\bar{\epsilon}(\bar{\kappa}_{1})$ are jointly tuned to ensure that the closed-loop system is $(\kappa_{*},\gamma_{*})$-strongly sequentially stable where $\kappa_{*}:=\sqrt{2\|P_{*}\|/\alpha_{0}}$ and $\gamma_{*}=\kappa_{*}^{-2}/2$, as demonstrated in Theorem 1. The rationale for choosing Gaussian perturbation noise with a covariance matrix $\Gamma_{t}$ such that $\|\Gamma_{t}\|=\mathcal{O}(1/\sqrt{t})$ is to ensure stability and achieve an overall regret of order $\mathcal{O}(\sqrt{t})$. Formally, the reasoning behind this choice is elaborated in Proposition 2. The algorithm enjoys any-time regret guarantees, as neither the initial estimate $\Theta_{0}$, the regularization parameter $\lambda$, nor the input perturbation noise $\eta_{t}$ depend on the time horizon of the algorithm’s execution.
3.1 Stability Analysis of ARSLO algorithm
In this subsection, we provide sufficient conditions for closed-loop stability under the deployment of control policies generated by Algorithm 1. The following theorem specifies how to tune the parameters $\Gamma_{t}$, $\lambda$, and $\bar{\epsilon}(\bar{\kappa}_{1})$ to guarantee that the system is $(\kappa_{*},\gamma_{*})$–strong sequentially stable.
Inputs: $\Theta_{0}$, $\bar{\kappa}_{1}$, and $\bar{\epsilon}(\bar{\kappa}_{1})$ (provided by Algorithm 2), $\delta$, $\vartheta$
Compute $\bar{c}$ using (19).
Set $\lambda=\frac{\sigma_{\omega}^{2}\bar{c}}{40}$
Initialize $V_{1}=\lambda I$, $r_{1}=\lambda\bar{\epsilon}^{2}(\bar{\kappa}_{1})$, $\hat{\Theta}_{1}=\Theta_{0}$, and construct $\mathcal{C}_{1}$.
for t=1,2,,Tt=1,2,...,T do
 if $\det(V_{t})>2\det(V_{\tau})$ or $t=1$ then
      Calculate $\mu_{t}=r_{t}+\sqrt{r_{t}}\vartheta\|V_{t}\|^{1/2}$
     Compute $K(\mathcal{C}_{t})$ by (13) and through solving the relaxed primal SDP problem (12)
     Compute $P(\mathcal{C}_{t})$ through solving the relaxed dual SDP problem (17)
     Let $V_{\tau}=V_{t}$
 else
     Set $K(\mathcal{C}_{t})=K(\mathcal{C}_{t-1})$ and $P(\mathcal{C}_{t})=P(\mathcal{C}_{t-1})$
   end if
  Calculate $\Gamma_{t}$ by (22) and sample $\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big)$
  Play $u_{t}=K(\mathcal{C}_{t})x_{t}+\eta_{t}$
  Observe $x_{t+1}$, save $(x_{t+1},z_{t})$ to the data set
  Construct $X_{t}$ and $Z_{t}$ and compute $V_{t}$
  Calculate $\hat{\Theta}_{t}$ by (6) and $r_{t}$ by (8) and build confidence ellipsoid $\mathcal{C}_{t}$
end for
Algorithm 1 Any-time Regret SDP-based LQ cOntrol (ARSLO)
Theorem 1 (Sequential Strong Stability of Algorithm 1)
Consider Algorithm 1, where the regularization parameter $\lambda$ is set as
$\lambda$ $:=\frac{\sigma_{\omega}^{2}\bar{c}}{80},$ (18)
$\bar{c}$ $:=a^{2}\Big(2\log\Big(\frac{4a^{2}n}{\delta}\Big)+1\Big)^{2},$ (19)
where
$a=\frac{40960\,\bar{\vartheta}^{2}_{B_{*}}\bar{\kappa}_{1}^{10}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}c_{z}(\bar{\kappa}_{1})}{\sigma_{\omega}^{2}},\quad c_{z}^{2}(\kappa)={64\kappa^{6}(1+\kappa^{2})}\bigg(\frac{\|x_{1}\|}{\log\frac{1}{\delta}}+\sqrt{20n\sigma_{\omega}^{2}}\bigg)^{2}$ (20)
and
$\bar{\kappa}_{1}$ $:=\sqrt{\kappa_{1}^{2}+\frac{1}{2\kappa_{1}^{2}}},\qquad\bar{\vartheta}_{B_{*}}:=\max\{1,\vartheta_{B_{*}}\},\quad\vartheta_{B_{*}}\geq\|B_{*}\|.$ (21)
Let the additive perturbation noise applied to the feedback control designed via the relaxed primal SDP be
$\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big),\qquad\Gamma_{t}:=\frac{\bar{p}_{t}\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}\Big)}{\sqrt{t+\bar{c}}},$ (22)
where
$\bar{p}_{t}=\frac{5120\,\kappa_{t}^{10}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$ (23)
Provided with an initial estimate $\Theta_{0}$ satisfying
$\|\Theta_{0}-\Theta_{*}\|_{F}\leq\bar{\epsilon}(\bar{\kappa}_{1}),\qquad\bar{\epsilon}(\bar{\kappa}_{1}):=\frac{\sigma_{\omega}\sqrt{2n(n+m)}}{\sqrt{\lambda}},$ (24)
the closed-loop system under the policies generated by Algorithm 1 is $(\kappa_{*},\gamma_{*})$-sequentially strongly stable, with
$\kappa_{*}:=\sqrt{\frac{2\|P_{*}\|}{\alpha_{0}}},\qquad\gamma_{*}:=\frac{1}{2\kappa_{*}^{2}}.$ (25)
Note that $\bar{\epsilon}(\bar{\kappa}_{1})$, given by (24), is parametrized by $\bar{\kappa}_{1}$, which in turn depends on the solution of the relaxed dual SDP $P(\mathcal{C}_{1})$ at first time step, or equivalently on $\kappa_{1}:=\sqrt{2\|P(\mathcal{C}_{1})\|/\alpha_{0}}$. The purpose of this parametrization, which determines how tight the initial estimate $\Theta_{0}$ should be, is to eliminate the need for an a priori bound on $\|P_{*}\|$, a limitation present in existing OFU-based algorithms such as abeille2020efficient and cohen2019learning. This relaxation is achieved by designing a warm-up phase that iteratively constructs a confidence ellipsoid (ball), solves the relaxed dual SDP for that, computes $\bar{\epsilon}(\bar{\kappa}_{1})$ from its solution, and checks if the confidence ball falls within the desired $\bar{\epsilon}(\bar{\kappa}_{1})$ neighborhood, and then out puts such an appropriate $\Theta_{0}$ and $\bar{\epsilon}(\bar{\kappa}_{1})$ to be used in ARSLO algorithm.
This relaxation is primarily implied by Lemma 15, which shows that the ARSLO algorithm always satisfies
$P(\mathcal{C}_{t})\preceq P_{*}\preceq P(\mathcal{C}_{t})+\frac{\alpha_{0}}{4\kappa_{t}^{2}}I,\quad\forall t.$ (26)
We refer the reader to the proof of Theorem 1 for further details.
For a sequence of policies generated by Algorithm 1 that guarantees $(\kappa_{*},\gamma_{*})$-strong sequential stability of the closed-loop system, one can show that the state norm remains bounded. The following proposition summarizes this statement.
Proposition 1
The state norm of the closed-loop system under the deployment of the ARSLO algorithm is bounded as
$\|x_{t}\|$ $\leq\kappa_{*}e^{-{\gamma_{*}}(t-1)/2}\|x_{1}\|+\frac{2\kappa_{*}}{\gamma_{*}}\sqrt{20\,\sigma_{\omega}^{2}n\log\frac{t}{\delta}}$ (27)
with probability at least $1-\delta$.
As an immediate consequence of Theorem 1, the following corollary concludes the required tightness of the initial estimate $\Theta_{0}$.
Corollary 1
For the ARSLO algorithm to generate strongly sequentially stabilizing policies, the initial estimate $\Theta_{0}$ should lie within a $\bar{\epsilon}(\bar{\kappa}_{1})$-neighborhood of $\Theta_{*}$ of order $\mathcal{O}\!\left({1}/{\kappa_{*}^{14}}\right)$.
3.2 Regret Bound Analysis of ARSLO algorithm
In this section, we provide an upper bound for the regret of the ARSLO algorithm, denoted by $R_{\text{ARSLO}}(t)$, and present discussions and comparisons with selected leading works from the literature. The following theorem summarizes the regret bounds for ARSLO algorithm.
Theorem 2
Fix $\delta\in(0,1)$. With probability at least $1-\delta$, the following statement holds true for all $t\geq-200\log\delta$:
$R_{\text{ARSLO}}(t)\leq\mathcal{O}\big(\sqrt{\;n^{2}(n+m)^{3}\|P_{*}\|^{16}\;t\log^{4}\frac{t}{\delta}}\big)$ (28)
It is worth mentioning that the regret terms involving $\lambda$ appear only in expressions of order $\mathcal{O}(\log\frac{t}{\delta})$. Consequently, they do not contribute to the $\mathcal{O}(\sqrt{t})$ terms, which play a pivotal role in the regret upper bound. A detailed expression of the regret upper bound can be found in Appendix H.
Since our proposed algorithm belongs to the class of OFU-based methods, it is natural to compare it with existing convex (i.e., computationally efficient) OFU-based algorithms in the literature. To the best of our knowledge, the only two algorithms of this type are those proposed in cohen2019learning and abeille2020efficient. The former, known as the OSLO algorithm, is technically involved and requires a warm-up phase of length $\mathcal{O}(\sqrt{T})$ for implementation up to round $T$. In addition, its regret bound exhibits a higher-order dependence on the dimension, namely $\mathcal{O}\left(\sqrt{n^{4}(n+m)^{3}}\right)$, and depends on an a priori bound of $\|P_{*}\|$. A similar issue exists in the latter work, which proposes the LAG-LQ algorithm, achieving the regret bound
$R_{{LAG-LQ}}\leq\mathcal{O}\left({D^{\frac{3}{2}}}\sqrt{{n(n+m)^{2}}{\|P_{*}\|^{4}}{T}\log^{3}\frac{T}{\delta}}\right)$
where $\operatorname{Tr}(P_{*})\leq{D}$ is an a priori known bound. Unlike the LAG-LQ algorithm, our method does not require such an a priori bound; accordingly, no such quantity appears in the regret bound of ARSLO. This is indeed achieved by jointly leveraging the relaxed primal and dual SDP solutions when tuning the input perturbation covariance matrix, and by employing a carefully designed warm-up phase that repeatedly solves the relaxed dual SDP problem while using the computed upper bound for the operator norm of the DARE solution as given in (26). In this respect, our algorithm offers a clear advantage over the method of abeille2020efficient, in addition to its primary strength of coming with anytime regret guarantees. Moreover, the use of both primal and dual solutions further facilitates non-isotropic exploration. As for the dependence on the system ambient dimension and the operator norm of the DARE solution, note that, even in favor of the LAQ-LQ algorithm, if we let $D=\operatorname{Tr}(P_{*})\leq n\|P_{*}\|$, it has a dependence of $\mathcal{O}\!\bigl(\sqrt{n^{4}(n+m)^{2}\|P_{*}\|^{7}}\bigr)$. On the other hand, the dependence of the regret of the ARSLO algorithm on the dimension and on the solution of the DARE is $\mathcal{O}\bigl(\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{16}}\bigr)$. Leaving aside the fact that we conduct this comparison in favor of the LAG-LQ algorithm by setting $D=\operatorname{Tr}(P_{*})$, the results nevertheless indicate that the dependence on the dimension for ARSLO is of comparable order, or in some cases even better. However, our algorithm exhibits a stronger dependence on $\|P_{*}\|$, suggesting that there may still be room for improvement.
We further compare our proposed algorithm with another class of methods, namely the Certainty-Equivalence (CE) based algorithms. For instance, simchowitz2020naive achieves a high-probability regret bound of
$R_{\text{CE}}\leq\mathcal{O}\left(\sqrt{nm^{2}\|P_{*}\|^{11}T\log\frac{T}{\delta}}\right).$ (29)
Unlike our proposed algorithm, their method does not provide an anytime regret guarantee. However, its dependence on the system dimensions $n$ and $m$, as well as on the operator norm of the DARE solution $P_{*}$, is lower than that of our method. Another CE-based approach proposed by jedra2022minimal can run in an anytime fashion, but it only provides guarantees on the expected cost. It also improves the dependence on $\|P_{*}\|$ compared to simchowitz2020naive, reducing the exponent from $5.5$ to $5.25$.
While these CE-based methods ensure Lyapunov stability, they do not explicitly control the evolution of the state with high probability. In contrast, a specific advantage of our algorithm is that it provides an explicit high-probability bound on the state trajectory. Such bounds are motivated by the fact that policy updates may induce transient spikes in the state norm—a phenomenon where switching from one stabilizing controller to another can temporarily increase the state magnitude. Although these spikes have little impact on the overall regret, as their contribution is absorbed into the accumulated cost, they may be undesirable in systems with safety constraints. Both the OSLO algorithm of cohen2019learning and our ARSLO algorithm incorporate the notion of strong sequential stability, which ensures that consecutively generated policies remain sufficiently close. This mechanism provides control over the state trajectory and may be particularly valuable when transferring the algorithm to different control domains where such guarantees are critical. While this feature improves trajectory control, it may not yet be the most efficient in terms of the regret bound’s dependence on $\|P_{*}\|$.
In the next section, we pursue the idea of removing the notion of strong sequential stability—which introduces conservatism into the algorithm design—while still ensuring that the closed-loop system remains stable with an explicit high-probability bound on the state-norm trajectory, along with an improved (i.e., lower) dependence of the regret bound on $\|P_{*}\|$.
[…]
The strong sequential-stability requirement mandates that each individually generated policy be stabilizing and that consecutively generated policies under the ARSLO algorithm remain sufficiently close, namely, $\|P^{-1/2}(\mathcal{C}_{t+1})P^{1/2}(\mathcal{C}_{t})\|\leq 1+\gamma_{*}/2$. Satisfying these condition necessitates a conservative algorithmic design, implemented by enforcing
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{32{\kappa_{t}^{8}}}I$ (30)
for any time $t$, where $\kappa_{t}=\sqrt{2\|P(\mathcal{C}_{t})\|/\alpha_{0}}>1$. However, this enforcement is particularly restrictive, as it induces a higher-order dependence of the regret upper bound on $\|P_{*}\|$.
In this section, we address this limitation by relaxing the sequentiality requirement (30) to the following condition:
$\mu_{t}\,\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I.$ (31)
for any $t$ and some $\bar{\rho}\in[0,\,8)$. Enforcing this condition guarantees that any policy generated by the algorithm is $(\kappa_{*},\gamma_{*})$–strongly stabilizing and that consecutively generated policies remain within a controlled distance of one another. However, this distance requirement is weaker than that imposed by strong sequential stability, which corresponds to the case $\bar{\rho}=8$. Under this relaxation, the closed-loop system’s state norm may not remain adequately controlled and can exhibit undesired spikes under arbitrary switching. As established in switched-systems control theory, switching among individually stabilizing policies can still induce instability or lead to temporary explosions of the state norm. A standard remedy in this setting is to sufficiently slow down the switching. Motivated by this idea, for a given $\bar{\rho}\in[0,,8)$, the frequency of policy updates must be appropriately limited. This is achieved by tuning the exploration–exploitation trade-off parameter $\beta(\bar{\rho})$ in the update criterion
$\det(V_{t})>(1+\beta(\bar{\rho}))\det(V_{\tau}).$
Viewing $\bar{\rho}=8$, strong sequential stability as one extreme, the case $\bar{\rho}=0$ represents the opposite extreme, imposing only the minimal requirement that each generated policy be individually strongly stabilizing. The key question, then, is which value of $\bar{\rho}$ achieves the optimal regret. Intuitively, setting $\bar{\rho}$ to a smaller value increases the distance between consecutively generated policies $\|P^{-1/2}(\mathcal{C}_{t^{\prime}})P^{1/2}(\mathcal{C}_{t})\|\leq 1+2{\kappa_{*}^{3-\bar{\rho}/2}}$ (by Lemma 24 and Remark 4), which requires a longer dwell-time between policy updates, i.e., larger value of $\beta(\bar{\rho})$, potentially increasing the upper bound of the regret term $R_{4}(t)$. Conversely, a larger value of $\bar{\rho}$ increases the order of dependence of the perturbation-noise covariance matrix $\Gamma_{t}$ on $\|P_{*}\|$, which raises the upper bound of the regret term $R_{6}(t)$. This clearly illustrates the inherent trade-off and motivates a general analysis to identify the optimal value $\bar{\rho}_{*}$ that achieves the best regret bound in terms of its dependence on $\|P_{*}\|$. This analysis also allows for a careful examination of the trade-off between the regret bound and the high-probability state-norm guarantees. The rationale for specifying $\beta(\bar{\rho})$ is provided in (40) of Theorem 4.
[…]
Theorem 3
Fix $\bar{\rho}\in[0,\,8)$. Let Algorithm 3, set the regularization parameter $\lambda$ as
$\lambda$ $=\frac{\sigma_{\omega}^{2}\bar{c}}{80}(n+m)\log\frac{1}{\delta},$ (32)
$\bar{c}$ $=a^{2}\Big(2\log\Big(\frac{4a^{2}n}{\delta}\Big)+1\Big)^{2},$ (33)
where
$a$ $=\frac{10240\bar{\vartheta}_{B_{*}}^{2}\bar{\kappa}_{1}^{\bar{\rho}+4}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}c_{z}(\bar{\kappa}_{1},\bar{\rho})}{\sigma_{\omega}^{2}},$
$c_{z}^{2}(\kappa,\bar{\rho})$ $={64\kappa^{6}(1+\kappa^{2})}(1+\xi(\kappa,\bar{\rho}))^{2}\bigg(\frac{\|x_{1}\|}{\log\frac{1}{\delta}}+\sqrt{20n\sigma_{\omega}^{2}}\bigg)^{2},$ (34)
$\xi(\kappa,\bar{\rho})=2\kappa^{3-\bar{\rho}/2}$ and
$\bar{\kappa}_{1}(\bar{\rho}):=$ $\sqrt{\kappa_{1}^{2}+4\kappa_{1}^{6-\bar{\rho}}},\quad\bar{\vartheta}_{B_{*}}:=\max\{1,\vartheta_{B_{*}}\},\quad\vartheta_{B_{*}}\geq\|B_{*}\|.$ (35)
Further, let the additive perturbation noise applied to the feedback control designed via the relaxed primal SDP be
$\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big),\qquad\Gamma_{t}:=\frac{\bar{p}_{t}\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}\Big)}{\sqrt{t+\bar{c}}},$ (36)
where
$\bar{p}_{t}=\frac{640\kappa_{t}^{\bar{\rho}+2}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$ (37)
Provided with an initial estimate $\Theta_{0}$ satisfying
$\|\Theta_{0}-\Theta_{*}\|_{F}\leq\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho}),\qquad\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho}):=\frac{\sigma_{\omega}\sqrt{2n(n+m)}}{\sqrt{\lambda}},$ (38)
any policy generated by Algorithm 1 is $(\kappa_{*},\gamma_{*})$-strongly stabilizing, with
$\kappa_{*}:=\sqrt{\frac{2\|P_{*}\|}{\alpha_{0}}},\qquad\gamma_{*}:=\frac{1}{2\kappa_{*}^{2}}.$ (39)
The rationale for expressing the parameter $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ in terms of $\bar{\kappa}_{1}$ follows the same reasoning as in the ARSLO algorithm; this parametrization ultimately allows us to eliminate the need for any a priori bound on $\|P_{*}\|$ through a carefully designed warm-up phase.
The following theorem characterizes the appropriate choice of the policy update parameter $\beta(\bar{\rho})$ that guarantees boundedness of the closed-loop system state norm.
Theorem 4
[…]
$\beta(\bar{\rho})=\begin{cases}2(\kappa_{1}^{2}+\kappa_{1}^{(6-\bar{\rho})})^{\frac{(6-\bar{\rho})}{4}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8\end{cases}.$ (40)
where $\kappa_{1}=\sqrt{2\|P(\mathcal{C}_{1})\|/\alpha_{0}}$. Then, under the above update mechanism, the closed-loop state is bounded as follows:
$\|x_{t}\|\leq\kappa_{*}(1-\gamma_{*})^{\frac{t-1}{2}}\|x_{1}\|+{\frac{2\kappa_{*}(1+\xi(\kappa_{*},\bar{\rho}))}{\gamma_{*}}\sqrt{20n\sigma_{\omega}^{2}\log\frac{2t}{\delta}}}$ (41)
where $\xi(\kappa_{*},\bar{\rho})=2{\kappa_{*}^{3-\bar{\rho}/2}}$.
Remark 1
Note that for $\bar{\rho}\geq 8$, the closed-loop stability guarantees hold for any choice of $\beta(\bar{\rho})>0$. As we observed earlier, in the case $\bar{\rho}=8$, which corresponds exactly to the strong sequential stability scenario, we arbitrarily set $\beta(\bar{\rho})=1$. In fact, it can be set to any positive value if our goal is solely to ensure stability in the sense of a bounded state norm. However, as we will discuss later, the choice of $\beta(\bar{\rho})$ contributes additive terms to the regret.
Corollary 2
[…]
$h=\frac{\big(\max\{3-\frac{\bar{\rho}}{2},\;1\}\big)(22+\bar{\rho})}{2}.$ (42)
[…]
Theorem 5
Fix $\delta\in(0,1)$ and let $\bar{\rho}\in[0,\,8)$. Then, with probability at least $1-\delta$, the following holds for all $t\geq-200\log\delta$:
$R_{\text{ARSLO}^{+}(\bar{\rho})}(t)\leq$ $\mathcal{O}\Big(\sqrt{\max\bigg\{\|P_{*}\|^{9-\frac{\bar{\rho}}{2}}\,\beta_{*}^{2}\big(\|P_{*}\|,\bar{\rho}\big)\,\;,\|P_{*}\|^{(11+\frac{\bar{\rho}}{2})}\bigg\}n^{2}(n+m)^{3}\,t\;\log^{4}\frac{t}{\delta}}\,\Big)$ (43)
where
$\beta_{*}\big(\|P_{*}\|,\bar{\rho}\big)=\begin{cases}\big(\|P_{*}\|+\|P_{*}\|^{(6-\bar{\rho})/2}\big)^{\frac{(6-\bar{\rho})}{4}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8.\end{cases}$
For $\bar{\rho}_{*}=2$, the regret bound achieves its optimum:
$R_{\text{ARSLO}^{+}(\bar{\rho}_{*})}(t)\leq$ $\mathcal{O}\Big(\sqrt{\,n^{2}(n+m)^{3}\;\|P_{*}\|^{12}\;t\;\log^{4}\frac{t}{\delta}}\,\Big).$ (44)
[…]
5 Warm-up phase
[…]
$\|\Theta_{0}-\Theta_{*}\|_{F}\leq\epsilon.$
For ARSLO, it has been shown that $\epsilon=\bar{\epsilon}(\bar{\kappa}_{1})$, where $\bar{\epsilon}(\bar{\kappa}_{1})$ is given in (24), and
$\bar{\kappa}_{1}=\sqrt{\kappa_{1}^{2}+\frac{1}{2\kappa_{1}^{2}}},$
as defined in (21).
[…]
$\bar{\kappa}_{1}=\sqrt{\kappa_{1}^{2}+4\kappa_{1}^{6-\bar{\rho}}}.$
For both cases, $\kappa_{1}=\sqrt{{2\|P(\mathcal{C}_{1})\|}/{\alpha_{0}}}$ where $P(\mathcal{C}_{1})$ is the solution of the relaxed dual SDP program associated with the initial ellipsoid $\mathcal{C}_{1}$. This ellipsoid is precisely the desired confidence ball $\|\Theta-\Theta_{0}\|_{F}\leq\epsilon$. Therefore, the warm-up phase must perform sufficient system identification to produce such a confidence ball and verify that this ball lies inside the required initial neighborhood. The following algorithm describes this warm-up procedure.
Algorithm 2, employs an initial strongly stabilizing policy $K_{0}$ and
[…]
Set $\varepsilon=c$// $c$ is chosen sufficiently small and is used only for initialization
while $\epsilon_{w}(t)>\varepsilon$ do
  Receive state $x_{t}$
  Play input $u_{t}=K_{0}x_{t}+\nu_{t}$
  Observe $x_{t+1}$
  Compute the least-squares estimate $\hat{\Theta}_{t}^{w}$ via (45)
  Compute the radius $r_{w}(t)$ via (46) and use it to compute $\epsilon_{w}(t)$ by (47)
  Construct the confidence ellipsoid (ball) $\mathcal{C}_{t}^{w}$ centered at $\hat{\Theta}_{t}^{w}$ with radius $\epsilon_{w}(t)$
  Solve the dual SDP for $P(\mathcal{C}_{t}^{w})$
  Compute $\kappa_{1}=\sqrt{{2\|P(\mathcal{C}_{t}^{w})\|}/{\alpha_{0}}}$
[…]
   For ARSLO Compute $\bar{\kappa}_{1}$ and $\bar{\epsilon}(\bar{\kappa}_{1})$ using (21) and (24) then set $\varepsilon=\bar{\epsilon}(\bar{\kappa}_{1})$
[…]
  Set $\Theta_{0}=\hat{\Theta}_{t}^{w}$
end while
Output: $\Theta_{0}$, $\bar{\kappa}_{1}$, and $\bar{\epsilon}(\bar{\kappa}_{1})$ (or $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$)
Algorithm 2 Warm-Up Phase
applies the input $u_{t}=K_{0}x_{t}+\nu_{t}$, where $\nu_{t}\sim\mathcal{N}(0,2\sigma_{\omega}^{2}\kappa_{0}^{2}I)$. It utilizes the saved data $z_{t}$ and $x_{t}$ to construct a confidence ellipsoid, similar to the one described in Section 2.2, with the difference that the least-squares estimation is regularized with respect to $\|\Theta\|^{2}_{F}$ using the regularization parameter $\rho$. The center of this ellipsoid is computed by
$\hat{\Theta}_{t}^{w}$ $=\bar{V}_{t}^{-1}\left(Z_{t}^{\top}X_{t}\right)$ (45)
$\bar{V}_{t}$ $=\rho I+Z_{t}^{\top}Z_{t}$
and
$r^{w}_{t}=\bigg(\sigma_{\omega}\sqrt{2n\log\frac{n\det(\bar{V}_{t})}{\delta\det(\lambda I)}}+\sqrt{\rho}\psi_{\Theta_{*}}\bigg)^{2}.$ (46)
where $Z_{t}$ and $X_{t}$ are constructed as described in the Section 2.2. Also, $\|\Theta_{*}\|_{F}\leq\psi_{\Theta_{*}}$, where one may set $\psi_{\Theta_{*}}=\sqrt{n}\vartheta$. The regularization parameter is chosen as $\rho=\sigma_{\omega}^{2}\vartheta^{-2}$.
[…]
Theorem 6
Let $K_{0}$ be a $(\kappa_{0},\gamma_{0})$-strongly stabilizing policy and let $\nu_{t}\sim\mathcal{N}(0,2\sigma_{\omega}^{2}\kappa_{0}^{2}I)$ be an input perturbation. Then, Algorithm 2 (the ‘warm-up phase’), by applying the control input $u_{t}=K_{0}x_{t}+\nu_{t}$, outputs $\Theta_{0}$ such that $\|\Theta_{0}-\Theta_{*}\|_{F}\leq\varepsilon$ if the following condition met
$\epsilon_{w}(t)=\sqrt{\frac{r^{w}_{t}}{\lambda_{min}(V_{t})}}\leq\varepsilon$ (47)
where
[…]
The following remark specifies the scaling of the warm-up duration $t_{w}$.
Remark 2
From the preceding analysis in Theorem 7, we have $\lambda_{\min}(\bar{V}_{t})=\mathcal{O}(t)$ and $r_{t}^{w}(t)=\mathcal{O}(\log t)$. Hence, there exists a finite time $t_{w}$ such that condition (47) is satisfied. Moreover, from (47), it follows directly that the warm-up phase duration satisfies $t_{w}=\mathcal{O}(1/\varepsilon^{2})$.
[…]
Theorem 7
With probability at least $1-\delta$, the following statements hold.
The regret incurred during the warm-up phase of the ARSLO algorithm satisfies
$R_{\text{warm-up}}(t_{w})\leq\mathcal{O}\big(\frac{\kappa_{0}^{6}\,\alpha_{1}\,\bar{\vartheta}_{B_{*}}^{8}}{\gamma_{0}^{2}}(n+m)^{5}\|P_{*}\|^{14}\big),$
[…]
$R_{\text{warm-up}}(t_{w},\bar{\rho})\leq\mathcal{O}\big(\frac{\kappa_{0}^{6}\,\alpha_{1}\,\bar{\vartheta}_{B_{*}}^{8}}{\gamma_{0}^{2}}(n+m)^{6}\|P_{*}\|^{h}\big)$
where $h$ is defined in (42).
6 Discussion and Conclusion
In this section, we discuss the conceptual foundations underlying the findings of this paper to provide a clearer understanding of the challenges involved in algorithm design. One specific discussion concerns the rationale behind the particular choice of perturbation noise, Gaussian $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, with $\|\Gamma_{t}\|=\mathcal{O}\!\big(1/\sqrt{t}\big)$. The following proposition formally justifies why this order of the covariance matrix is desirable for achieving a favorable regret upper bound, and explains how deviations from this order can worsen the regret upper-bound. This choice of scaling has previously been employed in self-tuning–regulator–based algorithms, including CE-based approaches such as simchowitz2020naive and jedra2022minimal.
Proposition 2
Consider the perturbation noise $\eta_{t}\sim\mathcal{N}\!\left(0,\Gamma_{t}\right)$ applied to the designed feedback control in the proposed algorithms. Any deviation of $\|\Gamma_{t}\|$ from the order $\mathcal{O}\!\left(1/\sqrt{t}\right)$ either prevents satisfaction of the sufficient stability conditions or increases the regret upper bound beyond order $\mathcal{O}(\sqrt{t})$.
In most existing algorithms, due to their specific regret decomposition, the state norm does not directly affect the dominant terms of the regret and usually appears only in $\mathcal{O}(\log T)$ terms. As a result, these works typically do not aim to derive tight high-probability bounds on the state norm trajectory. Their analyses show that stability in the Lyapunov sense is sufficient to control the regret bound, and they often provide only crude bounds on the state norm by arguing that as the length of epochs increases, the effect of policy updates diminishes in the long run. In contrast, we found that carefully studying the boundedness of the state norm is necessary, as it provides strong control over the state trajectory and makes the algorithm, with its anytime regret guarantee, transferable to a wider range of control settings. In our work, we observed that while imposing a strong sequentiality notion can help to explicitly talk about teh state norm bound, the resulting regret bound might not be optimal. This is because strong sequentiality requires additional constraints on the algorithm to keep subsequently generated policies close to each other. We proposed an alternative idea: let the algorithm generate only stabilizing policies, and handle concerns about state explosion due to policy updates using a dwell-time–inspired approach. We noticed that if the focus is purely on minimizing regret, this extreme approach is also inefficient. The optimal regret bound is achieved when the algorithm ensures that two subsequently generated policies are neither too far apart nor too close. This reveals a fundamental trade-off between state norm control and regret upper bounds. Our algorithm lies at the intersection of two broad classes of methods: self-tuning regulator–based algorithms, such as CE-based approaches mania2019certainty; simchowitz2020naive; jedra2022minimal, and OFU-based algorithms, such as abeille2020efficient; cohen2019learning. While our method is in the spirit of OFU, it also incorporates elements of input perturbation methods. This combination allowed us to overcome one of the remaining challenges of computationally efficient OFU-based algorithms: the requirement for an a priori bound on the optimal average expected cost, i.e., $\|P_{*}\|\leq D$. Such a requirement is unnecessary for self-tuning regulator–based algorithms, and by bridging these two approaches, we were able to remove this assumption for OFU-based methods.
Acknowledgments and Disclosure of Funding
This work was supported in part by a C3.AI DTI Research Grant.
Content of Appendices
Appendix A Discussion of Related Works
The other useful and practical categorization of regret guaranteed algorithms for LQ control, also adopted by jedra2022minimal, is to group existing methods under two main umbrellas: (i) self-tuning regulators and (ii) OFU-based approaches.
Self-tuning regulator–based algorithms operate as follows. At each step, they estimate the unknown matrices $(A_{*},B_{*})$, treat these estimates as the true parameters, solve the DARE, and design the control accordingly. To ensure sufficient excitation of the system—and thus enable learning of $(A_{*},B_{*})$—the control inputs are typically perturbed with appropriately tuned noise. Many existing works, including mania2019certainty, faradonbeh2020input, simchowitz2020naive, and jedra2022minimal, fall within this category, and these methods are computationally efficient.
The second class consists of OFU-based algorithms, in which a high-probability confidence set for $(A_{*},B_{*})$ is constructed, and a control is designed by playing optimistically with respect to this set. Compared to robust control schemes, which follow a min–max paradigm, these methods can be viewed as min–min algorithms. OFU-based approaches were initially and predominantly developed in the context of bandits auer2002finite; abbasi2011improved. The first application of such ideas to the LQ control setting appears in campi1998adaptive, which provides only asymptotic guarantees. Along this line of research, abbasi2011online proposed the first regret-guaranteed algorithm for LQ control. Later works aimed to relax certain assumptions or burdens of this algorithm. For example, faradonbeh2017finite extends the approach to stabilizable systems and relaxes the Gaussian process noise assumption to a class of heavy-tailed noise with arbitrary correlation structures. lale2022reinforcement reduces the dimensional dependence of the regret upper bound from exponential to polynomial. However, neither of these works addresses the most fundamental drawback of abbasi2011online, namely its computational inefficiency. In response to this, cohen2019learning and abeille2020efficient proposed computationally efficient algorithms. The former is based on a relaxation of the SDP-based formulation of LQ control using the constructed confidence ellipsoid, while the latter relies on extended value iteration (EVI) followed by a Lagrangian relaxation. To the best of our knowledge, these two works constitute the only computationally efficient OFU-based algorithms for LQ control.
Our work sits precisely at their intersection of these categories. On one hand, it embodies the spirit of OFU-based algorithms, building a confidence ellipsoid and playing optimistically with respect to this set to design the controller. On the other hand, our proposed algorithm applies carefully tuned input perturbation noise, which enables guaranteed regret bounds and stability in real time.
We are primarily interested in conducting the comparison within the category of computationally efficient algorithms. Our contributions can be summarized under three main points: (i) establishing an anytime regret guarantee; (ii) providing a high-probability bound on the state trajectory and discussing the trade-offs that arise when relaxing strong sequential stability; and (iii) removing the need for any a priori bound on the optimal average expected cost $J_{*}$, or equivalently, on $\|P_{*}\|$.
One important aspect of our algorithm is its ability to provide an anytime regret guarantee. Most existing methods are designed for a fixed horizon, and their extension to the infinite-horizon setting typically relies on the doubling trick. However, this approach does not fully resolve the issue, since each epoch still operates with a fixed horizon. As discussed earlier, the foundational line of regret-guaranteed algorithms begins withabbasi2011regret, which offers an anytime guarantee but suffers from high computational complexity and an exponential dependence on the system dimension. Although more recent methods have addressed the computational challenges of that approach, they remain horizon-dependent, whereas abbasi2011regret inherently provides an anytime guarantee. In contrast, our method is both computationally efficient and anytime. Among existing Self-tuning regulator–based works, the only method that has anytime-style guarantee is the one by jedra2022minimal, but its guarantees apply only to the expected regret. In the line of OFU-based algorithms, to the best of our knowledge, the only two computationally efficient existing methods are those of cohen2019learning and abeille2020efficient; however, both are horizon-dependent. This makes our approach the first computationally efficient OFU-style algorithm that also provides an anytime regret guarantee.
The next important guarantee concerns whether an algorithm can provide a fully system-theoretic high-probability upper bound on the state norm trajectory. While overall stability is ensured by all the existing algorithms which is appeared to be enough to guarantee the promised regret bound, however, an explicit high probability bound on state norm is still need attention. In CE-based methods jedra2022minimal and simchowitz2020naive the overall stability is established using a common Lyapunov function. However, the former work does not provide an explicit high probability upper-bound on state norm, while the latter introduces a crude bound which is not precise. We refer the readers to Lemma 5.3 of simchowitz2020naive and its proof, where the crude bound is obtained by apply $\max_{k\geq 1}\;k(1-\rho)^{k}\;\lesssim\;\frac{1}{\rho}$ for some $\rho<1$ but we know that $\lesssim$ absorbs a constant that might be dependent on $\rho$. This crude bound does not affect the regret bound as it appears in $\mathcal{O}(\log T)$ terms. Even when an algorithm generates stabilizing policies, updating from one policy to another can cause temporary spikes, since the next generated policy is not necessarily better than the current one unless the algorithm is explicitly designed to prevent this. Such an enforcement has been proposed by faradonbeh2017finite and abeille2020efficient. In these approaches, an initial system identification phase is used to obtain a sufficiently accurate estimate, followed by the fulfillment of explicit accuracy conditions to guarantee sequential stability. While the former approach does not provide an explicit state norm bound, the latter provides a state norm bound that relies on an a priori bound on $J_{*}$, making it not fully system-theoretic and therefore less reliable. Similarly, cohen2019learning introduced the notion of strong sequential stability. In this framework, the algorithm is designed to generate stabilizing policies while ensuring that any two consecutively generated policies remain sufficiently close, thereby preventing spikes in the state norm during policy updates. This technique requires an appropriate initialization and the fulfillment of explicit conditions during the implementation of the main algorithm. Overall, the OFU-based algorithms in faradonbeh2017finite, abeille2020efficient, and cohen2019learning share a common strategy: obtaining a sufficiently accurate initial estimate, followed by the enforcement of explicit conditions thereafter. While we provide an anytime and fully system-theoretic version of the guarantee of cohen2019learning, we also show that this strategy is not efficient in terms of the regret bound. By incorporating dwell-time techniques, we demonstrate how to improve the regret bound while maintaining explicit control over the state trajectory.
We also compare algorithms in terms of the a priori information or bounds required for their implementation. In particular, we focus on requirements such as an upper bound on $\|P_{*}\|$ and the horizon, which are explicitly indicated in a separate column of the table specifying whether the algorithm is anytime or not. Notably, in our proposed algorithm, the parameter $\delta$ is fully independent of the implementation time and can be set to any arbitrary value, making it effectively non-restrictive. Almost all computationally efficient algorithms require an initial stabilizing policy $K_{0}$. In this work, we eliminate the need for prior access to the optimal average cost $J_{*}$, a requirement commonly present in existing computationally efficient OFU-based algorithms. While this issue has been addressed in certainty-equivalence methods, as studied in simchowitz2020naive; jedra2022minimal, it remains unresolved in current computationally efficient OFU-based algorithms, including abeille2020efficient; cohen2019learning. Such a requirement makes the performance of these algorithms dependent on an external bound that may be loose, which is reflected in their regret upper bounds. We address this limitation by introducing a warm-up phase that iteratively solves the dual SDP problem to detect when the parameter estimate enters the desired ball, whose radius depends on $\|P_{*}\|$.
Table 1 compares our algorithm with several notable algorithms from the literature. We specifically selected these works because they either provide an explicit regret upper bound or their estimated regret upper bound has been reported by other studies.
Table 1: Comparison of existing methods.
Paper Method Regret upper-bound w.p. $1-\delta$ Any-time State norm w.p.
Required
information
simchowitz2020naive CE $\widetilde{\mathcal{O}}\!\left(\sqrt{nm^{2}\|P_{*}\|^{11}T\log(1/\delta)}\right)$ $\times$ $\times$ $K_{0}$
jedra2022minimal CE $\widetilde{\mathcal{O}}\!\left(\sqrt{nm^{2}\|P_{*}\|^{10.5}T\log^{2}T}\right)$ $\checkmark$ $\times$ $K_{0}$
cohen2019learning OFU $\widetilde{\mathcal{O}}\!\left(\sqrt{n^{4}(n+m)^{3}D^{10}T\log^{4}(T/\delta)}\right)$ $\times$ $\checkmark$ $K_{0},Tr(P_{*})\leq D$
abeille2020efficient OFU $\widetilde{\mathcal{O}}\!\left(\sqrt{n(n+m)^{2}D^{3}\|P_{*}\|^{4}T\log^{3}(T/\delta)}\right)$ $\times$ $\times$ $K_{0},Tr(P_{*})\leq D$
ARSLO OFU $\widetilde{\mathcal{O}}\!\left(\sqrt{n^{3}(n+m)^{2}\|P_{*}\|^{16}T\log^{4}(T/\delta)}\right)$ $\checkmark$ $\checkmark$ $K_{0}$
[…]
  • This regret bound holds in expectation.
  • These works provide only a crude high-probability bound on the state norm.
Appendix B Preliminaries
B.1 Primal and Dual-SDP formulation of LQR
Consider the standard LQR problem of (1) and (2). Assuming controllability of $(A_{*},B_{*})$, it is known that the optimal sequence minimizing average expected cost (3) is $u^{*}=\{u^{*}_{t}\}_{t=1}^{T-1}$, where $u^{*}_{t}=K_{*}(\Theta*)x_{t}$ and
$K_{*}(\Theta_{*})=-(B_{*}^{\top}P(\Theta_{*})B_{*}+R)^{-1}B_{*}^{\top}P(\Theta_{*})A_{*}.$ (48)
In (48), $P(\Theta_{*})$ denotes the unique solution for the Discrete Algebraic Riccati Equation (DARE), computed by
$P(\Theta_{*})=Q+A_{*}^{\top}P(\Theta_{*})A_{*}-A_{*}^{\top}P(\Theta_{*})B_{(}B_{*}^{\top}P(\Theta_{*})B_{*}+R)^{-1}B_{*}^{\top}P(\Theta_{*})A_{*}.$ (49)
It is known that the optimal average expected cost is $J_{*}=P(\Theta_{*})\bullet W$.
This problem can be transformed into the Semi-Definite Programming (SDP) formulation in the following manner.
Observing that a steady-state joint state and input distribution $(x_{\infty},u_{\infty})$ exists for any stabilizing policy $\pi$, we denote the covariance matrix of this joint distribution as $\mathcal{E}(\pi)$, which is defined as follows:
$\mathcal{E}(\pi)=\mathbb{E}\begin{pmatrix}x_{\infty}x_{\infty}^{\top}&x_{\infty}u_{\infty}^{\top}\\ u_{\infty}x_{\infty}^{\top}&u_{\infty}u_{\infty}^{\top}\end{pmatrix}$
where $u_{\infty}=\pi(x_{\infty})$ and the average expected cost of the policy is given as follows
$J(\pi)=\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}\bullet\mathcal{E}(\pi).$
Now, for a linear stabilizing policy $K$ that maps $u_{\infty}=Kx_{\infty}$, the covariance matrix of the joint distribution takes the following form:
$\mathcal{E}(K)=\begin{pmatrix}X&XK^{\top}\\ KX&KXK^{\top}\end{pmatrix}$. (50)
where $X=\mathbb{E}[x_{\infty}x_{\infty}^{\top}]$. Then the average expected cost is computed as follows
$J(K)=\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}\bullet\mathcal{E}(K)=(Q+K^{\top}RK)\bullet X.$
Now given $\Theta_{*}=(A_{*}\;B_{*})^{\top}$, the LQR control problem can be formulated in a SDP form, as follows:
$\textrm{minimize}_{\Sigma}\;\;\;\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}\bullet\Sigma$
$\textrm{S.t.}\;\;\;\Sigma_{xx}={\Theta}_{*}^{\top}\Sigma\Theta_{*}+W$
$\Sigma\succ 0$ (51)
where any feasible solution $\Sigma$ can be represented as follows:
$\Sigma=\begin{pmatrix}\Sigma_{xx}&\Sigma_{xu}\\ \Sigma_{ux}&\Sigma_{uu}\end{pmatrix}$ (52)
in which $\Sigma_{xx}\in\mathbb{R}^{n\times n}$, $\Sigma_{uu}\in\mathbb{R}^{m\times m}$, and $\Sigma_{ux}=\Sigma_{xu}\in\mathbb{R}^{m\times n}$. The optimal value of (51) coincides with the average expected cost $J_{*}$, and for $W\succ 0$ (ensuring $\Sigma_{xx}\succ 0$), the optimal policy of the system, denoted as $K_{*}$ and associated with the optimum solution $\Sigma^{*}$, can be derived from $K_{*}=\Sigma^{*}_{{ux}}\Sigma^{*^{-1}}_{{xx}}$ considering (50).
The dual SDP problem is written as follows
[…]
Let $P_{*}$ denote the optimal solution of (56) and let the optimal input be $u_{t}^{*}=K_{*}x_{t}$. Then define
$z_{t}=\begin{pmatrix}I\\ K_{*}\end{pmatrix}x_{t}.$
Multiplying both sides of the optimal equality (56) by $z_{t}^{\top}$ and $z_{t}$ yields
$P_{*}=Q+K_{*}^{\top}RK_{*}+(A_{*}+B_{*}K_{*})^{\top}P_{*}(A_{*}+B_{*}K_{*})$ (57)
which is the prominent optimal Riccati equation. Therefore, the dual solution $P_{*}$ is indeed the solution of the optimal Riccati equation.
B.2 $(\kappa,\gamma)$-stability
To gain a deeper understanding of $(\kappa,\gamma)$-stability, we present illustrative proofs from Lemma B.1 of cohen2018online. The claim to be shown is that for any linear time-invariant system defined by the pair $(A_{*},B_{*})$, a stabilizing policy $K$ ensures $(\kappa,\gamma)$-strong stability in the sense of Definition 1. The closed-loop system achieves stability if $\rho(A_{*}+B_{*}K)=1-\gamma$, for some $0<\gamma<1$. Now defining matrix $Q=(1-\gamma)^{-1}(A_{*}+B_{*}K)$, the stability of $Q$ is guaranteed if a positive definite matrix $P$ exists such that
$Q^{\top}PQ\preceq P.$ (58)
By definition, when $Q$ is stable, then $(A_{*}+B_{*}K)$ is also stable. Inequality (58) can be further rewritten as:
$(A_{*}+B_{*}K)^{\top}P(A_{*}+B_{*}K)\preceq(1-\gamma)^{2}P.$ (59)
Pre and post multiplying both sides of (59) by $P^{-\frac{1}{2}}$ and $P^{\frac{1}{2}}$ yields
$L^{\top}L\preceq(1-\gamma)^{2}.$ (60)
where $L=P^{\frac{1}{2}}(A_{*}+B_{*}K)P^{-\frac{1}{2}}$. Then defining $H=P^{-\frac{1}{2}}$, one can write
$HLH^{-1}=A+BK.$
Setting the condition number of $P^{-\frac{1}{2}}$ to $\kappa$, i.e., $\|H\|\|H^{-1}\|\leq\kappa$ and having $\|L\|\leq 1-\gamma$ from (60) complete proof.
Appendix C Helpful Lemmas
Lemma 1
(Lemma 25 of cohen2019learning) Let $X$ and $Z$ denote matrices of equal size and let $Y$ denote a $(\kappa,\gamma)$ stable matrix such that $X\preceq Y^{\top}XY+Z$, then $X\preceq\frac{\kappa^{2}}{\gamma}\|Z\|I$.
The following lemma, adapted from abbasi2011online gives a self-normalized bound for scalar-valued martingales.
Lemma 2
Let $F_{k}$ be a filtration, $z_{k}$ be a stochastic process adapted to $F_{k}$ and $\omega^{i}_{k}$ (where $\omega^{i}_{k}$ is the $i$-th element of noise vector $\omega_{k}$) be a real-valued martingale difference, again adapted to filtration $F_{k}$ which satisfies the conditionally sub-Gaussianity assumption (Assumption 1) with known constant $\sigma_{\omega}$. Consider the martingale and co-variance matrices:
$S^{i}_{t}:=\sum_{k=1}^{t}z_{k-1}\omega^{i}_{k},\quad V_{t}=\lambda I+\sum_{s=1}^{t-1}z_{t}z_{t}^{\top}$
then with probability of at least $1-\delta$, $0<\delta<1$ we have,
$\left\lVert S^{i}_{t}\right\rVert^{2}_{V_{t}^{-1}}\leq 2\sigma_{\omega}^{2}\log\bigg(\frac{\det(V_{t})}{\delta\det(\lambda I)}\bigg)$
Proof The proof follows from Theorem 3 and Corollary 1 in abbasi2011regret which gives.
$\left\lVert S^{i}_{t}\right\rVert^{2}_{V_{t}^{-1}}\leq 2\sigma_{\omega}^{2}\log\bigg(\frac{\det(V_{t})^{1/2}\det(\lambda I)^{-1/2}}{\delta}\bigg)$
Then, we have
$\log\bigg(\frac{\det(V_{t})}{\det(\lambda I)}\bigg)\geq\log\bigg(\frac{\det(V_{t})^{1/2}}{\det(\lambda I)^{1/2}}\bigg),$
true which completes proof.
The following corollary generalizes the Lemma 3 for a vector-valued martingale which will be later used in Theorem 8 to derive the confidence ellipsoid of estimates.
Corollary 3
Under the same assumptions as in Lemma 3 and defining
$S_{t}=Z^{\top}_{t}W_{t}=\sum_{k=1}^{t}z_{k-1}\omega^{\top}_{k}$
with probability at least $1-\delta$,
$\operatorname{Tr}(S^{\top}_{t}V^{-1}_{t}S_{t})$ $\leq 2\sigma_{\omega}^{2}n\log\bigg(\frac{n\det(V_{t})}{\delta\det(\lambda I)}\bigg)$ (61)
Proof Applying Lemma 3 for i=1,,ni=1,...,n yields
$\left\lVert S^{i}_{t}\right\rVert^{2}_{V_{t}^{-1}}\leq 2\sigma_{\omega}^{2}\log\bigg(\frac{n\det(V_{t})}{\delta\det(\lambda I)}\bigg)$
with probability at least $1-\delta/n$. Furthermore, we have
$\operatorname{Tr}(S^{\top}_{t}V^{-1}_{t}S_{t})=\sum_{i=1}^{n}{S^{i}_{t}}^{\top}V^{-1}_{t}S^{i}_{t}$
applying union bound, (61) holds with probability at least $1-\delta$.
Lemma 3
(cohen2019learning) For some $M\succ 0$ and vector $z$ the following statement holds true.
$\min\{z^{\top}M^{-1}z,1\}\leq 2\log\frac{\det(zz^{\top}+M)}{\det(M)}$
Lemma 4
Let $a\geq 1$ and $b\geq e$. Then the following inequality
$\frac{a\log(bx)}{\sqrt{x+c}}\leq 1$
holds for any $x\geq 1$, provided that
$c=a^{2}\Big(2\log(2a^{2}b)+1\Big)^{2}$
Proof Equivalently we look to specify $c$ such that
$a^{2}\log^{2}bx\leq x+c,\quad\forall x\geq 1.$
Let $y=\log bx$, and define the function
$g(y):=a^{2}y^{2}-\frac{e^{y}}{b},\quad y\geq\log b.$
Then the problem reduces to finding a $c$ such that $g(y)\leq c$ for all $y\geq\log b$.
The critical points of $g(y)$ must satisfy $g^{\prime}(y)=0$ which equivalently is
$\frac{e^{y}}{2a^{2}b}=y.$
Now suppose $y_{*}$ be any solution of this equation in $\log\beta,\;\infty)$. If there was no such $y_{*}$ then the natural choice for $c$ is $\log b$. Now we aim to upper-bound $y_{*}$ if it exists. We let
$D:=\log(2a^{2}b)$
and we define the candidate $Y$ such that $y_{*}\leq Y$ as follows
$Y:=D+\log D+1$
Now we define the following function which is continuous for any $y>0$
$H(y):=\frac{e^{y}}{2a^{2}by}.$
Then we have
$H^{\prime}(y)=\frac{e^{y}(y-1)}{2a^{2}by^{2}}\geq 0$
which shows that it is non-decreasing for any $y\in[1,\;\infty)$. Now lets us evaluate $H(y)$ in $y=Y$.
$H(Y)=\frac{e^{Y}}{2a^{2}\beta Y}=\frac{e\;e^{D}D}{2a^{2}b(D+\log D+1)}=\frac{eD}{D+\log D+1}$ (62)
Now notice that
$\frac{(e-1)D}{\log D+1}\geq 1,\quad\textit{for}\;D\geq\log 2$
which results in
$\frac{eD}{\log D+D+1}=\frac{(e-1)D+D}{\log D+1+D}\geq\frac{(e-1)D}{\log D+1}\geq 1.$
Applying this result in (62) means $H(Y)\geq 1$.
Now, notice that since $H(y_{*})=1$ and $H$ is nondecreasing, it follows that
$y_{*}\leq Y=D+\log D+1.$
Now consider the function $g(y)$ and the fact that it attains its maximum in $y_{*}$. Then for any admissible $y$ we can write
$g(y)\leq g(y_{*})=a^{2}y_{*}^{2}-\frac{e^{y_{*}}}{b}\leq a^{2}y_{*}^{2}\leq a^{2}Y^{2}=a^{2}(D+\log D+1)^{2}\leq a^{2}(2D+1)^{2}.$
Recalling that a $c$ such that $g(y)\leq c$ indeed guarantees $a^{2}\log^{2}bx\leq x+c$ for all $x\geq 1$, then we choose such $c$ as follows
$c=a^{2}\Big(2\log(2a^{2}b)+1\Big)^{2}.$
Lemma 5
(Azuma, 1967 — Theorem 22 in cohen2019learning) Let $X_{1},\dots,X_{N}$ be a martingale‑difference sequence (with respect to some filtration) such that
$|X_{i}|\;\leq\;c\quad\text{for all }i=1,\dots,N.$
Then for all $t>0$,
$\Pr\!\Big(\sum_{i=1}^{N}X_{i}>t\Big)\;\leq\;\exp\!\Big(-\frac{t^{2}}{2Nc^{2}}\Big).$
Lemma 6
(Hanson–Wright inequality, Theorem 21 in cohen2019learning) Let $x\sim\mathcal{N}(0,I_{n})$ be a standard Gaussian random vector, and let $A\in\mathbb{R}^{m\times n}$. Then for all $z>0$,
$\Pr\Big(\|Ax\|-\|A\|_{F}>2\|A\|\,\|A\|_{F}\sqrt{z}+2\|A\|^{2}z\Big)<e^{-z}.$
In particular, if $x\sim\mathcal{N}(0,\Sigma)$ for some covariance matrix $\Sigma$, then $\mathbb{E}\|Ax\|^{2}=\operatorname{Tr}(A\Sigma A^{\top})$, and for any $z\geq 1$,
$\Pr\Big(\|Ax\|^{2}-\mathbb{E}\|Ax\|^{2}>4z\|\Sigma\|\,\|A\|\,\|A\|_{F}\Big)<e^{-z}.$
Appendix D Confidence Ellipsoid Construction
Consider the linear system (5), that evolves up to time $t$. Then, we can express the dynamics as follows:
$X_{t}$ $=Z_{t}\Theta_{*}+W_{t}$
where $W$ is the vertical concatenation of $\omega_{1}^{\top},...,\omega_{t}^{\top}$ and $X_{t}$ and $Z_{t}$ are matrices constructed by rows $x^{\top}_{1},...,x^{\top}_{t}$ and ${z}^{\top}_{1},...,{z}^{\top}_{t-1}$ respectively.
Given an initial estimate $\Theta_{0}$ of $\Theta_{*}$ we define $e(\Theta)$ as follows:
$e(\Theta)$ $=\lambda\operatorname{Tr}\big((\Theta-\Theta_{0})^{\top}(\Theta-\Theta_{0})\big)+\sum_{s=0}^{t-1}\operatorname{Tr}\big((x_{s+1}-{\Theta}^{\top}z_{s})(x_{s+1}-{\Theta}^{\top}z_{s})^{\top}\big)$
which has a regularization term $\operatorname{Tr}\big((\Theta-\Theta_{0})^{\top}(\Theta-\Theta_{0})\big)$ that penalizes deviation from the estimate $\Theta_{0}$. Then $l_{2}$-regularized least square estimation is obtained as follows:
$\hat{\Theta}_{t}=\operatorname*{argmin}_{\Theta}e(\Theta)={V}_{t}^{-1}\big(Z_{t}^{\top}X_{t}+\lambda\Theta_{0}\big)$ (63)
where
$V_{t}=\lambda I+\sum_{s=0}^{t-1}z_{s}z_{s}^{\top}=\lambda I+Z_{t}^{\top}Z_{t},$
In a different scenario when there is no such an initial estimate $\Theta_{0}$, the regularization is done with respect to $\operatorname{Tr}\big(\Theta^{\top}\Theta\big)$ that results in least square estimation
$\hat{\Theta}_{t}={V}_{t}^{-1}Z_{t}^{\top}X_{t}.$ (64)
The following theorem provides an upper-bound on estimation error for both of the mentioned scenarios.
Theorem 8
(abbasi2011regret,cohen2019learning) The least square estimation error $(\hat{\Theta}_{t}-\Theta_{*})$ satisfies
$\operatorname{Tr}\big((\hat{\Theta}_{t}-\Theta_{*})^{\top}V_{t}(\hat{\Theta}_{t}-\Theta_{*})\big)\leq r_{t}$ (65)
where
  1. 1.
    when there is an initial estimate $\Theta_{0}$ such that $\|\Theta_{0}-\Theta_{*}\|_{*}\leq\epsilon$
    $r_{t}=\bigg(\sigma_{\omega}\sqrt{2n\log\frac{n\det(V_{t})}{\delta\det(\lambda I)}}+\sqrt{\lambda}\epsilon\bigg)^{2}$ (66)
    and $\hat{\Theta}_{t}$ is computed by (63).
  2. 2.
    And when there is no such $\Theta_{0}$
    $r_{t}=\bigg(\sigma_{\omega}\sqrt{2n\log\frac{n\det(V_{t})}{\delta\det(\lambda I)}}+\sqrt{\lambda}\vartheta\bigg)^{2}.$ (67)
    and $\hat{\Theta}_{t}$ is computed by (64).
Appendix E Relaxed Primal and Dual SDP
When incorporating the confidence ellipsoid, the primal and dual SDPs (51 and 56) can be relaxed to be used for control design and regret bound analysis. In order to incorporate the estimation error in SDP formulation, we first need the following perturbation lemma. This lemma, a generalized version of Lemma 24 of cohen2019learning, can be used to incorporate un-normalized confidence ellipsoids into SDP formulation.
Lemma 7 (Perturbation bound with trace-based confidence ellipsoid)
Let $X$, $\Delta$ denotes matrices of matching size, and let $P\succeq 0$, and $V\succ 0$ be symmetric matrices. Suppose the trace-based confidence ellipsoid satisfies
$\operatorname{Tr}(\Delta V\Delta^{\top})\leq r$
Then we have the following symmetric positive semidefinite bound:
$\mu\,\|P\|V^{-1}\preceq(X+\Delta)^{\top}P(X+\Delta)-X^{\top}PX\preceq\mu\,\|P\|V^{-1},$
where
$\mu:=2\,\|V\|^{1/2}\,\|X\|\,\sqrt{r}+r.$
Proof Expanding the quadratic difference gives
$(X+\Delta)^{\top}P(X+\Delta)-X^{\top}PX=X^{\top}P\Delta+\Delta^{\top}PX+\Delta^{\top}P\Delta.$
Define
$M:=V^{1/2}(X^{\top}P\Delta+\Delta^{\top}PX)V^{1/2}$
which is a positive semi definite matrix.
Using the triangle inequality:
$\|M\|\leq\|V^{1/2}X^{\top}P\Delta V^{1/2}\|+\|V^{1/2}\Delta^{\top}PXV^{1/2}\|.$
Since $\|A^{\top}\|=\|A\|$ for any matrix $A$, the two terms are equal as P is PSD, giving
$\|M\|\leq 2\,\|V^{1/2}X^{\top}P\Delta V^{1/2}\|.$
For any matrices $A,B,C$,
$\|ABC\|\leq\|A\|\,\|B\|\,\|C\|_{F}.$
Set $A=V^{1/2}X^{\top}$, $B=P$, $C=\Delta V^{1/2}$. Then
$\|V^{1/2}X^{\top}P\Delta V^{1/2}\|\leq\|V^{1/2}X^{\top}\|\,\|P\|\,\|\Delta V^{1/2}\|_{F}.$
Hence,
$\|M\|\leq 2\,\|V^{1/2}X^{\top}\|\,\|P\|\,\|\Delta V^{1/2}\|_{F}.$
By the definition of the Frobenius norm,
$\|\Delta V^{1/2}\|_{F}^{2}=\operatorname{Tr}((\Delta V^{1/2})(\Delta V^{1/2})^{\top})=\operatorname{Tr}(\Delta V\Delta^{\top})\leq r.$
Thus,
$\|M\|\leq 2\,\|V^{1/2}\|\|X\|\,\|P\|\,\sqrt{r}\leq 2\|V\|^{1/2}\|X\|\,\|P\|\,\sqrt{r}.$
in which we applied the fact that $\|V^{1/2}\|=\|V\|^{1/2}$.
Since
$X^{\top}P\Delta+\Delta^{\top}PX=V^{-1/2}MV^{-1/2},$
and for any symmetric matrix $M$,
$V^{-1/2}MV^{-1/2}\preceq\|M\|\,V^{-1},$
we obtain
$(X^{\top}P\Delta+\Delta^{\top}PX)\preceq 2\|V\|^{1/2}\|X\|\,\|P\|\,\sqrt{r}V^{-1}.$ (68)
We now upper-bound the second term $\Delta^{\top}P\Delta$. Define
$\bar{M}=(V^{1/2}\Delta^{\top})\,P\,(\Delta V^{1/2}).$
It follows that
$\|\bar{M}\|\leq\|P\|\,\|\Delta V^{1/2}\|_{F}^{2}=\|P\|\,\operatorname{Tr}(\Delta V\Delta^{\top})\leq\|P\|\,r.$
Therefore,
$\Delta^{\top}P\Delta$ $=V^{-1/2}\bar{M}V^{-1/2}\preceq\|\bar{M}\|V^{-1}$
$\preceq\|P\|\,rV^{-1}.$ (69)
Combining the bounds (68) and (69) gives
$(X+\Delta)^{\top}P(X+\Delta)-X^{\top}PX\preceq\mu\,\|P\|V^{-1}.$
where $\mu:=2\,\|V\|^{1/2}\,\|X\|\,\sqrt{r}+r$.
The other side of inequality can be easily shown by following a similar argument.
Now having the perturbation lemma we can write the following lemma to relax the SDPs (51 and 56).
Lemma 8
Let $\Theta$ be an estimate of $\Theta_{*}$ such that $\operatorname{Tr}\big((\Theta-\Theta_{*})^{\top}V(\Theta-\Theta_{*})\big)\leq r$ for some $V\succ 0$ and $r>0$. Then the the primal and dual SDPs’ equalities (51) and (56) hold if the following inequalities hold respectively.
$\Sigma_{xx}\succeq\Theta^{\top}\Sigma\Theta+W-\mu(\Sigma\bullet V^{-1})I$ (70)
and
$\begin{pmatrix}Q-P&0\\ 0&R\end{pmatrix}+\Theta P{\Theta}^{\top}\succeq\mu\|P\|V^{-1}$ (71)
where $\mu\leq r+2\vartheta\sqrt{r}\|V\|^{1/2}$.
Proof Noting that $\operatorname{Tr}\big((\Theta-\Theta_{*})^{\top}V(\Theta-\Theta_{*})\big)\leq r$ implies $(\Theta-\Theta_{*})^{\top}V(\Theta-\Theta_{*})\big)\preceq rV^{-1}$, the proof follows by applying perturbation lemma, Lemma 7.
To show (70), consider (51), and then apply Lemma 7 by treating $\Theta$ as $X$ and $\Theta_{*}-\Theta$ as $\Delta$, i.e.,
$\Sigma_{xx}$ $={\Theta_{*}}^{\top}\Sigma\Theta_{*}+W$
$=({\Theta_{*}-\Theta+\Theta})^{\top}\Sigma({\Theta_{*}-\Theta+\Theta})+W$
$\succeq\Theta^{\top}\Sigma\Theta-\mu\|\Sigma\|_{*}V^{-1}+W$ (72)
$\succeq\Theta^{\top}\Sigma\Theta-\mu\|\Sigma\|_{*}\|V^{-1}\|_{*}I+W$ (73)
$=\Theta^{\top}\Sigma\Theta-\mu\|\Sigma V^{-1}\|_{*}I+W$ (74)
$=\Theta^{\top}\Sigma\Theta+W-\mu(\Sigma\bullet V^{-1})I$ (75)
where in (72) we applied Lemma 7, inequality (73) holds due to the fact that $V^{-1}\preceq\|V^{-1}\|_{*}I$ as $V^{-1}\succ 0$. The equality (74) is due to the fact that tow matrices $\|\Sigma\|_{*}I$ and $\|V^{-1}\|_{*}I$ are permutation matrices, and for any two permutation matrices $A$ and $B$, i.e., $AB=BA$ we have $\|AB\|_{*}=\|A\|_{*}\|B\|_{*}$. Finally, the equality (75) holds by definition.
Proof of (71) follows similar procedure.
Conclusively, given the parameter estimate in the form of confidence ellipsoid (7, 8) by this construction the relaxed primal SDP is formulated as follows:
[…]
where $\mu_{t}=r_{t}+\sqrt{r_{t}}\vartheta\|V_{t}\|^{1/2}$. Denoting optimal solution of program (79) by $\Sigma(\mathcal{C}_{t})$, the control signal extracted from solving relaxed primal SDP, $u=K(\mathcal{C}_{t})x$ is deterministic and linear in state where
$K(\mathcal{C}_{t})=\Sigma_{ux}(\mathcal{C}_{t}){\Sigma_{xx}^{-1}(\mathcal{C}_{t})}.$
Furthermore, the relaxed dual SDP formulation is given as follows:
[…]
where we denote its optimal solution by $P(\mathcal{C}_{t})$.
Appendix F Lower-bound on minimum eigenvalue of covariance matrix
In this section, we aim to provide a lower bound for the minimum eigenvalue of covariance matrix $V_{t}\succ 0$ during implementation of Algorithm 1, the so-called ARSLO. This bound is highly useful for stability guarantee and regret bound analysis. We build upon the methodology of Theorem 20 in cohen2019learning, which is specifically designed for the warm-up phase with a fixed policy $K_{0}$ and input perturbation with a time-invariant covariance matrix. However, what we obtain is tailored to the ARSLO algorithm, which incorporates input perturbation with a time-varying Gaussian noise covariance and does not rely on a fixed policy.
We begin with the following lemma which is borrowed from cohen2019learning.
Lemma 9
[…]
$\eta_{t}\sim\mathcal{N}\left(0,\frac{\sigma_{\omega}^{2}\bar{p}_{t}\left(K_{t}K_{t}^{\top}+\frac{\|P_{t}\|}{\alpha_{0}}I\right)}{\sqrt{t+\bar{c}}}\right),$
where $\bar{p}_{t}=\mathcal{O}(\log t)$ and $P_{t}$ (with some abuse of notation) is the solution of the relaxed dual SDP (17). Then we have $\mathbb{E}[z_{t}z_{t}^{\top}]\succeq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{2\sqrt{t}}I$ for all $t$’s such that
$\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}-\frac{1}{2}\geq 0.$
Proof Given the dynamics and the assumption on the process noise $\omega_{t}$ it is straightforward to show that $\mathbb{E}[x_{t}x_{t}^{\top}|\mathcal{F}_{t-1}]\succeq\sigma_{\omega}^{2}I$. Furthermore by definition $u_{t}=K_{t}x_{t}+\eta_{t}$ and considering the fact that input perturbation $\eta_{t}$ is independent than $\omega_{t+1}$ we can write:
$\mathbb{E}[z_{t}z_{t}^{\top}|\mathcal{F}_{t-1}]=\begin{pmatrix}I\\ K_{t}\end{pmatrix}\mathbb{E}[x_{t}x_{t}^{\top}|\mathcal{F}_{t-1}]\begin{pmatrix}I\\ K_{t}\end{pmatrix}^{\top}+\begin{pmatrix}0&0\\ 0&\mathbb{E}[\eta_{t}\eta_{t}^{\top}|\mathcal{F}_{t-1}]\end{pmatrix}$.
By lower-bounding right hand side we can further write
$\mathbb{E}[z_{t}z_{t}^{\top}|\mathcal{F}_{t-1}]\succeq$ $\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{\sqrt{t+\bar{c}}}\begin{pmatrix}\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}I&\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}\\ \frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}^{\top}&\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}K_{t}^{\top}+K_{t}K_{t}^{\top}+\frac{\|P_{t}\|}{\alpha_{0}}\end{pmatrix}$
$\succeq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{\sqrt{t+\bar{c}}}\begin{pmatrix}(\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}-\frac{1}{2}+\frac{1}{2})I&\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}\\ \frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}^{\top}&\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}K_{t}^{\top}+K_{t}K_{t}^{\top}+\frac{I}{2}\end{pmatrix}$
$\succeq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{\sqrt{t+\bar{c}}}\begin{pmatrix}\frac{I}{2}&0\\ 0&\frac{I}{2}\end{pmatrix}+\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{\sqrt{t+\bar{c}}}\underbrace{\begin{pmatrix}(\overbrace{\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}-\frac{1}{2})I}^{\bar{\Gamma}_{11}}&\overbrace{\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}}^{\bar{\Gamma}_{12}}\\ \underbrace{\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}^{\top}}_{\bar{\Gamma}_{21}}&\underbrace{\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}K_{t}K_{t}^{\top}+K_{t}K_{t}^{\top}}_{\bar{\Gamma}_{22}}\end{pmatrix}}_{\bar{\Gamma}}$
$\succeq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{2\sqrt{t+\bar{c}}}I\quad\quad\forall t\quad\textit{such that}\quad\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}-\frac{1}{2}\geq 0,$
where in the second inequality we used the fact that $\|P_{t}\|/\alpha_{0}\geq 1/2$ (see (119) in the proof of Lemma 14). The last inequality holds because $\bar{\Gamma}\succ 0$ by Schur complement noting that $\bar{\Gamma}_{11}\succ 0$ and $\bar{\Gamma}_{22}-\bar{\Gamma}_{12}\bar{\Gamma}_{11}^{-1}\bar{\Gamma}_{21}\succ 0$.
Remark 3
[…]
$B_{*}\frac{\bar{p}_{t}\,\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|I}{\alpha_{0}}\Big)}{\sqrt{t+\bar{c}}}B_{*}^{\top}\preceq\sigma_{\omega}^{2}I.$
[…]
$2\kappa_{t}^{2}\bar{\vartheta}_{B_{*}}^{2}\frac{\bar{p}_{t}}{\sqrt{t+\bar{c}}}\leq 1,$
which yields
$\frac{\bar{p}_{t}}{\sqrt{t+\bar{c}}}\leq 1\quad\text{as }\kappa_{t}\geq 1.$ (84)
This ensures the existence of a small value $s$ such that
$\frac{\sqrt{t+\bar{c}}}{\bar{p}_{t}}-\frac{1}{2}\geq 0\quad\text{for all }t\geq s,$
and consequently, the minimum time step for which the result of Lemma 9 holds is small.
Lemma 10
(Lemma 35 cohen2019learning) Let $a\in\mathbb{R}^{n+m}$ be a unit vector, and define $S_{t}=a^{\top}z_{t}$. Consider the random indicator variable $I_{t}$ defined by
$I_{t}=\begin{cases}1&\text{if }S_{t}^{2}>\dfrac{\sigma_{\omega}^{2}\bar{p}_{t}}{4\sqrt{t+\bar{c}}},\\ 0&\text{otherwise}.\end{cases}$
Then, it holds that $\mathbb{E}[I_{t}\mid\mathcal{F}_{t-1}]\geq\dfrac{1}{5}$.
Now, we are in a position to introduce the key lemma which will be useful for stability analysis purpose.
Lemma 11
Let $a\in\mathbb{R}^{n+m}$ be a unit vector, then
$a^{\top}\big(\sum_{k=1}^{t}z_{k}z_{k}^{\top}\big)a\geq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{40}\sqrt{t+\bar{c}}-{\frac{\sigma_{\omega}^{2}\bar{c}}{40}}$ (85)
for $t\geq 200\log 1/\delta$ with probability at least $1-\delta$ where $\delta\in(0,\;1)$.
Proof By defining $\mathcal{M}_{t}=I_{t}-\mathbb{E}[I_{t}|\;\mathcal{F}_{t-1}]$ and noting that it is a martingale difference and the fact that $|\mathcal{M}_{t}|\leq 1$, we can apply Azuma inequality as follows:
$\sum_{k=1}^{t}\mathcal{M}_{k}\geq-\sqrt{-2t\log\delta}\geq-\frac{t}{10}$
where the last inequality holds for $t\geq-200\log\delta$. By definition of $\mathcal{M}_{k}$ this implies
$\sum_{k=1}^{t}I_{k}\geq\sum_{k=1}^{t}\mathcal{M}_{k}+\sum_{k=1}^{t}\mathbb{E}[I_{k}|\mathcal{F}_{k-1}]\geq-\frac{t}{10}+\sum_{k=1}^{t}\mathbb{E}[\mathcal{M}_{k}|\mathcal{F}_{k-1}]\geq-\frac{t}{10}+\frac{t}{5}\geq\frac{t}{10}$
where the last inequality holds by Lemma 10. We further have $S_{k}^{2}\geq I_{k}\frac{\sigma_{\omega}^{2}\bar{p}_{k}}{4\sqrt{k+\bar{c}}}$ which results in
$a^{\top}\big(\sum_{k=1}^{t}z_{k}z_{k}^{\top}\big)a$ $=\sum_{k=1}^{t}S_{k}^{2}\geq\sum_{k=1}^{t}I_{k}\frac{\sigma_{\omega}^{2}\bar{p}_{k}}{4\sqrt{k+\bar{c}}}\geq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{4\sqrt{t+\bar{c}}}\sum_{k=1}^{t}I_{k}$
$\geq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{40\sqrt{t+\bar{c}}}t=\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{40}\sqrt{t+\bar{c}}-\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{40\sqrt{t+\bar{c}}}\bar{c}$
$\geq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{40}\sqrt{t+\bar{c}}-\frac{\sigma_{\omega}^{2}\bar{c}}{40}$
where in the last inequality we applied (84). This completes the proof.
Now we use the result of Lemma 11, to find a lower-bound on the minimum eigenvalue of co-variance matrix $\sum_{k=1}^{t}z_{k}z_{k}^{\top}$. The following Lemma concludes the result.
Lemma 12
Fix $\delta\in(0,\,1)$. Then with probability at least $1-\delta$ we have
$\lambda_{min}\big(\sum_{k=1}^{t}z_{k}z_{k}^{\top}\big)\geq\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\underbrace{\frac{\sigma_{\omega}^{2}\bar{c}}{80}}_{=:\bar{C}}$ (86)
for $t\geq 400(n+m+\log 1/\delta)$.
Proof We establish the result using a standard covering argument over the unit sphere, following the general strategy of cohen2019learning, but we include it here for completeness.
Let $\mathcal{N}(1/4)$ be a minimal $1/4$-net of the unit sphere $\mathbb{S}^{n+m-1}$, and define
$\bar{V}_{t}:=\sum_{k=1}^{t}z_{k}z_{k}^{\top}.$
Consider the set
$\mathcal{M}\;=\;\left\{\frac{\bar{V}_{t}^{-1/2}u}{\|\bar{V}_{t}^{-1/2}u\|}\;:\;u\in\mathcal{N}(1/4)\right\}.$
Assume that $t\;\geq\;200\log\!\left({|\mathcal{M}|}/{\delta}\right)$. Using (85) and applying a union bound over all $n\in\mathcal{M}$, we obtain that with probability at least $1-\delta$,
$n^{\top}\bar{V}_{t}n\;\geq\;\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{40}\sqrt{t+\bar{c}}-\frac{\sigma_{\omega}^{2}\bar{c}}{40}\;=:\;\mathcal{A}_{t}.$
By construction, this implies that for every $u\in\mathcal{N}(1/4)$,
$u^{\top}\bar{V}_{t}^{-1}u\;\leq\;\frac{1}{\mathcal{A}_{t}}.$ (87)
Let $z$ be a unit eigenvector corresponding to the smallest eigenvalue of $\bar{V}_{t}$. By the definition of the $1/4$-net, there exists $u_{z}\in\mathcal{N}(1/4)$ such that $\|z-u_{z}\|\leq 1/4$. We then have
$\|\bar{V}_{t}^{-1}\|=z^{\top}\bar{V}_{t}^{-1}z$ $\leq u_{z}^{\top}\bar{V}_{t}^{-1}u_{z}+(z-u_{z})^{\top}\bar{V}_{t}^{-1}(z+u_{z})$
$\leq u_{z}^{\top}\bar{V}_{t}^{-1}u_{z}+\|z-u_{z}\|\,\|\bar{V}_{t}^{-1}\|\,\|z+u_{z}\|$
$\leq\frac{1}{\mathcal{A}_{t}}+\frac{2}{4}\|\bar{V}_{t}^{-1}\|,$
where we used (87) and the fact that $\|z\|=\|u_{z}\|=1$.
Rearranging yields
$\|\bar{V}_{t}^{-1}\|\;\leq\;\frac{2}{\mathcal{A}_{t}}.$
Finally, note that $|\mathcal{M}|=|\mathcal{N}(1/4)|$, and standard bounds give $|\mathcal{N}(1/4)|\leq 12^{n+m}$. Hence,
$t\;\geq\;400\bigl(n+m+\log(1/\delta)\bigr)$
ensures the required condition on $t$. Since $\|\bar{V}_{t}^{-1}\|=1/\lambda_{\min}(\bar{V}_{t})$, the claim follows.
Appendix G Stability Analysis of ARSLO algorithm
To establish stability, we first need the following lemma, which is equivalent to Lemma 16 of cohen2019learning, with the only difference being that we do not require the selection of a regularization parameter $\lambda$ of $\mathcal{O}(\sqrt{T})$. This distinction is indeed the core objective of our study.
Lemma 13
Suppose $\Sigma(\mathcal{C}_{t})$ and $P(\mathcal{C}_{t})$ be the solution of the relaxed primal and dual SDPs (12) and (83) respectively. Having $\mu_{t}\geq r_{t}+2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}$ then $\Sigma_{xx}(\mathcal{C}_{t})$ is invertible and for $K(\mathcal{C}_{t})=\Sigma_{ux}(\mathcal{C}_{t})\Sigma_{xx}(\mathcal{C}_{t})^{-1}$ we have
$P(\mathcal{C}_{t})=$ $\;Q+K^{\top}(\mathcal{C}_{t})RK(\mathcal{C}_{t})+(\hat{A}_{t}+\hat{B}_{t}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})(\hat{A}_{t}+\hat{B}_{t}K(\mathcal{C}_{t}))$
$-\mu_{t}\|P(\mathcal{C}_{t})\|_{*}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}$ (88)
Proof The proof follows the same steps as for Lemma 16 of (cohen2019learning) with only difference that we need to show that
$\mu_{t}\big(\Sigma(\mathcal{C}_{t})\bullet V_{t}^{-1}\big)I\preceq\sigma_{\omega}^{2}I$ (89)
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V^{-1}_{t}\preceq\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}$ (90)
using which all the claims follows. By appropriately tuning the parameters $\lambda$, $\bar{p}_{t}$, and $\bar{\epsilon}$ according to Theorem 1 it is shown that
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I,$ (91)
which together with the fact that $Q,R\preceq\alpha_{0}$ (90) holds true. On the other hand, noting that $\|\Sigma(\mathcal{C}_{t})\|_{*}\preceq\nu/\alpha_{0}$ and (91)
$\mu_{t}\big(\Sigma(\mathcal{C}_{t})\bullet V_{t}^{-1}\big)I$ $=\operatorname{Tr}\big(\frac{1}{\|P(\mathcal{C}_{t})\|}\Sigma(\mathcal{C}_{t})^{\top}\mu_{t}\|P(\mathcal{C}_{t})\|V_{t}^{-1}\big)I$
$\preceq\operatorname{Tr}\big(\frac{1}{\|P(\mathcal{C}_{t})\|}\Sigma(\mathcal{C}_{t})^{\top}\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}\big)I$
$=\frac{1}{\|P(\mathcal{C}_{t})\|}\Sigma(\mathcal{C}_{t})\bullet\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}$
$=\frac{1}{\|P(\mathcal{C}_{t})\|}P(\mathcal{C}_{t})\bullet W$
$\preceq\frac{1}{\|P(\mathcal{C}_{t})\|}\|P(\mathcal{C}_{t})\|\sigma_{\omega}^{2}I=\sigma_{\omega}^{2}I$
that completes the proof for (89).
Now we are in a position to provide the proof for Theorem 1.
G.1 Proof of Theorem 1
The proof of the theorem consists of two steps. First, we establish strong stability using Lemma 14. Then, we prove the sequential property of the generated policies in Lemma 16.
Lemma 14
Any feedback gain $K(\mathcal{C}_{t})$ designed by the ARSLO algorithm is $(\kappa_{*},\gamma_{*})$-strongly stabilizing, where $\kappa_{*}=\sqrt{\frac{2\|P_{*}\|}{\alpha_{0}}}$ and $\gamma_{*}=\kappa_{*}^{-2}/2$.
Proof One can write
$(\hat{A}_{t}+\hat{B}_{t}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})(\hat{A}_{t}+\hat{B}_{t}K(\mathcal{C}_{t}))$ $=\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}\hat{\Theta}_{t}P(\mathcal{C}_{t})\hat{\Theta}_{t}^{\top}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}$
$\preceq\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}{\Theta}_{*}P(\mathcal{C}_{t}){\Theta}_{*}^{\top}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}$ (92)
$-\mu_{t}\|P(\mathcal{C}_{t})\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V_{t}^{-1}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}$
$=({A}_{*}+{B}_{*}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})({A}_{*}+{B}_{*}K(\mathcal{C}_{t}))$
$-\mu_{t}\|P(\mathcal{C}_{t})\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V_{t}^{-1}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}$ (93)
where in the inequality (92) we applied the perturbation lemma, Lemma 7. Substituting (93) into (88), yields
$P(\mathcal{C}_{t})\succeq$ $\;Q+K^{\top}(\mathcal{C}_{t}))RK(\mathcal{C}_{t}))+({A}_{*}+{B}_{*}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})({A}_{*}+{B}_{*}K(\mathcal{C}_{t}))$
$-2\mu_{t}\|P(\mathcal{C}_{t})\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}$. (94)
Now we aim to tune parameters $\lambda$ and $\bar{p}_{t}$ and specify $\bar{\epsilon}$, the error upper-bound of $\Theta_{0}$ to be provided by the warm-up algorithm, such that
$\mu_{t}\|P(\mathcal{C}_{t})\|V_{t}^{-1}\preceq\frac{\alpha_{0}}{32\kappa_{t}^{8}}I.$ (95)
where
$\kappa_{t}=\sqrt{\frac{2\|P(\mathcal{C}_{t})\|}{\alpha_{0}}}.$ (96)
Note that the subscript $t$ in $\kappa_{t}$ corresponds to the subscript $t$ in $\mathcal{C}_{t}$, which denotes the confidence ellipsoid constructed using data up to time $t$.
Applying (96), the condition (95) can be rewritten as follows
$\mu_{t}V_{t}^{-1}\preceq\frac{1}{16\kappa_{t}^{10}}I.$ (97)
By definition $\mu_{t}=r_{t}+2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}$ the sufficient condition for (97) is
$r_{t}V_{t}^{-1}\preceq\frac{1}{32\kappa_{t}^{10}}I$ (98)
$2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}V_{t}^{-1}\preceq\frac{1}{32\kappa_{t}^{10}}I.$ (99)
By applying the result of Lemma 12, namely the inequality (86) one can write:
$2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}V_{t}^{-1}$ $\preceq\frac{2\vartheta\sqrt{r_{t}}\|V_{t}\|^{\frac{1}{2}}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}I$
where
$\bar{C}:=\frac{\sigma_{\omega}^{2}\bar{c}}{80}.$
Then the sufficient condition for (99) is
$\frac{2\vartheta\sqrt{r_{t}}\|V_{t}\|^{\frac{1}{2}}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}$ $\leq\frac{1}{32\kappa_{t}^{10}}$ (100)
and similarly for (98) is
$\frac{r_{t}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}\leq\frac{1}{32\kappa_{t}^{10}}$ (101)
Let further set
$\lambda\geq\bar{C}$ (102)
then the condition for (100) can be written as follows
$\frac{2\vartheta\sqrt{r_{t}}\|V_{t}\|^{\frac{1}{2}}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}$ $\leq\frac{2\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}}\leq\frac{1}{32\kappa_{t}^{10}}$
To satisfy the inequality above, it suffices to require
$\bar{p}_{t}\geq\frac{5120\,\kappa_{t}^{10}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$
Similarly, a sufficient condition for (101) is given by
$\bar{p}_{t}\geq\frac{2560\,r_{t}\,\kappa_{t}^{10}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$
We then choose $\bar{p}_{t}$ as
$\bar{p}_{t}\geq\max\bigg\{\frac{5120\,\kappa_{t}^{10}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}},\,\frac{2560\,r_{t}\,\kappa_{t}^{10}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}\bigg\}.$ (103)
Note that, one can show that $\bar{p}_{t}$ is determined by the first term for all $t$.
Next, we proceed to specify $\bar{c}$ such that
$B_{*}\frac{\bar{p}_{t}\sigma_{\omega}^{2}\big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|I}{\alpha_{0}}\big)}{\sqrt{t+\bar{c}}}B_{*}^{\top}\preceq\sigma_{\omega}^{2}I.$ (104)
This condition ensures that the effect of input perturbation noise on the closed-loop system does not exceed the magnitude of the process noise.
A sufficient condition for (104) can be written as
$\frac{2\,\kappa_{t}^{2}\,\bar{\vartheta}_{B_{*}}^{2}\bar{p}_{t}}{\sqrt{t+\bar{c}}}\leq 1$ (105)
where we define $\bar{\vartheta}_{B_{*}}=\max\{1,\vartheta_{B_{*}}\}$ and $\|B_{*}\|\leq\vartheta_{B_{*}}$.
To derive a suitable upper bound for $\bar{p}_{t}$, we first establish the following inequalities.
We have
$\log\frac{n\det(V_{t})}{\delta\det(\lambda I)}$ $\leq(n+m)\log\frac{n}{\delta}\Big(1+t\log\frac{t}{\delta}\Big)$ (106)
$\leq 2(n+m)\log\frac{2nt}{\delta}.$ (107)
where (106) holds under the choice
$\lambda\geq\frac{\sum_{k=1}^{z}\|z_{k}z_{k}^{\top}\|^{2}}{t\log\frac{2t}{\delta}}.$ (108)
By choosing $\epsilon$ appropriately such that
$\sqrt{\lambda}\epsilon\leq\sigma_{\omega}\sqrt{2n(n+m)}$ (109)
it then follows from (107) and (109) that
$r_{t}\leq 16\sigma_{\omega}^{2}n(n+m)\log\frac{2nt}{\delta}.$ (110)
Furthermore, from Lemma 17, we have
$\|z_{t}\|^{2}\leq c^{2}_{z}(\kappa_{*})\log\frac{t}{\delta}.$
To obtain a computable value for $\bar{c}$, we require an upper bound on $\kappa_{*}$, which can be derived directly from Lemma 15:
$\kappa_{t}^{2}\leq\kappa_{*}^{2}\leq\kappa_{t}^{2}+\frac{1}{2\kappa_{t}^{2}}:=\bar{\kappa}_{t}^{2},\quad\forall t.$
From this, we can define
$\kappa_{t}\leq\kappa_{*}\leq\bar{\kappa}_{1}:=\sqrt{\kappa_{1}^{2}+\frac{1}{\kappa_{1}^{2}}}.$
Combining these results, we can now derive and upper-bound for the left hand side of (105) as follows
$\frac{40960\,\bar{\vartheta}^{2}_{B_{*}}\kappa_{t}^{10}\vartheta\sqrt{r_{t}}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}\big(\lambda+c_{z}^{2}(\kappa_{*})\log\frac{t}{\delta}\big)^{1/2}\leq$ $\frac{40960\,\bar{\vartheta}^{2}_{B_{*}}\bar{\kappa}_{1}^{10}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)\log\frac{t}{\delta}}}{\sigma_{\omega}^{2}{t+\bar{c}}}\sqrt{\lambda}$ (111)
$+\frac{40960\,\bar{\vartheta}^{2}_{B_{*}}\bar{\kappa}_{1}^{10}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}}{\sigma_{\omega}^{2}{t+\bar{c}}}c_{z}(\bar{\kappa}_{1})\log\frac{2nt}{\delta}$ (112)
We enforce the following condition:
$\frac{40960\,\bar{\vartheta}^{2}_{B_{*}}\bar{\kappa}_{1}^{10}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}c_{z}(\bar{\kappa}_{1})\log\frac{2nt}{\delta}\leq\frac{1}{2}.$ (113)
Applying Lemma 4, an appropriated choice of $\bar{c}$ that satisfies this condition is
$\bar{c}=a^{2}\Big(2\log\Big(\frac{4a^{2}n}{\delta}\Big)+1\Big)^{2}.$ (114)
where
$a=\frac{40960\,\bar{\vartheta}^{2}_{B_{*}}\bar{\kappa}_{1}^{10}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}c_{z}(\bar{\kappa}_{1})}{\sigma_{\omega}^{2}}.$
With this choice of $\bar{c}$, condition (113) ensures that (112) is satisfied since $t+\bar{c}>1$. Moreover, by choosing $\lambda=\bar{C}=\frac{\sigma_{\omega}^{2}\bar{c}}{80}$ we also satisfy (111).
Recalling (108), which provides an additional condition for tuning $\lambda$ based on the upper bound of the co-state, we require
$\lambda\geq c_{z}^{2}(\bar{\kappa}_{1}).$
Therefore, we select
$\lambda=\max\bigg\{\frac{\sigma_{\omega}^{2}\bar{c}}{80},\,c_{z}^{2}(\bar{\kappa}_{1})\bigg\}=\frac{\sigma_{\omega}^{2}\bar{c}}{80}$ (115)
where the equality holds due to the magnitude of $\bar{c}$ in (114).
With $\bar{c}$ specified, $\lambda$ is determined, then, from (103), $\bar{p}_{t}$ is obtained as
$\bar{p}_{t}:=\frac{5120\,\kappa_{t}^{10}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$ (116)
Recalling (109), the appropriate choice of $\bar{\epsilon}(\bar{\kappa}_{1})$ is given by
$\bar{\epsilon}(\bar{\kappa}_{1}):=\frac{\sigma_{\omega}\sqrt{2n(n+m)}}{\sqrt{\lambda}}.$ (117)
With $\bar{\epsilon}(\bar{\kappa}_{1})$ given by (117) and $\lambda$ and $\bar{p}_{t}$ tuned according to (115) and (116), respectively, (97) holds true, resulting in the satisfaction of
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I.$ (118)
Applying this into (94) yields
$P(\mathcal{C}_{t})$ $\succeq\frac{\alpha_{0}}{2}I+\frac{\alpha_{0}}{2}K^{\top}(\mathcal{C}_{t})K(\mathcal{C}_{t})+(A_{*}+B_{*}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})(A_{*}+B_{*}K(\mathcal{C}_{t}))$ (119)
where we applied $Q,R\succeq\alpha_{0}I$. Furthermore, (119) implies
$(A_{*}+B_{*}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})(A_{*}+B_{*}K(\mathcal{C}_{t}))\preceq P(\mathcal{C}_{t})-\frac{\alpha_{0}}{2}I.$ (120)
By pre and post multiplying both sides of (120) with $P^{-1/2}(\hat{\Theta}_{t})$ we can further write:
${P}^{-1/2}(\mathcal{C}_{t})(A_{*}+B_{*}K(\mathcal{C}_{t}))^{\top}P(\mathcal{C}_{t})(A_{*}+B_{*}K(\mathcal{C}_{t})){P}^{-1/2}(\mathcal{C}_{t})$ $\preceq I-\frac{\alpha_{0}}{2}{P}^{-1}(\hat{\Theta}_{t})$
$\preceq\quad(1-\frac{\alpha_{0}}{2\|P(\mathcal{C}_{t})\|})I$ (121)
where in (121) we applied the fact that $P(\mathcal{C}_{t})\preceq\|P(\hat{\Theta}_{t}\|I$.
Let
$L_{t}=$ ${P}^{-\frac{1}{2}}(\hat{\Theta}_{t})(A_{*}+B_{*}K(\mathcal{C}_{t})){P}^{\frac{1}{2}}(\hat{\Theta}_{t})$ (122)
noting that $P^{-1/2}(\hat{\Theta}_{t})$ exists by $P(\mathcal{C}_{t})\succ 0$. Then by (121) we have
$\|L_{t}\|\leq\sqrt{1-\frac{\alpha_{0}}{2\|P(\mathcal{C}_{t})\|}}$ $\leq 1-\frac{1}{2}(\frac{\alpha_{0}}{2\|P(\mathcal{C}_{t})\|})=:1-\gamma_{t}$ (123)
$\leq 1-\frac{1}{2}(\frac{\alpha_{0}}{2\|P_{*}\|})=:1-\gamma_{*}$ (124)
where inequality (123) follows from the fact that $\alpha_{0}<2|P(\mathcal{C}_{t})|$, and inequality (124) follows from $P(\hat{\Theta}t)\succeq P*$, as established in Lemma 15.
Furthermore, by
$\frac{\alpha_{0}}{2}{K}^{\top}(\mathcal{C}_{t})K(\mathcal{C}_{t})\preceq P(\mathcal{C}_{t})$
which follows from (119), we directly deduce that
$\|K(\mathcal{C}_{t})\|$ $\leq\sqrt{\frac{2\|P(\mathcal{C}_{t})\|}{\alpha_{0}}}:={\kappa}_{t}$ (125)
$\leq\sqrt{\frac{2\|P_{*}\|}{\alpha_{0}}}:=\kappa_{*}.$
By (122) we can write
$A_{*}+B_{*}K(\mathcal{C}_{t})=H_{t}L_{t}H_{t}^{-1}$
where $H_{t}={P}^{1/2}(\mathcal{C}_{t})$. We have already demonstrated upper bounds on $\|L_{t}\|$ and $\|K(\mathcal{C}_{t})\|$ by (123) and (125) respectively. By (119) we know that $P(\mathcal{C}_{t})\succeq\frac{\alpha_{0}}{2}I$ which yields $\|H_{t}^{-1}\|\leq\sqrt{2/\alpha_{0}}:=1/b_{0}$. Furthermore, since $\|H_{t}\|=\|P(\mathcal{C}_{t})\|^{1/2}:=B_{0}$ we obtain
$\|H_{t}^{-1}\|\|H_{t}\|$ $\leq\frac{1}{b_{0}}B_{0}=\sqrt{\frac{2\|P(\mathcal{C}_{t})\|}{\alpha_{0}}}$
$\leq\sqrt{\frac{2\|P_{*}\|}{\alpha_{0}}}.$
Conclusively, recalling parts (1) and (2) of Definition 1, we can conclude that any generated policies are $(\kappa,\gamma)$- strong stabilizing (also $(\kappa_{*},\gamma_{*})$- strong stabilizing) where $\gamma=\kappa^{-2}/2$ and $\gamma_{*}=\kappa_{*}^{-2}/2$.
In the next step, we prove that by tuning $\lambda$ and $\bar{p}_{t}$, and specifying $\bar{\epsilon}$ according to Lemma 14, the sequential property of the policies generated by the ARSLO algorithm (part (3) of Definition 1) is fulfilled. Beforehand, we require the following lemma, which is helpful.
Lemma 15
The following statement holds with probability at least $1-\delta$.
$P(\mathcal{C}_{t})\preceq P_{*}\preceq P(\mathcal{C}_{t+1})+\frac{\alpha_{0}\gamma_{t+1}}{2}.$
where $P_{*}$ is the optimal solution of dual program (56) and $\gamma_{t+1}=\kappa_{t+1}^{-2}/2$.
Proof The proof follows steps similar to those in Lemma 19 of cohen2019learning, with only minor differences. For the sake of completeness, we provide it here.
First we show that $P(\mathcal{C}_{t})\preceq P_{*}$.
By reapplying the perturbation lemma, Lemma 7, to the inequality of the relaxed dual SDP (83), we obtain:
$\begin{pmatrix}Q-P(\mathcal{C}_{t})&0\\ 0&R\end{pmatrix}+{\Theta}_{*}P(\mathcal{C}_{t}){\Theta}_{*}^{\top}\succeq 0$ (126)
By post multiplying (126) by $\begin{pmatrix}I\\ K_{*}\end{pmatrix}$ and pre-multiplying by its transpose, we get:
$Q-P(\mathcal{C}_{t})+K_{*}^{\top}RK_{*}+(A_{*}+B_{*}{K_{*}})^{\top}P(\mathcal{C}_{t})(A_{*}+B_{*}K_{*})\succeq 0.$ (127)
Combining (127) with (57), the optimal Riccati solution, yields:
$P(\mathcal{C}_{t})-P_{*}\preceq(A_{*}+B_{*}K_{*})^{\top}(P(\mathcal{C}_{t})-P_{*})(A_{*}+B_{*}K_{*})$ (128)
Utilizing Lemma 1 on (128), we conclude $P(\mathcal{C}_{t})\preceq P_{*}$.
Now we show the second inequality. From bertsekas2012dynamic we have:
$P_{*}\preceq Q+K^{\top}(\mathcal{C}_{t})RK(\mathcal{C}_{t})+(A_{*}+B_{*}K(\mathcal{C}_{t}))^{\top}P_{*}(A_{*}+B_{*}K(\mathcal{C}_{t}))$
Combining this with (94) and applying Lemma 1, along with the fact that the generated policies are $(\kappa,\gamma)$-strongly stabilizing, implies
$P_{*}-P(\mathcal{C}_{t})$ $\preceq\frac{16\|P(\mathcal{C}_{t})\|^{2}\mu_{t}}{\alpha_{0}^{2}}\|P(\mathcal{C}_{t})\|\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I$
$=4\kappa_{t}^{4}\mu_{t}\|P(\mathcal{C}_{t})\|\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I.$ (129)
By (97) of Lemma 14, we have
$\mu_{t}\|P(\mathcal{C}_{t})\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}$ $\preceq\frac{\alpha_{0}}{16{\kappa_{t}}^{6}}I$ (130)
in which we applied the fact that
$\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|\leq 1+\frac{2\|P(\mathcal{C}_{t})\|}{\alpha_{0}}\leq\frac{4\|P(\mathcal{C}_{t})\|}{\alpha_{0}}=2\kappa_{t}^{2}$
as $2\|P(\mathcal{C}_{t})\|>\alpha_{0}$.
Substituting (130) into (129) we can write
$P_{*}-P(\mathcal{C}_{t})$ $\preceq 4\kappa^{4}\mu_{t}\|P(\mathcal{C}_{t})\|\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I$
$\preceq 4\kappa_{t}^{4}\frac{\alpha_{0}}{16\kappa_{t}^{6}}I=\frac{\alpha_{0}{\gamma_{t}}}{2}I.$ (131)
Now, we provide the following lemma, which indeed demonstrates that the control designed by Algorithm 1 satisfies property (3) in Definition 1."
Lemma 16
Let $H^{2}_{t}=P(\mathcal{C}_{t})$ be the dual relaxed SDP solution. Then $\|H_{t+1}H_{t}\|\leq 1+\gamma_{t+1}/2$.
Proof The proof directly follows cohen2019learning, but we have chosen to redo it, as we will require a slightly modified version of this lemma for the proof of Theorem 4. Recalling definition of $H_{t}=P^{1/2}(\mathcal{C}_{t})$ we can write:
$\|H_{t+1}^{-1}H_{t}\|^{2}$ $=\|P^{-\frac{1}{2}}(\mathcal{C}_{t+1})P^{\frac{1}{2}}(\mathcal{C}_{t})\|^{2}$
$=\|P^{-\frac{1}{2}}(\mathcal{C}_{t+1})P(\mathcal{C}_{t})P^{-\frac{1}{2}}(\mathcal{C}_{t+1})\|$
$\leq\|P^{-\frac{1}{2}}(\mathcal{C}_{t+1})(P(\mathcal{C}_{t+1})+\frac{{\gamma_{t+1}}\alpha_{0}}{2}I)P^{-\frac{1}{2}}(\mathcal{C}_{t+1})\|$ (132)
$=\|I+\frac{{\gamma_{t+1}}\alpha_{0}}{2}P^{-1}(\mathcal{C}_{t+1})\|$
$\leq 1+\frac{{\gamma_{t+1}}\alpha_{0}}{2}\|P^{-1}(\mathcal{C}_{t+1})\|$
$\leq 1+{\gamma_{t+1}}$
where (132) follows by Lemma 15 and the last inequality holds because $\|P^{-1}(\mathcal{C}_{t+1})\|\leq 2/\alpha_{0}$. This concludes that $\|H_{t+1}^{-1}H_{t}\|\leq\sqrt{1+{\gamma_{t+1}}}\leq 1+{\gamma_{t+1}}/2$.
G.2 Proof of Proposition 1
Proof Proof follows by cohen2019learning with an only difference that now we have an input perturbation as well. Let $M=A_{*}+B_{*}K_{t}$ where $K_{t}^{\prime}$s are strong sequential stabilizing policies generated by ARSLO algorithm. The closed loop dynamics then is written as follows:
$x_{t}=M_{1}x_{1}+\sum_{s=1}^{t-1}M_{s}(B_{*}\nu_{s}+\omega_{s+1})$
where
$M_{t}=I,\;\;M_{s}=M_{s+1}(A_{*}+B_{*}K_{s})=\prod_{j=s}^{t-1}(A_{*}+B_{*}K_{j}),\;\forall 1\leq s\leq t-1.$
Recalling Definition 1, we can set $A_{*}+B_{*}K_{j}=H_{j}L_{j}H_{j}^{-1}$, then for all $1\leq s\leq t$ where $\|L_{j}\|\leq 1-\gamma_{j}$ and $\|H^{-1}_{j+1}H_{j}\|\leq 1+\gamma_{j+1}$ we can write
$\|M_{s}\|\leq$ $\lVert(H_{t-1}L_{t-1}H^{-1}_{t-1})...(H_{s+1}L_{s+1}H^{-1}_{s+1})(H_{s}L_{s}H^{-1}_{s})\rVert$
$\leq$ $\|H_{t-1}\|\;\|(L_{t-1}H^{-1}_{t-1}H_{t-2})...(L_{s+2}H^{-1}_{s+2}H_{s+1})(L_{s+1}H^{-1}_{s+1}H_{s})\|\>\|L_{s}\|\>\|H_{s}\|^{-1}$
$\leq$ $\|H_{t-1}\|\bigg(\prod_{j=s+1}^{t-1}\|L_{j}\|\;\|H_{j}^{-1}H_{j-1}\|\bigg)\|H_{s}^{-1}\|$
$\leq$ $\sqrt{P(\mathcal{C}_{t-1})}\bigg(\prod_{j=s+1}^{t-1}(1-\gamma_{j})\;(1+\gamma_{j}/2)\bigg)(1-\gamma_{s})\sqrt{\frac{2}{\alpha_{0}}}$
$\leq$ $\kappa_{{t-1}}\bigg(\prod_{j=s+1}^{t-1}(1-\gamma_{j}/2)\;\bigg)(1-\gamma_{s}/2)$
$\leq$ ${\kappa_{*}}(1-{\gamma_{*}}/2)^{t-s}$
where $\|H_{j}^{-1}H_{j-1}\|\leq(1+\gamma_{j}/2)$ holds for all $j$ by the sequentiality property of the generated policies. We also applied the fact that $H_{t}=P^{1/2}(\mathcal{C}_{t})\succeq\sqrt{\alpha_{0}/2}\,I$. Finally, we used the fact that $\kappa_{t}\leq\kappa_{*}$, as well as the inequality $(1-\gamma_{j}/2)\leq(1-\gamma_{*}/2)$, where the latter holds by definition of $\gamma_{*}:=\kappa_{*}^{-2}/2$.
Upper-bounding state norm yields
$\|x_{t}\|$ $\leq\|M_{1}\|\|x_{1}\|+\sum_{s=1}^{t-1}\|M_{s+1}\|\|B_{*}\eta_{s}+\omega_{s}\|$
$\leq\kappa_{*}(1-\gamma_{*}/2)^{t-1}\|x_{1}\|+\kappa_{*}\sum_{s=1}^{t-1}(1-{\gamma_{*}}/2)^{t-s-1}\|B_{*}\eta_{s}+\omega_{s}\|$
$\leq\kappa_{*}e^{-{\gamma_{*}}(t-1)/2}\|x_{1}\|+\frac{2{\kappa_{*}}}{{\gamma_{*}}}\max_{1\leq s\leq t}\|B_{*}\eta_{s}+\omega_{s}\|.$
Recalling the input perturbation noise definition, one can write:
$B_{*}\eta_{s}+\omega_{s}\sim\mathcal{N}\bigg(0,B_{*}\frac{\bar{p}_{t}\sigma_{\omega}^{2}\big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}\big)}{\sqrt{t+\bar{c}}}B_{*}^{\top}+\sigma_{\omega}^{2}I\bigg).$
Now, with the choice of $\bar{c}$ in (114), which guarantees
$B_{*}\frac{\bar{p}_{t}\sigma_{\omega}^{2}\big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|I}{\alpha_{0}}\big)}{\sqrt{t+\bar{c}}}B_{*}^{\top}\preceq\sigma_{\omega}^{2}I$
by applying the Hanson–Wright concentration inequality, we obtain
$\max_{1\leq s\leq t}\|B_{*}\eta_{s}+\omega_{s}\|\leq\sqrt{20\sigma_{\omega}^{2}n\log\frac{t}{\delta}}.$
Consequently, we can write
$\|x_{t}\|$ $\leq\kappa_{*}e^{-{\gamma_{*}}(t-1)/2}\|x_{1}\|+\frac{2\kappa_{*}}{\gamma_{*}}\sqrt{20\sigma_{\omega}^{2}n\log\frac{t}{\delta}}.$
Lemma 17
The following holds with probability at least $1-\delta$:
$\|x_{t}\|^{2}\leq 4\frac{\kappa_{*}^{2}}{\gamma_{*}^{2}}\big(\frac{\|x_{1}\|}{\sqrt{\log\frac{1}{\delta}}}+\sqrt{20n\sigma_{\omega}^{2}}\big)^{2}\log\frac{t}{\delta}:=X^{2}_{t}(\kappa_{*})$ (133)
and
$\|z_{t}\|^{2}\leq{64\kappa_{*}^{6}(1+\kappa_{*}^{2})}\bigg(\frac{\|x_{1}\|}{\log\frac{1}{\delta}}+\sqrt{20n\sigma_{\omega}^{2}}\bigg)^{2}\log\frac{t}{\delta}:=c_{z}^{2}(\kappa_{*})\log\frac{t}{\delta}=:\zeta_{t}^{2}(\kappa_{*}).$ (134)
Proof From (27), we can express
$\|x_{t}\|^{2}\leq 4\frac{\kappa_{*}^{2}}{\gamma_{*}^{2}}\big(\frac{\|x_{1}\|}{\sqrt{\log\frac{1}{\delta}}}+\sqrt{20n\sigma_{\omega}^{2}}\big)^{2}\log\frac{t}{\delta}:=X_{t}.$
Recalling that
$z_{t}=\begin{pmatrix}x_{t}\\ K(\mathcal{C}_{t})x_{t}+\eta_{t}\end{pmatrix}$
we obtain
$\|z_{t}\|^{2}\leq$ $2(1+\kappa_{*}^{2})\|x_{t}\|^{2}+2\|\eta_{t}\|^{2}.$
Furthermore, by the definition of $\bar{c}$ in (114), it holds that
$\|\eta_{t}\|\leq\sqrt{10n\sigma_{\omega}^{2}\log\frac{t}{\delta}}$
which in turn yields
$\|z_{t}\|^{2}\leq\frac{16\kappa^{2}(1+\kappa_{*}^{2})}{\gamma_{*}^{2}}\bigg(\frac{\|x_{1}\|}{\log\frac{1}{\delta}}+\sqrt{20n\sigma_{\omega}^{2}}\bigg)^{2}\log\frac{t}{\delta}.$
G.3 Proof of Corollary 1
Proof The proof follows from the definition of $\lambda$ in (20), which has the same order as $a^{2}$, namely $\mathcal{O}(\bar{\kappa}_{1}^{28})$. Moreover, by the definition of $\bar{\kappa}_{1}$ in (21), it is of the same order as $\kappa_{1}$, which is upper-bounded by $\kappa_{*}$. Finally, since $\bar{\epsilon}(\bar{\kappa}_{1})$ defined in (24) is of order $\mathcal{O}(1/\sqrt{\lambda})$, the claim follows.
Appendix H Regret Bound Analysis of ARSLO algorithm
H.1 Regret Decomposition
We begin by decomposing the regret bound. The structure of the decomposition follows the approach of cohen2019learning, with modifications necessitated by the presence of input perturbation. These changes slightly alter the form of the regret decomposition and introduce additional terms contributing to the final bound.
For the instantaneous regret we have:
$r_{k}=x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k}-J^{\star}\leq$ $x_{k}^{\top}\big(Q+K^{\top}(\mathcal{C}_{k})RK(\mathcal{C}_{k})\big)x_{k}+2\eta_{k}^{\top}RK(\mathcal{C}_{k})x_{k}+\eta^{\top}_{k}R\eta_{k}-\sigma_{\omega}^{2}\|P_{k}\|_{*}.$
Here, we used the following fact about the average expected cost:
$J_{*}\geq\sigma_{\omega}^{2}\|P(\mathcal{C}_{t})\|_{*}$
which follows from $J_{*}\geq\sigma_{\omega}^{2}\|P_{*}\|_{*}$ together with $\|P_{*}\|_{*}\geq\|P(\mathcal{C}_{t})\|_{*}$.
By Lemma 13 we can write
$x_{k}^{\top}\big(Q+K^{\top}(\mathcal{C}_{k})RK(\mathcal{C}_{k})\big)x_{k}=$ $x_{k}^{\top}P(\mathcal{C}_{k})x_{k}-x_{k}^{\top}(\hat{A}_{k}+\hat{B}_{k}K(\mathcal{C}_{k}))^{\top}P(\mathcal{C}_{k})(\hat{A}_{k}+\hat{B}_{k}K(\mathcal{C}_{k}))x_{k}$
$+\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|_{*}\bar{z}_{k}^{\top}V^{-1}_{\tau(k)}\bar{z}_{k}$ (135)
where
$\bar{z}_{k}:=\begin{pmatrix}x_{k}\\ K(\mathcal{C}_{k})x_{k}\end{pmatrix}$.
Under the ARSLO algorithm, the policy update condition $\det(V_{k})>(1+\beta)\det(V_{\tau(k)})$, where $\tau(k)$ denotes the last policy update time prior to $k$, implies that for all $k$ in the same epoch, $A_{k}=A_{\tau(k)}$, $B_{k}=B_{\tau(k)}$, and $K(\mathcal{C}_{k})=K(\mathcal{C}_{\tau(k)})$.
Using (135), we obtain
$r_{k}\leq$ $x_{k}^{\top}P(\mathcal{C}_{k})x_{k}-x_{k+1}^{\top}P(\mathcal{C}_{k})x_{k+1}$
$+x_{k}^{\top}(A_{*}+B_{*}K(\mathcal{C}_{k}))^{\top}P(\mathcal{C}_{k})(A_{*}+B_{*}K(\mathcal{C}_{k}))x_{k}$
$-x_{k}^{\top}(\hat{A}_{k}+\hat{B}_{k}K(\mathcal{C}_{k}))^{\top}P(\mathcal{C}_{k})(\hat{A}_{k}+\hat{B}_{k}K(\mathcal{C}_{k}))x_{k}$ (136)
$+2\omega_{k}^{\top}P(\mathcal{C}_{k})(A_{*}+B_{*}K(\mathcal{C}_{k}))x_{k}+\omega_{k}^{\top}P(\mathcal{C}_{k})\omega_{k}-\sigma_{\omega}^{2}\|P(\mathcal{C}_{t})\|_{*}+\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|\bar{z}_{k}^{\top}V^{-1}_{\tau(k)}\bar{z}_{k}$
$+2\eta_{k}^{\top}RK(\mathcal{C}_{k})x_{k}+\eta^{\top}_{k}R\eta_{k}.$ (137)
The terms in (137) are explicit additional contributions arising from the input perturbation. Also, in deriving (136), we simply added and subtracted
$x_{k+1}^{\top}P(\mathcal{C}_{k})x_{k+1}=$ $x_{k}^{\top}(A_{*}+B_{*}K(\mathcal{C}_{k}))^{\top}P(\mathcal{C}_{k})(A_{*}+B_{*}K(\mathcal{C}_{k}))x_{k}$
$+2\omega_{k}^{\top}P(\mathcal{C}_{k})(A_{*}+B_{*}K(\mathcal{C}_{k}))x_{k}+\omega_{k}^{\top}P(\mathcal{C}_{k})\omega_{k}.$
Next, applying Lemma 7 yields
$\bar{z}_{k}^{\top}\bigg(\Theta_{*}P(\mathcal{C}_{k})\Theta_{*}^{\top}-\hat{\Theta}_{k}P(\mathcal{C}_{k})\hat{\Theta}^{\top}_{k}\bigg)\bar{z}_{k}\leq\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|\;\bar{z}_{k}^{\top}V^{-1}_{\tau(k)}\bar{z}_{k}.$
which implies
$x_{k}^{\top}(A_{*}+B_{*}K(\mathcal{C}_{k}))^{\top}P(\mathcal{C}_{k})(A_{*}+B_{*}K(\mathcal{C}_{k}))x_{k}$
$-x_{k}^{\top}(\hat{A}_{k}+\hat{B}_{k}K(\mathcal{C}_{k}))^{\top}P(\mathcal{C}_{k})(\hat{A}_{k}+\hat{B}_{k}K(\mathcal{C}_{k}))x_{k}\leq\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|\;\bar{z}_{k}^{\top}V^{-1}_{\tau(k)}\bar{z}_{k}$
providing an upper bound for (136).
Since $V_{\tau(k)}$ is constructed using $z_{k}$ and not $\bar{z}_{k}$, and
$\bar{z}_{k}=z_{k}-\begin{pmatrix}0\\ \eta_{k}\end{pmatrix}$.
we obtain the structural bound
$\bar{z}_{k}^{\top}V^{-1}_{\tau(k)}\bar{z}_{k}\leq 2z_{k}^{\top}V^{-1}_{\tau(k)}z_{k}+2\eta_{k}^{\top}V^{-1}_{\tau(k)}\eta_{k}.$
Combining these results gives
$r_{k}\leq$ $x_{k}^{\top}P(\mathcal{C}_{k})x_{k}-x_{k+1}^{\top}P(\mathcal{C}_{k})x_{k+1}$
$+2\omega_{k}^{\top}P(\mathcal{C}_{k})(A_{*}+B_{*}K(\mathcal{C}_{k}))x_{k}$
$+\omega_{k}^{\top}P(\mathcal{C}_{k})\omega_{k}-\sigma_{\omega}^{2}\|P(\mathcal{C}_{t})\|_{*}$
$+4\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|z_{k}^{\top}V^{-1}_{\tau(k)}z_{k}$ (138)
$+2\eta_{k}^{\top}RK(\mathcal{C}_{k})x_{k}$
$+\eta^{\top}_{k}\big(R+4\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|V^{-1}_{\tau(k)}\big)\eta_{k}.$
To further bound term (138), note that $\tau(k)$ is the last policy update time, and the algorithm ensures $\det(V_{k})\leq(1+\beta)\det(V_{\tau(k)})$. Consequently, $z_{k}^{\top}V^{-1}_{\tau(k)}z_{k}\leq(1+\beta)z_{k}^{\top}V^{-1}_{k}z_{k}$ and hence
$4\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|z_{k}^{\top}V^{-1}_{\tau(k)}z_{k}\leq 4(1+\beta)\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|z_{k}^{\top}V^{-1}_{k}z_{k}.$
For the regret analysis, define the good event:
$\mathcal{E}_{t}=\left\{\Theta_{*}\in\mathcal{C}_{t}(\delta)\;\text{and}\;\|z_{t}\|^{2}\leq\zeta_{t}^{2}(\kappa_{*})\right\},$
where $\mathcal{C}_{t}(\delta)$ is the confidence set in (7) and $\zeta_{t}^{2}(\kappa_{*})$ is the bound in (134).
Finally, the accumulated regret decomposes as
$R_{\textit{ARSLO}}(t)\leq\sum_{j=1}^{6}R_{j}(t)$
where:
$R_{1}(t)=\sum_{k=1}^{t}\big(x_{k}^{\top}P(\mathcal{C}_{k})x_{k}-x_{k+1}^{\top}P(\mathcal{C}_{k})x_{k+1}\big)1_{\mathcal{E}_{k}},$ (139a)
$R_{2}(t)=\sum_{k=1}^{t}w^{\top}_{k}P(\mathcal{C}_{k})\big(A_{*}+B_{*}K(\mathcal{C}_{k})\big)x_{k}1_{\mathcal{E}_{k}},$ (139b)
$R_{3}(t)=\sum_{k=1}^{t}\big(w^{\top}_{k}P(\mathcal{C}_{k})w_{k}-\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|\big)1_{\mathcal{E}_{k}},$ (139c)
$R_{4}(t)=\sum_{k=1}^{t}4(1+\beta)\|P(\mathcal{C}_{k})\|\mu_{\tau(k)}\big(z^{\top}_{k}V_{k}^{-1}z_{k}\big)1_{\mathcal{E}_{k}`},$ (139d)
$R_{5}(t)=\sum_{k=1}^{t}2\eta_{k}^{\top}RK(\mathcal{C}_{k})x_{k}1_{\mathcal{E}_{k}},$ (139e)
$R_{6}(t)=\sum_{k=1}^{t}\eta^{\top}_{k}\big(R+4\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|V^{-1}_{\tau(k)}\big)\eta_{k}1_{\mathcal{E}_{k}}.$ (139f)
H.2 Proof of Theorem 2
Proof
To derive the regret bound, we upper-bound each term in (139a-139f) separately, setting $\beta=1$ to remain consistent with the ARSLO algorithm (Algorithm 1). Throughout the regret analysis, we make use of the norm upper bounds established in Lemma 17.
Lemma 18
With probability at least $1-\delta$, the following upper bound holds for $|R_{1}(t)|$:
$|R_{1}(t)|\lesssim\|P_{*}\|^{4}n(n+m)\log_{2}^{2}\frac{nt}{\delta}$
Proof We begin by rewriting $R_{1}(t)$ as
$R_{1}(t)=$ $\sum_{k=1}^{t}(x^{\top}_{k}P(\mathcal{C}_{k})x_{k}-x^{\top}_{k+1}P(\mathcal{C}_{k})x_{k+1})1_{\mathcal{E}_{k}}$
$=$ $\sum_{k=2}^{t-1}(x^{\top}_{k}(P(\mathcal{C}_{k})-P(\mathcal{C}_{k-1})x_{k})1_{\mathcal{E}_{k}}$
$+x_{1}^{\top}P(\hat{\Theta}_{1})x_{1}-x_{t+1}^{\top}P(\mathcal{C}_{t})x_{t+1}.$
The term $x_{k}^{\top}(P(\mathcal{C}_{k})-P(\mathcal{C}_{k-1}))x_{k}$ is nonzero only when the policy is updated. Thus, to upper-bound $R_{1}$, it suffices to bound the total number of policy updates. Let $N(t)$ denote the number of updates up to time $t$.
Recalling the policy-update rule with $\beta=1$, we have
$\det(V_{\tau(t)})\geq 2^{N(t)}\det(\lambda I)$
where $\tau(t)$ is the last update time prior to $t$.
This yields
$N(t)\leq\log_{2}(\frac{\det(V_{\tau(t)})}{\det(\lambda)})\leq\log_{2}(\frac{\det(V_{t})}{\det(\lambda)}).$
From the definition of $\lambda$ in (115) and similar to what we carried out in (107) we have
$\log_{2}(\frac{\det(V_{t})}{\det(\lambda)})\leq 2(n+m)\log_{2}\frac{2t}{\delta}.$
This yields
$N(t)\leq$ $2(n+m)\log_{2}\frac{2t}{\delta}.$
Using $\|P(\mathcal{C}_{k})\|\leq\|P_{*}\|$ and $X_{t}(\kappa_{*}):=\max_{k\leq t}\|x_{k}\|$, we obtain
$|R_{1}|\leq\|P_{*}\|X_{t}^{2}(\kappa_{*})+2\|P_{*}\|(n+m)X_{t}^{2}(\kappa_{*})\log_{2}\frac{2t}{\delta}$
Finally, incorporating the dependence of $X_{t}$ on $\|P_{*}\|$ and $(n,m)$ gives the stated bound, completing the proof.
Lemma 19
With probability at least $1-\delta/4$, the following bound holds:
$|R_{2}(t)|\lesssim\sqrt{\|P_{*}\|^{5}n\,t\,\log^{3}\frac{t}{\delta}}.$
Proof Applying Lemma 30 from cohen2019learning, we obtain
$R_{2}(t)=\sum_{k=1}^{t}\underbrace{x^{\top}_{k}\big(A_{*}+B_{*}K(\mathcal{C}_{k})\big)^{\top}P(\mathcal{C}_{k})}_{\bar{v}^{\top}_{k}}w_{k}\leq 2\sigma_{\omega}\bar{D}\sqrt{\frac{2}{\delta}},$
where $\sum_{k=1}^{t}\bar{v}_{k}^{\top}\bar{v}_{k}\leq\bar{D}^{2}$.
We now derive an upper bound on this quantity. Observe that
$\sum_{k=1}^{t}\bar{v}_{k}^{\top}\bar{v}_{k}$ $\leq\|P_{*}\|^{2}\sum_{k=1}^{t}x_{k}^{\top}\big(A_{*}+B_{*}K(\mathcal{C}_{k})\big)^{\top}\big(A_{*}+B_{*}K(\mathcal{C}_{k})\big)x_{k}$
$\leq\|P_{*}\|^{2}(1-\gamma_{*})^{2}\sum_{k=1}^{t}x_{k}^{\top}x_{k}$
$\leq\|P_{*}\|^{2}\,X_{t}^{2}(\kappa_{*})\,t:=\bar{D}^{2}$
where we used the bounds $\|P(\mathcal{C}_{t})\|\leq\|P_{*}\|$ and $\|A_{*}+B_{*}K(\mathcal{C}_{t})\|\leq 1-\gamma_{*}<1$. Hence,
$|R_{2}(t)|\leq 2\sigma_{\omega}\|P_{*}\|X_{t}(\kappa_{*})\sqrt{t\frac{2}{\delta}}.$
Finally, substituting the expression for $X_{t}(\kappa_{*})$ from (133) yields the desired result, completing the proof.
Lemma 20
The term $R_{3}(t)$ is bounded as follows:
$R_{3}(3)\lesssim\|P_{*}\|_{*}\sqrt{t\log^{3}\frac{t}{\delta}}$
with probability at least $1-\delta$.
Proof The proof follows directly from cohen2019learning; for completeness, we provide it here.
Let
$X_{k}:=\omega_{k}^{\top}P(\mathcal{C}_{k})\omega_{k}-\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|_{*},\quad k=1,\dots,t.$
Since $P(\mathcal{C}_{k})$ is $\mathcal{F}_{k-1}$-measurable and $\omega_{k}$ is independent of $\mathcal{F}_{k-1}$ with $\mathbb{E}[\omega_{k}\omega_{k}^{\top}]=\sigma_{\omega}^{2}I$, we have
$\mathbb{E}[X_{k}\mid\mathcal{F}_{k-1}]=0.$
So $X_{k}$’s forms a martingale difference sequence. Moreover, since $P(\mathcal{C}_{k})\succeq 0$ and $\|P(\mathcal{C}_{k})\|_{*}\leq\|P_{*}\|_{*}$ we have
$X_{k}\geq-\sigma^{2}\|P_{*}\|_{*}\quad\text{a.s.}$
However, $X_{k}$ is not almost surely bounded from above, which prevents a direct application of Azuma’s inequality. To address this, we define the truncated random variable
$\widetilde{X}_{k}:=X_{k}\cdot\mathbf{1}\{X_{k}\leq\Gamma\},\qquad\Gamma:=5\|P_{*}\|_{*}\sigma_{\omega}^{2}\log\frac{4t}{\delta}.$
Then $\sigma_{\omega}^{2}\|P_{*}\|_{*}\leq\widetilde{X}_{k}\leq\Gamma$, allowing Azuma’s inequality to apply.
Let $\widetilde{Y}_{k}:=\widetilde{X}_{k}-\mathbb{E}[\widetilde{X}_{k}\mid\mathcal{F}_{k-1}]$. Then $|\widetilde{Y}_{k}|\leq\Gamma$ a.s. By Azuma’s inequality, Lemma 5 , we, with probability at least $1-\delta/2$ , can write
$\sum_{k=1}^{t}\widetilde{Y}_{k}\leq\Gamma\sqrt{2t\log\frac{1}{\delta}}$
which implies
$\sum_{k=1}^{t}\widetilde{X}_{k}\leq\sum_{k=1}^{t}\mathbb{E}[\widetilde{X}_{k}]+8\sigma_{\omega}^{2}\|P_{*}\|_{*}\sqrt{t\log\frac{1}{\delta}\log^{2}\frac{4t}{\delta}}.$
Next, we control the truncation probability using the Hanson–Wright inequality, Lemma 6.
For each $k$, it implies that for $\delta^{\prime}\in(0,1/e)$:
$\Pr\big(w_{k}^{\top}P(\mathcal{C}_{k})w_{k}\geq\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|_{*}+4\sigma^{2}\|P^{1/2}(\mathcal{C}_{k})\|\,\|P^{1/2}(\mathcal{C}_{k})\|_{F}\log(1/\delta^{\prime})\big)\leq\delta^{\prime}.$
Using the facts that $\|P^{1/2}(\mathcal{C}_{k})\|_{F}\|P^{1/2}(\mathcal{C}_{k})\|\leq\|P(\mathcal{C}_{k})\|_{*}$ and $\|P(\mathcal{C}_{k})\|_{*}\leq\|P_{*}\|_{*}$ we get
$w_{k}^{\top}P(\mathcal{C}_{k})w_{k}\leq\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|_{*}+4\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|_{*}\log(1/\delta^{\prime})\leq 5\sigma_{\omega}^{2}\|P_{*}\|_{*}\log(1/\delta^{\prime}).$
Choosing $\delta^{\prime}=\delta/(2t)$, we obtain
$\Pr(X_{k}>\Gamma)\leq\frac{\delta}{2t}.$
By the union bound, with probability at least $1-\delta/2$, $X_{k}\leq\Gamma$ for all $k$, i.e., $\widetilde{X}_{k}=X_{k}$.
Azuma inequality holds for $\widetilde{X}_{k}$ with probability $1-\delta/2$, and no-truncation event holds with probability $1-\delta/2$. By union bound, both hold simultaneously with probability at least $1-\delta$. On this event:
$\sum_{k=1}^{t}X_{k}=\sum_{k=1}^{t}\widetilde{X}_{k}\leq 8\sigma_{\omega}^{2}\|P_{*}\|_{*}\sqrt{t\log\frac{1}{\delta}\log^{2}\frac{4t}{\delta}}.$
This implies
$\sum_{k=1}^{t}\omega_{k}^{\top}P_{k}\omega_{k}-\sigma_{\omega}^{2}\|P_{k}\|_{*}\leq\|P_{*}\|_{*}\sqrt{t\log\frac{1}{\delta}\log^{2}\frac{4t}{\delta}}$
which completes the proof.
Lemma 21
The term $R_{4}(t)$ has the following upper bound.
$R_{4}(t)\lesssim\vartheta\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{6}\,t\,\log^{4}\frac{nt}{\delta}}$
Proof Recall that $\tau(t)$ denotes the most recent policy update time prior to time $t$. Using the definition $\mu_{t}=r_{t}+2\vartheta\sqrt{r_{t}}|V_{t}|^{1/2}$ and setting $\beta=1$ to align with the ARSLO algorithm, we can write
$R_{4}(t):=$ $4\sum_{k=1}^{t}(1+\beta)\|P(\mathcal{C}_{k})\|\mu_{\tau(k)}\big({z}^{\top}_{k}V_{k}^{-1}{z}_{k}\big)1_{\mathcal{E}_{k}`}$
$\leq$ $8\mu_{t}\|P_{*}\|\sum_{k=1}^{t}\big({{z}_{k}}^{\top}{V_{k}}^{-1}{{z}_{k}}\big)1_{\mathcal{E}_{k}}$
where the inequality uses $\mu_{\tau(t)}\leq\mu_{t}$ and $\|P(\mathcal{C}_{k})\|\leq\|P_{*}\|$ for all $k\geq 1$.
We decompose $\mu_{t}$ as follows:
$\mu_{t}=(r_{t}+2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2})\leq$ $\bigg(r_{t}+2\vartheta\,\sqrt{r_{t}}\sqrt{\lambda+t\zeta_{t}^{2}(\kappa_{*})}\bigg)$
$\leq$ $r_{t}+2\vartheta\sqrt{r_{t}\,\lambda}+2\vartheta\sqrt{r_{t}\,\zeta_{t}^{2}(\kappa_{*})\,t}.$
Moreover, by using (110) as an upper bound for $r_{t}$ and applying (134), we obtain the following upper bounds for the aforementioned terms.
$\mu_{t}\leq$ $16\sigma_{\omega}^{2}n(n+m)\log\frac{2nt}{\delta}$
$+8\vartheta\sigma_{\omega}^{2}\sqrt{\frac{\bar{c}n(n+m)}{10}\log\frac{2nt}{\delta}}$
$+8\vartheta c_{z}(\kappa_{*})\sigma_{\omega}\sqrt{n(n+m)\,t\,\log\frac{t}{\delta}\,\log\frac{2nt}{\delta}}$
$\lesssim$ $\vartheta\sqrt{n^{2}(n+m)\|P_{*}\|^{4}t\log^{2}\frac{nt}{\delta}}$
We observe that
${z_{k}}^{\top}{V_{k}}^{-1}{z_{k}}\leq\frac{z_{k}^{\top}z_{k}}{\lambda+\frac{\bar{p}_{t}\sigma_{\omega}^{2}\sqrt{k+\bar{c}}}{40}-\bar{C}}\leq\frac{\zeta_{k}^{2}}{\frac{\bar{p}_{k}\sigma_{\omega}^{2}\sqrt{k+\bar{c}}}{40}}<1.$ (140)
The second inequality follows from the specific choice of $\bar{c}$ that satisfies (105).
Since the condition (140) holds, we may apply Lemma 3. Consider an epoch within the interval $\tau_{i},\;\tau_{i+1})$. Over this epoch
$\sum_{t=\tau_{i}}^{\tau_{i+1}}z_{t}^{\top}V_{t}^{-1}z_{t}\leq 2\log\frac{\det(V_{\tau_{i+1}})}{\det(V_{\tau_{i}})}.$
Applying this bound across all epochs starting at $\tau_{1},...,\tau_{N(t)}$, and letting $N(t)$ denote the total number of policy updates prior to time $t$, we can write
$\sum_{t=1}^{t-1}z_{t}^{\top}V_{t}^{-1}z_{t}$ $\leq\sum_{i=1}^{N(t)-1}2\log\frac{\det(V_{\tau_{i+1}})}{\det(V_{\tau_{i}})}+2\log\frac{\det(V_{t})}{\det(V_{\tau_{N(t)}})}$
$\leq 2\log\frac{\det(V_{t})}{\det(\lambda I)}$ (141)
where the last inequality follows by telescoping the logarithmic terms.
By definition of $\lambda$ we have also
$\log\frac{\det(V_{t})}{\det(\lambda_{1}I)}\leq 2(n+m)\log\frac{2nt}{\delta}\lesssim(n+m)\log\frac{nt}{\delta}.$
as already shown in (107). Combining all bounds, we obtain
$R_{4}(t)\lesssim\vartheta\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{6}t\log^{4}\frac{nt}{\delta}}$
Lemma 22
The term $R_{5}$ has the following upper bound
$R_{5}(t)\lesssim\alpha_{1}\big(n^{4}(n+m)\big)^{1/4}\,t^{1/4}\,\sqrt{\vartheta\|P_{*}\|_{*}\|P_{*}\|^{11}\log^{3}\frac{nt}{\delta}}$
Proof Define
$\bar{v}_{k}^{\top}=2x_{k}^{\top}K^{\top}(\mathcal{C}_{t})R^{\top}.$
Then
$R_{5}(t)=\sum_{k=1}^{t}\bar{v}_{k}^{\top}\eta_{k}1_{\mathcal{E}_{k}}.$
Let $Y_{k}:=\bar{v}_{k}^{\top}\eta_{k}$. Conditioned on all measurements up to time $k$, $Y_{k}$ is a zero-mean Gaussian random variable. We may write $Y_{k}=m_{k}\,f_{k}$, where $f_{k}\sim\mathcal{N}(0,1)$ and $\textit{Var}(Y_{k})=m_{k}^{2}$.
Applying Theorem 23 of cohen2019learning with $V=\beta$ yields
$\frac{(\sum_{k=1}^{t}f_{k}m_{k})^{2}}{\beta+\sum_{k=1}^{t}m_{k}^{2}}\leq 2\log\frac{1+\beta^{-1}\sum_{k=1}^{t}m_{k}^{2}}{\delta}.$ (142)
We now upper-bound the quadratic term:
$\sum_{k=1}^{t}m_{k}^{2}$ $\leq\sum_{k=1}^{t}\|\tilde{v}_{k}\|^{2}\|\Gamma_{k}\|^{2}\leq\sum_{k=1}^{t}\|\bar{v}_{k}\|^{2}\frac{2\sigma_{\omega}^{2}\bar{p}_{k}\kappa_{*}^{2}}{\sqrt{k+\bar{c}}}$
$\leq 2\big(\max_{1\leq k\leq t}\|\bar{v}_{k}\|^{2}\big)\bar{p}_{t}\sigma_{\omega}^{2}\kappa_{*}^{2}\sum_{k=1}^{t}\frac{1}{\sqrt{k+\bar{c}}}$
$\leq 8\alpha^{2}_{1}\bar{p}_{t}\sigma_{\omega}^{2}\kappa_{*}^{4}X_{t}^{2}(\kappa_{*})\sqrt{t+\bar{c}}=:D.$ (143)
Substituting the bound (143) into (142) and choosing $\beta=D$, we obtain, with probability at least $1-\delta/2$
$\underbrace{(\sum_{k=1}\bar{v}_{k}^{\top}\eta_{k})^{2}}_{R_{5}^{2}(t)}=(\sum_{k=1}^{t}f_{k}m_{k})^{2}\leq 4D\log\frac{4}{\delta}.$
Equivalently,
$R_{5}\leq 4\sqrt{2\bar{p}_{t}\log\frac{4}{\delta}}\alpha_{1}\sigma_{\omega}\kappa_{*}^{2}X_{t}(\kappa_{*})(t+\bar{c})^{1/4}.$ (144)
From (116), $\bar{p}_{t}$ satisfies the upper-bound
$\bar{p}_{t}\lesssim\vartheta\|P_{*}\|^{7}\sqrt{n^{2}(n+m)}\log\frac{nt}{\delta}$ (145)
and for $X^{2}_{t}(\kappa_{*})$ we have
$X^{2}_{t}(\kappa_{*})\lesssim\|P_{*}\|^{3}n\log\frac{t}{\delta}.$
Substituting these expressions into (144) gives
$R_{5}(t)\lesssim\alpha_{1}(n^{4}(n+m))^{1/4}\sqrt{\vartheta\|P_{*}\|^{12}\log^{3}\frac{nt}{\delta}}(t+\bar{c})^{1/4}.$
Lemma 23
The term $R_{6}(t)$ has the following upper bound.
$R_{6}(t)\lesssim\alpha_{1}\vartheta\sqrt{n^{2}(n+m)m^{2}\|P_{*}\|^{16}t\log^{4}\frac{nt}{\delta}}$
with probability at least $1-\delta$.
Proof The following inequality holds
$R_{6}(t)=$ $\sum_{k=1}^{t}\eta^{\top}_{k}\big(R+4\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|V^{-1}_{\tau(k)}\big)\eta_{k}1_{\mathcal{E}_{k}}$
$\leq$ $\sum_{k=1}^{t}\eta^{\top}_{k}\big(R+\alpha_{0}I\big)\eta_{k}1_{\mathcal{E}_{k}}$
$\leq$ $2\alpha_{1}\sum_{k=1}^{t}\eta^{\top}_{k}\eta_{k}1_{\mathcal{E}_{k}}.$ (146)
The second inequality follows from (118), namely
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I\quad\forall t$
which is ensured by appropriate tuning of $\lambda$, $\bar{p}_{t}$, and $\bar{\epsilon}$, as followed by the proof of Lemma 14.
Then problem is reduced to upper-bound (146).
We can re-express
$\eta_{k}=\Gamma_{k}^{1/2}z_{k}\quad z_{k}\sim\mathcal{N}(0,I_{m}).$
Then
$\sum_{k=1}^{t}\eta_{k}^{\top}\eta_{k}=\sum_{k=1}^{t}z_{k}^{\top}\Gamma_{k}z_{k}.$
No by applying Hanson–Wright inequality (Lemma 6), it yields
$z_{k}^{\top}\Gamma_{k}z_{k}\leq$ $\|\Gamma^{1/2}_{k}\|^{2}_{F}+2\|\Gamma^{1/2}\|_{F}\|\Gamma_{k}^{1/2}\|\sqrt{\log\frac{k}{\delta}}+2\|\Gamma_{k}^{1/2}\|^{2}\log\frac{k}{\delta}$
$\leq$ $5\|\Gamma^{1/2}_{k}\|^{2}_{F}\log\frac{k}{\delta}:=5\|\Gamma_{k}\|_{*}\log\frac{k}{\delta}$
with probability $1-\delta/k$.
Hence,
$\sum_{k=1}^{t}z_{k}^{\top}\Gamma_{k}z_{k}$ $\leq 5\sum_{k=1}^{t}\|\Gamma_{k}\|_{*}\log\frac{k}{\delta}\leq 10n\bar{p}_{t}\sigma_{\omega}^{2}\log\frac{t}{\delta}\kappa_{*}^{2}\sum_{k=1}^{t}\frac{1}{\sqrt{k+\bar{c}}}$
$\leq 20\,m\,\bar{p}_{t}\,\sigma_{\omega}^{2}\kappa_{*}^{2}\,\log\frac{t}{\delta}\big(\sqrt{t+\bar{c}}-\sqrt{\bar{c}}\big)$
in which we applied the fact that $\|\Gamma_{k}\|_{*}\leq m\|\Gamma_{k}\|$.
Finally applying (145) and definition of $\kappa_{*}$ yields
$R_{6}(t)\lesssim\alpha_{1}\vartheta\sqrt{n^{2}(n+m)m^{2}\|P_{*}\|^{16}t\log^{4}\frac{nt}{\delta}}.$
Considering all bounded terms, we observe that the two most significant terms of the regret are $R_{4}(t)$ and $R_{6}(t)$. Among these, $R_{6}(t)$ is dominant in terms of dependence on $\|P_{*}\|$, while $R_{4}(t)$ is slightly worse than $R_{6}(t)$ in terms of dimensional dependence, namely $\mathcal{O}(\sqrt{n^{2}(n+m)^{3}})$ versus $\mathcal{O}(\sqrt{n^{2}(n+m)m^{2}})$. By selecting the more conservative bound in terms of dimension, we report the overall regret as
$R_{\text{ARSLO}}(t)\leq\mathcal{O}\big(\sqrt{\;n^{2}(n+m)^{3}\|P_{*}\|^{16}\;t\log^{4}\frac{t}{\delta}}\big)$
which completes the proof.
[…]
Inputs: $\Theta_{0}$, $\bar{\kappa}_{1}$, and $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ for some $\bar{\rho}\in[0,\,8)$ (provided by Algorithm 2), $\delta$, $\vartheta$
Compute $\bar{c}$ using (33).
Set $\lambda=\frac{\sigma_{\omega}^{2}\bar{c}}{40}(n+m)\log\frac{1}{\delta}$
Initialize $V_{1}=\lambda I$, $r_{1}=\lambda\bar{\epsilon}^{2}$, $\hat{\Theta}_{1}=\Theta_{0}$, and construct $\mathcal{C}_{1}$.
Set
$\beta(\bar{\rho})=\begin{cases}2\bar{\kappa}_{1}^{\frac{(6-\bar{\rho})}{2}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8\end{cases}.$ (147)
for $t=1,2,$... do
 if $\det(V_{t})>(1+\beta(\bar{\rho}))\det(V_{\tau})$ or $t=1$ then
      Calculate $\mu_{t}=r_{t}+\sqrt{r_{t}}\vartheta\|V_{t}\|^{1/2}$
     Compute $K(\mathcal{C}_{t})$ by (13) and through solving the relaxed primal SDP problem (12)
     Compute $P(\mathcal{C}_{t})$ through solving the relaxed dual SDP problem (17)
     Let $V_{\tau}=V_{t}$
 else
     Set $K(\mathcal{C}_{t})=K(\mathcal{C}_{t-1})$ and $P(\mathcal{C}_{t})=P(\mathcal{C}_{t-1})$
   end if
  Calculate $\Gamma_{t}$ by (36)-(37) and sample $\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big)$
  Play $u_{t}=K(\mathcal{C}_{t})x_{t}+\eta_{t}$
  Observe $x_{t+1}$, save $(x_{t+1},z_{t})$ to the data set
  Construct $X_{t}$ and $Z_{t}$ and compute $V_{t}$ using regularization parameter $\lambda=\mathcal{G}\log\frac{2nt}{\delta}$
  Calculate $\hat{\Theta}_{t}$ by (6) and $r_{t}$ by (8) and build confidence ellipsoid $\mathcal{C}_{t}$
end for
[…]
J.1 Proof of Theorem 3
[…]
From the proof of the previous lemma, we concluded that ensuring a policy is $(\kappa_{*},\gamma_{*})$-strongly stabilizing requires
$\mu_{t}\,\|P(\mathcal{C}_{t})\|\,V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I.$ (148)
To simultaneously control stability and regret, we impose a slightly stronger requirement:
$\mu_{t}\,\|P(\mathcal{C}_{t})\|\,V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I,$ (149)
where $\kappa_{t}=\sqrt{\frac{2\|P(\mathcal{C}_{t})\|}{\alpha_{0}}}>1$ and $\bar{\rho}\geq 0$. This condition immediately implies (148). By definition, (149) is equivalent to
$\mu_{t}\,\,V_{t}^{-1}\preceq\frac{1}{2\kappa_{t}^{\bar{\rho}+2}}I.$ (150)
Our goal is to select $\lambda$, $\bar{p}_{t}$, and to specify $\bar{\epsilon}$—the warm-up estimator error bound—so that (150) holds.
Recall that $\mu_{t}=r_{t}+2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}$, so sufficient conditions for (150) are
$r_{t}V_{t}^{-1}\preceq\frac{1}{4\kappa_{t}^{\bar{\rho}+2}}I$ (151)
$2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}V_{t}^{-1}\preceq\frac{1}{4\kappa_{t}^{\bar{\rho}+2}}I.$ (152)
By applying the result of Lemma 12, namely inequality (86), one can write
$2\vartheta\sqrt{r_{t}}\|V_{t}\|^{1/2}V_{t}^{-1}$ $\preceq\frac{2\vartheta\sqrt{r_{t}}\|V_{t}\|^{\frac{1}{2}}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}I.$
where
$\bar{C}:=\frac{\sigma_{\omega}^{2}\bar{c}}{80}.$ (153)
Then a sufficient condition for (152) is
$\frac{2\vartheta\sqrt{r_{t}}\|V_{t}\|^{\frac{1}{2}}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}$ $\leq\frac{1}{4\kappa_{t}^{\bar{\rho}+2}}$ (154)
and similarly, a sufficient condition for (151) is
$\frac{r_{t}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}\leq\frac{1}{4\kappa_{t}^{\bar{\rho}+2}}.$ (155)
Let us choose $\lambda$ such that
$\lambda\geq\bar{C}.$
Under this choice, a sufficient condition for (154) can be derived as follows:
$\frac{2\vartheta\sqrt{r_{t}}\|V_{t}\|^{\frac{1}{2}}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}-\bar{C}+\lambda}\leq$ $\frac{2\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\frac{\sigma_{\omega}^{2}\bar{p}_{t}}{80}\sqrt{t+\bar{c}}}$
$\leq$ $\frac{1}{4\kappa_{t}^{\bar{\rho}+2}}.$
To satisfy the inequality above, it suffices to require
$\bar{p}_{t}\geq\frac{640\kappa_{t}^{\bar{\rho}+2}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$
Similarly, a sufficient condition for (155) is given by
$\bar{p}_{t}\geq\frac{320\,r_{t}\,\kappa_{t}^{\bar{\rho}+2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$
We then choose $\bar{P}_{t}$ as
$\bar{p}_{t}:=\max\bigg\{\frac{640\kappa_{t}^{\bar{\rho}+2}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}},\;\frac{320\,r_{t}\,\kappa_{t}^{\bar{\rho}+2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}\bigg\}.$ (156)
Note that, for the specific choice of $\bar{c}$ introduced later, one can show that $\bar{p}_{t}$ is determined by the first term for all $t$’s.
Next, we proceed to specify $\bar{c}$ such that
$B_{*}\frac{\bar{p}_{t}\sigma_{\omega}^{2}\big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|I}{\alpha_{0}}\big)}{\sqrt{t+\bar{c}}}B_{*}^{\top}\preceq\sigma_{\omega}^{2}I.$ (157)
This condition ensures that the effect of input perturbation noise on the closed-loop system does not exceed the magnitude of the process noise.
A sufficient condition for (157) can be written as
$\frac{2\,\kappa_{t}^{2}\,\bar{\vartheta}_{B_{*}}^{2}\bar{p}_{t}}{\sqrt{t+\bar{c}}}\leq 1$ (158)
where we define $\bar{\vartheta}_{B_{*}}=\max\{1,\vartheta_{B_{*}}\}$ and $\|B_{*}\|\leq\vartheta_{B_{*}}$.
To derive a suitable upper bound for $\bar{p}_{t}$, we first establish the following inequalities.
We have
$\log\frac{n\det V_{t}}{\delta\,\det(\lambda I)}$ $\leq(n+m)\log\frac{n}{\delta}\Big(1+t\log\frac{t}{\delta}\Big)$ (159)
$\leq 2(n+m)\log\frac{2nt}{\delta}.$ (160)
where (159) holds under the choice
$\lambda\geq\frac{\sum_{k=1}^{z}\|z_{k}z_{k}^{\top}\|^{2}}{t\log\frac{2t}{\delta}}.$ (161)
By choosing $\epsilon$ appropriately such that
$\sqrt{\lambda}\epsilon\leq\sigma_{\omega}\sqrt{2n(n+m)}$ (162)
it then follows from (160) and (162) that
$r_{t}\leq 16\sigma_{\omega}^{2}n(n+m)\log\frac{2nt}{\delta}.$
Furthermore, from Lemma 25, we have
$\|z_{t}\|^{2}\leq c^{2}_{z}(\kappa_{*},\bar{\rho})\log\frac{2t}{\delta}.$
To obtain a computable value for $\bar{c}$, we require an upper bound on $\kappa_{*}$, which can be derived directly from Lemma 24 and, in particular, from (172):
$\kappa_{t}^{2}\leq\kappa_{*}^{2}\leq\kappa_{t}^{2}+4\kappa_{t}^{6-\bar{\rho}},\quad\text{for all }t.$ (163)
From this, we can define
$\kappa_{*}\leq\bar{\kappa}_{1}(\bar{\rho}):=\sqrt{\kappa_{1}^{2}+4\kappa_{1}^{6-\bar{\rho}}}.$
Combining these results, and noting that $\kappa_{t}\leq\bar{\kappa}_{1}$ as guaranteed by (163), we can now derive and upper-bound for the left hand side of (158) as follows
$\frac{1280\bar{\vartheta}_{B_{*}}^{2}\kappa_{t}^{\bar{\rho}+4}\vartheta\sqrt{r_{t}}}{\sigma_{\omega}^{2}(t+\bar{c})}\big(\lambda+c_{z}^{2}(\kappa_{*},\bar{\rho})\log\frac{2t}{\delta})\big)^{1/2}\leq$ $\frac{5120\bar{\vartheta}_{B_{*}}^{2}\bar{\kappa}_{1}^{\bar{\rho}+4}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)\log\frac{2nt}{\delta}}}{\sigma_{\omega}^{2}(t+\bar{c})}\sqrt{\lambda}$ (164)
$+\frac{5120\bar{\vartheta}_{B_{*}}^{2}\bar{\kappa}_{1}^{\bar{\rho}+4}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}}{\sigma_{\omega}^{2}(t+\bar{c})}c_{z}(\bar{\kappa}_{1},\bar{\rho})\log\frac{2nt}{\delta}.$ (165)
We enforce the following condition:
$\frac{5120\bar{\vartheta}_{B_{*}}^{2}\bar{\kappa}_{1}^{\bar{\rho}+4}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}c_{z}(\bar{\kappa}_{1},\bar{\rho})\log\frac{2nt}{\delta}\leq\frac{1}{2}.$ (166)
Applying Lemma 4, an appropriate choice of $\bar{c}$ that satisfies this condition is
$\bar{c}=a^{2}\left(2\log\!\left(\frac{4na^{2}}{\delta}\right)+1\right)^{2},$ (167)
where
$a=\frac{10240\bar{\vartheta}_{B_{*}}^{2}\bar{\kappa}_{1}^{\bar{\rho}+4}\vartheta\sqrt{\sigma_{\omega}^{2}n(n+m)}c_{z}(\bar{\kappa}_{1},\bar{\rho})}{\sigma_{\omega}^{2}}.$
With this choice of $\bar{c}$, condition (166) ensures that (165) is satisfied since $t+\bar{c}>1$. Moreover, by choosing $\lambda=\bar{C}=\frac{\sigma_{\omega}^{2}\bar{c}}{40}$, we also satisfy (164).
Recalling (161), which provides an additional condition for tuning $\lambda$ based on the upper bound of the co-state, we require
$\lambda\geq c_{z}^{2}(\bar{\kappa}_{1},\bar{\rho}).$
Therefore, the natural choice is to select
$\lambda\geq\max\left\{\frac{\sigma_{\omega}^{2}\bar{c}}{80},\,c_{z}^{2}(\bar{\kappa}_{1},\bar{\rho})\right\}=\frac{\sigma_{\omega}^{2}\bar{c}}{80},$ (168)
where the equality holds due to the magnitude of $\bar{c}$ in (167).
However, we slightly increase it to ensure closed-loop stability, as will be discussed in the proof of Theorem 4. We choose
$\lambda=\frac{\sigma_{\omega}^{2}\bar{c}}{80}(n+m)\log\frac{1}{\delta},$ (169)
With $\bar{c}$ specified, $\lambda$ is determined. Then, from (156), $\bar{p}_{t}$ is obtained as
$\bar{p}_{t}:=\frac{640\kappa_{t}^{\bar{\rho}+2}\vartheta\sqrt{r_{t}}(\lambda+\|\sum_{k=1}^{t}z_{k}z_{k}^{\top}\|)^{1/2}}{\sigma_{\omega}^{2}\sqrt{t+\bar{c}}}.$ (170)
Recalling (162), the appropriate choice of $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ is given by
$\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho}):=\frac{\sigma_{\omega}\sqrt{2n(n+m)}}{\sqrt{\lambda}}.$ (171)
With $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ defined in (171), and with $\lambda$ and $\bar{p}_{t}$ selected according to (169) and (170), respectively, the condition (150) is satisfied. Moreover, by following the same reasoning as in Lemma 14, it can be concluded that any generated policy is $(\kappa_{*},\gamma_{*})$-strongly stabilizing, where
$\kappa_{*}=\sqrt{\frac{2\|P(\mathcal{C}_{t})\|}{\alpha_{0}}},\quad\gamma_{*}=\frac{1}{2\kappa_{*}^{2}}.$
Lemma 24
[…]
$P(\mathcal{C}_{t})\preceq P_{*}\preceq\;P(\mathcal{C}_{t+1})+2\kappa_{t+1}^{6-\bar{\rho}}\alpha_{0}I\preceq\;P(\mathcal{C}_{t+1})+2\kappa_{*}^{6-\bar{\rho}}\alpha_{0}I,$ (172)
and
$\big\|H_{t+1}^{-1}H_{t}\big\|\;\leq\;1+2{\kappa_{*}^{3-\bar{\rho}/2}}\;=:\;{1+\xi(\kappa_{*},\bar{\rho}).}$ (173)
where $H_{k}=P^{1/2}(\mathcal{C}_{k})$.
Proof
We begin by proving the first claim. In Lemma 15, we have already established
$P(\mathcal{C}_{t})\preceq P_{*}$
and we have also shown that
$P_{*}-P(\mathcal{C}_{t})$ $\preceq\frac{16\|P(\mathcal{C}_{t})\|^{2}\mu_{t}}{\alpha_{0}^{2}}\|P(\mathcal{C}_{t})\|_{*}\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I$
$=4\kappa_{t}^{4}\mu_{t}\|P(\mathcal{C}_{t})\|_{*}\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I.$ (174)
Moreover, since the parameters $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$, $\lambda$, and $\bar{p}_{t}$ have been selected such that
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}\|V_{t}^{-1}\|I\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I$ (175)
which indeed is the sufficient condition for
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I$
which fulfills generation of $(\kappa_{*},\gamma_{*})$-strongly stabilizing policies for any $\bar{\rho}\in[0,\;8)$. We also have:
$\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|\leq 2\kappa_{t}^{2}.$ (176)
Substituting (175) and (176) into (174) yields
$P_{*}\preceq\;P(\mathcal{C}_{t})+2\kappa_{t}^{6-\bar{\rho}}\alpha_{0}I\;\preceq\;P(\mathcal{C}_{t})+2\kappa_{*}^{6-\bar{\rho}}\alpha_{0}I$ (177)
where the last inequality follows from $\kappa_{t}\leq\kappa_{*}$.
Applying similar steps as in (131) on (177) implies
$\|H_{t+1}^{-1}H_{t}\|^{2}\leq 1+4{\kappa_{*}^{6-\bar{\rho}}}$
which results in:
$\|H_{t+1}^{-1}H_{t}\|\leq\sqrt{1+4{\kappa^{6-\bar{\rho}}}}\leq 1+2{\kappa_{*}^{3-\bar{\rho}/2}}=:1+\xi(\kappa_{*},\bar{\rho}).$
Remark 4
The condition in (173) holds between any two policies because the upper and lower bounds on $P_{*}$ in (172) are valid for arbitrary time steps. This matrix inequality is, in fact, used to derive (173). Therefore, for any pair of time steps $(t^{\prime},t)$, we have
$\|H_{t^{\prime}}^{-1}H_{t}\|\leq 1+\xi(\kappa_{*},\bar{\rho}).$
J.2 Proof of Theorem 4
Proof Let $M=A_{*}+B_{*}K_{t}$ where $K_{t}^{\prime}$s are $(\kappa,\gamma)$- strong sequential stabilizing policies generated by ARSLO algorithm. The closed loop dynamics then is written as follows:
$x_{t}=M_{1}x_{0}+\sum_{s=1}^{t-1}M_{s+1}(B_{*}\nu_{s}+\omega_{s+1})$
where
$M_{t}=I,\;\;M_{s}=M_{s+1}(A_{*}+B_{*}K_{s})=\prod_{j=s}^{t-1}(A_{*}+B_{*}K_{j}),\;\forall 1\leq s\leq t-1.$
Recalling Definition 1, we can set $A_{*}+B_{*}K_{j}=H_{j}L_{j}H_{j}^{-1}$, then for all $1\leq s<t$ where $\|L_{j}\|\leq 1-\gamma_{j}$ one can write
$\|M_{s}\|\leq$ $\lVert(H_{t-1}L_{t-1}H^{-1}_{t-1})...(H_{s+1}L_{s+1}H^{-1}_{s+1})(H_{s}L_{s}H^{-1}_{s})\rVert$
$\leq$ $\|H_{t-1}\|\;\|(L_{t-1}H^{-1}_{t-1}H_{t-2})...(L_{s+2}H^{-1}_{s+2}H_{s+1})(L_{s+1}H^{-1}_{s+1}H_{s})\|\>\|L_{s}\|\>\|H_{s}\|^{-1}$
$\leq$ $\|H_{t-1}\|\bigg(\prod_{j=s+1}^{t-1}\|L_{j}\|\bigg)\bigg(\prod_{j=s+1}^{t-1}\|H_{j}^{-1}H_{j-1}\|\bigg)(1-\gamma_{s})\|H_{s}^{-1}\|$
$\leq$ $\sqrt{P(\mathcal{C}_{t-1})}\bigg(\prod_{j=s+1}^{t-1}(1-\gamma_{j})\bigg)\bigg(\prod_{j=s+1}^{t-1}\|H_{j}^{-1}H_{j-1}\|\bigg)(1-\gamma_{s})\sqrt{\frac{2}{\alpha_{0}}}$
$\leq$ $\kappa_{{t-1}}\bigg(\prod_{j=s+1}^{t-1}(1-\gamma_{j})\;\bigg)\bigg(\prod_{j=s+1}^{t-1}\|H_{j}^{-1}H_{j-1}\|\bigg)(1-\gamma_{s})$
$\leq$ ${\kappa_{*}}(1-{\gamma_{*}})^{t-s}\bigg(\prod_{j=s+1}^{t-1}\|H_{j}^{-1}H_{j-1}\|\bigg)$
$\leq$ ${\kappa_{*}}(1-{\gamma_{*}})^{t-s}\big(1+\xi(\kappa_{*},\bar{\rho})\big)^{N(t-1)-N(s+1)}$
$\leq$ ${\kappa_{*}}(1-{\gamma_{*}})^{t-s}\big(1+\xi(\kappa_{*},\bar{\rho})\big)^{N(t-s-1)}$
where we applied
[…]
$\kappa_{t-1}=\sqrt{2P(\mathcal{C}_{t-1})/\alpha_{0}}$ by definition, and that $\kappa_{t}\leq\kappa_{*}$ for any $t$. Also, we use the fact that $N(t)-N(s)\leq N(t-s)$ which follows from the monotonicity of epoch lengths. This monotonicity is a direct consequence of the determinant growth identity for positive definite matrices:
$\det(V_{k+1})=\det(V_{k})\big(I+z_{k}^{\top}V^{-1}_{k}z_{k}\big).$
Since $V_{k}\succeq V_{k-1}$, thus, the incremental growth of $\det(V_{k})$ decreases over time. Consequently, to satisfy the update condition $\det(V_{k})\;\geq\;\bigl(1+\beta(\bar{\rho})\bigr)\,\det(V_{\tau(k)})$ a longer interval between updates is required. In other words, epoch lengths are nondecreasing, which is an established fact for OFU-based algorithms with determinant-type policy update rules. This implies that update events become less frequent as time increases. Hence, for any interval $[s,t)$, the number of updates within this interval is at most the number of updates that could occur in an initial interval of the same length, establishing $N(t)-N(s)\;\leq\;N(t-s)$.
Given the policy update rule, for some tunable $\beta_{\tau(t)}(\bar{\rho})$, we aim to upper-bound the number of policy updates by time $s$. Similar to the analysis in Lemma 18, the number of policy updates by time $s$, denoted $N(s)$, can be upper-bounded as
$N(s)\leq\log_{1+\beta(\bar{\rho})}\bigg(\frac{\lambda+\|\sum_{k=1}^{s}z_{k}z_{k}^{\top}\|}{\lambda}\bigg)^{n+m}.$ (178)
The only difference compared to Lemma 18 is that $\beta_{\tau(k)}(\bar{\rho})$ is not constant; hence we use its minimum over all epochs to upper-bound $N(s)$.
Note that by Lemma 24, we have $\xi(\kappa_{*},\bar{\rho})=2\kappa_{*}^{3-\bar{\rho}/2}$ and moreover
$\kappa_{t}^{2}\leq\kappa_{*}^{2}\leq\kappa_{t}^{2}+4\kappa_{t}^{6-\bar{\rho}},\quad\text{for all }t$
and then
$\kappa_{1}^{2}\leq\kappa_{*}^{2}\leq\kappa_{1}^{2}+4\kappa_{1}^{6-\bar{\rho}}.$
Using this, we choose
$\beta(\bar{\rho})=\begin{cases}2(\kappa_{1}^{2}+\kappa_{1}^{(6-\bar{\rho})})^{\frac{(6-\bar{\rho})}{4}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8\end{cases}$ (179)
which guarantees that
$1+\beta(\bar{\rho})\geq 1+\xi(\kappa_{*},\bar{\rho}).$ (180)
Note that for $6\leq\bar{\rho}<8$, we have $\xi(\kappa_{*},\bar{\rho})\leq 2$. This is why, over this specific range, we choose $\beta_{\tau(k)}(\bar{\rho})=2$, which allows us to control the growth of the state norm, as shown below. With this choice, we obtain
$(1-\gamma_{*})^{\frac{s}{2}}(1+\xi(\kappa_{*},\bar{\rho}))^{N(s)}\leq$ $(1-\gamma_{*})^{\frac{s}{2}}\bigg(\frac{\lambda+\|\sum_{k=1}^{s}z_{k}z_{k}^{\top}\|}{\lambda}\bigg)^{n+m}$
where we used the property $a^{\log_{b}c}\leq c$ for $b\geq a>1$.
Using the standard bound $\big\|\sum_{k=1}^{s}z_{k}z_{k}^{\top}\big\|\leq c_{z}^{2}(\kappa_{*},\bar{\rho})\,t\log\frac{t}{\delta}$, for $\delta\leq 1/e$ we can further upper-bound
$\bigg(\frac{\lambda+\|\sum_{k=1}^{s}z_{k}z_{k}^{\top}\|}{\lambda}\bigg)^{n+m}$ $\leq\bigg(1+\frac{c^{2}_{z}(\kappa_{*},\bar{\rho})s\log\frac{s}{\delta}}{\lambda}\bigg)^{n+m}$
$\leq\bigg(1+\frac{c^{2}_{z}(\kappa_{*},\bar{\rho})s^{2}\log\frac{1}{\delta}}{\lambda}\bigg)^{n+m}$
$\leq\bigg(1+\frac{c_{z}(\kappa_{*},\bar{\rho})s\sqrt{\log\frac{1}{\delta}}}{\sqrt{\lambda}}\bigg)^{2(n+m)}.$
The second inequality follows from the property that $\log ab\leq a\log b$ for $a\geq 1$ and $b\geq e$. The third inequality uses the fact that $(1+x^{2})\leq(1+x)^{2}$ for any $x\geq 0$.
Recall from (169) that we chose
$\lambda=\frac{\sigma_{\omega}^{2}\bar{c}}{40}(n+m)\log\frac{1}{\delta}$
Considering the definition of $\bar{c}$ in (167), it is straightforward to verify that
$\lambda\geq 64(n+m)^{2}\bar{\kappa}_{1}^{8}c^{2}_{z}(\bar{\kappa}_{1}^{2},\bar{\rho})\log\frac{1}{\delta}.$
This then gives
$\bigg(1+\frac{c_{z}(\kappa_{*},\bar{\rho})s\sqrt{\log\frac{1}{\delta}}}{\sqrt{\lambda}}\bigg)^{2(n+m)}\leq\bigg(1+\frac{s}{8(n+m)\bar{\kappa}_{1}^{4}}\bigg)^{2(n+m)}\leq e^{\frac{s}{4\bar{\kappa}_{1}^{4}}}$
where the last inequality follows from the standard property $(1+\frac{x}{a})^{a}\leq e^{x/a}$ for any $x\geq 0$.
Then we have
$(1-\gamma_{*})^{\frac{s}{2}}(1+\xi(\kappa_{*},\bar{\rho}))^{N(s)}\leq$ $(1-\gamma_{*})^{\frac{s}{2}}e^{\frac{s}{4\bar{\kappa}_{1}^{4}}}$
$=$ $(1-\frac{1}{2\kappa_{*}^{2}})^{\frac{s}{2}}e^{\frac{s}{4\kappa_{*}^{4}}}$
$\leq$ $e^{\frac{-s}{4\kappa_{*}^{2}}}e^{\frac{s}{4\kappa_{*}^{4}}}=e^{-(\frac{1}{4\kappa_{*}^{2}}-\frac{1}{4\kappa_{*}^{4}})s}.$
Hence, we can upper bound
$\|M_{1}\|\leq\kappa_{*}(1-\gamma_{*})^{\frac{t-s}{2}}$
and
$\|M_{s}\|\leq\kappa_{*}(1+\xi(\kappa_{*},\bar{\rho}))(1-\gamma_{*})^{\frac{t-s}{2}}.$
Using this bound, the upper bound on the state norm becomes
$\|x_{t}\|\leq$ $\|M_{1}\|\|x_{1}\|+\sum_{s=1}^{t-1}\|M_{s+1}\|\|B_{*}\eta_{j}+\omega_{s}\|$
$\leq$ $\kappa_{*}(1-\gamma_{*})^{(t-1)/2}\|x_{1}\|+\kappa_{*}(1+\xi(\kappa_{*},\bar{\rho}))\sum_{s=1}^{t-1}(1-{\gamma_{*}})^{(t-s-1)/2}\|B_{*}\eta_{s}+\omega_{s}\|$
$\leq$ $\kappa_{*}e^{-{\gamma_{*}(t-1)}/{2}}\,\|x_{1}\|+\kappa_{*}(1+\xi(\kappa_{*},\bar{\rho}))\max_{1\leq s<t}\|B_{*}\eta_{s}+\omega_{s}\|\sum_{t=1}^{\infty}(1-{\gamma_{*}})^{t/2}$
$\leq$ $\kappa_{*}e^{-{\gamma_{*}(t-1)}/{2}}\,\|x_{1}\|+\frac{2\kappa_{*}}{\gamma_{*}}(1+\xi(\kappa_{*},\bar{\rho}))\max_{1\leq s<t}\|B_{*}\eta_{s}+\omega_{s}\|$
Now, with the choice of $\bar{c}$ in (167), which guarantees
$B_{*}\frac{\bar{p}_{t}\sigma_{\omega}^{2}\big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|I}{\alpha_{0}}\big)}{\sqrt{t+\bar{c}}}B_{*}^{\top}\preceq\sigma_{\omega}^{2}I$
and applying the Hanson–Wright concentration inequality, we obtain
$\max_{1\leq s\leq t}\|B_{*}\eta_{s}+\omega_{s}\|\leq\sqrt{20\sigma_{\omega}^{2}n\log\frac{t}{\delta}}.$
Consequently,
$\|x_{t}\|\leq\kappa_{*}e^{-{\gamma_{*}(t-1)}/{2}}\,\|x_{1}\|+\frac{2\kappa_{*}}{\gamma_{*}}(1+\xi(\kappa_{*},\bar{\rho}))\sqrt{20n\sigma_{\omega}^{2}\log\frac{t}{\delta}}.$ (181)
The following lemma gives appropriate upper-bounds for $\|x_{t}\|^{2}$ and $\|z_{t}\|^{2}$.
Lemma 25
The following statements hold with probability at least $1-\delta$:
$\|x_{t}\|^{2}\leq 4\frac{\kappa_{*}^{2}}{\gamma_{*}^{2}}(1+\xi(\kappa_{*},\bar{\rho}))^{2}\big(\frac{\|x_{1}\|}{\sqrt{\log\frac{1}{\delta}}}+\sqrt{20n\sigma_{\omega}^{2}}\big)^{2}\log\frac{t}{\delta}=:X_{t}^{2}(\kappa_{*},\bar{\rho})$ (182)
and
$\|z_{t}\|^{2}\leq{64\kappa_{*}^{6}(1+\kappa_{*}^{2})}(1+\xi(\kappa_{*},\bar{\rho}))^{2}\bigg(\frac{\|x_{1}\|}{\log\frac{1}{\delta}}+\sqrt{20n\sigma_{\omega}^{2}}\bigg)^{2}\log\frac{t}{\delta}=:c_{z}^{2}(\kappa_{*},\bar{\rho})\log\frac{t}{\delta}=:\zeta_{t}^{2}(\kappa_{*},\bar{\rho})$ (183)
Proof The proof proceeds exactly as in Lemma 17.
J.3 Proof of Corollary 2
Proof The proof follows from the definition of $\lambda$, noting that by (32), $\lambda$ has the same order as $a^{2}$, namely $\mathcal{O}(\bar{\kappa}_{1}^{\,22+\bar{\rho}})$. Moreover, by the definition of $\bar{\kappa}_{1}$ in (35), it is of order $\mathcal{O}(\max\{\kappa_{1},\;\kappa_{1}^{(6-\bar{\rho})/2}\})$, and $\kappa_{1}$ itself is upper-bounded by $\kappa_{*}$. Finally, since $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ defined in (38) is of order $\mathcal{O}(1/\sqrt{\lambda})$, the claim follows.
[…]
K.1 Proof of Theorem 5
[…]
Lemma 26
With probability at least $1-\delta$, the following upper bound holds for $|R_{1}(t)|$:
$|R_{1}(t)|\lesssim\|P_{*}\|^{7-\frac{\bar{\rho}}{2}}n(n+m)\log^{2}\frac{t}{\delta}$
Proof The proof follows the same structure as Lemma 18, with the only difference being that the policy update rule now satisfies $\det(V_{t})\geq(1+\beta_{\tau(k)}(\bar{\rho}))\det(V_{\tau(k)})$ where $\beta_{\tau(k)}(\bar{\rho})$ is defined in (179). The only remaining technicality is to upper-bound the number of policy updates $N(t)$, which we have already established in (178):
$N(s)\leq\log_{1+\beta(\bar{\rho})}\bigg(\frac{\lambda+\|\sum_{k=1}^{s}z_{k}z_{k}^{\top}\|}{\lambda}\bigg)^{n+m}.$
Moreover, by (179), for any $\bar{\rho}\in[0,8)$ we have the property (180)
$1+\beta(\bar{\rho})\geq 3\quad\forall k.$
which implies
$N(s)$ $\leq\log_{3}\bigg(\frac{\lambda+\|\sum_{k=1}^{s}z_{k}z_{k}^{\top}\|}{\lambda}\bigg)^{n+m}$
$\leq 4(n+m)\log_{1+\xi(\kappa_{*},\bar{\rho})}\frac{t}{\delta}$
where the last inequality holds by our choice of $\lambda$ in (168), which satisfies $\lambda\geq c_{z}^{2}(\kappa_{,}\bar{\rho})$.
Thus, similar to the ARSLO algorithm case,
$|R_{1}(t)|\leq$ $\|P_{*}\|X_{t}^{2}(\kappa_{*},\bar{\rho})+4\|P_{*}\|(n+m)X_{t}^{2}(\kappa_{*},\bar{\rho})\log_{3}\frac{t}{\delta}.$
Using the upper bound on $X_{t}^{2}(\kappa_{*},\bar{\rho})$ given in (182), and the fact that $\xi(\kappa_{*},\bar{\rho})\lesssim\|P_{*}\|^{(3-\bar{\rho}/2)/2}$ by (173) and the definition of $\kappa_{*}$, we obtain
$X^{2}_{t}(\kappa_{*},\bar{\rho})\lesssim\|P_{*}\|^{6-\frac{\bar{\rho}}{2}}n\log\frac{t}{\delta}.$ (184)
Therefore,
$|R_{1}(t)|\lesssim\|P_{*}\|^{7-\frac{\bar{\rho}}{2}}n(n+m)\log^{2}\frac{t}{\delta}.$
Lemma 27
The following upper bound holds for $|R_{2}(t)|$:
$|R_{2}(t)|\lesssim\|P_{*}\|^{4-\frac{\bar{\rho}}{4}}\sqrt{n\log^{2}\frac{t}{\delta}}$
with probability at least $1-\delta/4$.
Proof The argument follows the same steps as in Lemma 19. In particular, we have
$|R_{2}(t)|\leq 2\sigma_{\omega}\|P_{*}\|X_{t}(\kappa_{*},\bar{\rho})\sqrt{t\log\frac{2}{\delta}}.$
Applying the bound on $X_{t}(\kappa_{*},\bar{\rho})$ given in (184) yields the stated result.
Lemma 28
The term $R_{3}(t)$ admits the following bound:
$R_{3}(3)\lesssim\|P_{*}\|_{*}\sqrt{t\log^{3}\frac{t}{\delta}}$
with probability at least $1-\delta$.
Proof The proof follows directly from Lemma 20.
Lemma 29
Let $\bar{\rho}\in[0,\,8)$. Then the term $R_{4}(t)$ admits the upper bound
$R_{4}(t)\lesssim\,\vartheta\,\beta_{*}(\|P_{*}\|,\bar{\rho})\,\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{9-\bar{\rho}/2}\,t\,\log^{4}\frac{nt}{\delta}}$
with probability at least $1-\delta$, where
$\beta_{*}(\|P_{*}\|,\bar{\rho})=\begin{cases}\big(\|P_{*}\|+\|P_{*}\|^{(6-\bar{\rho})/2}\big)^{\frac{(6-\bar{\rho})}{4}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8\end{cases}.$
Proof The argument follows the same steps as in Lemma 21, with the main difference being that $\beta$ is no longer equal to $1$ and is instead tuned according to (179). Its upper bound satisfies
$\beta(\bar{\rho})\lesssim\beta_{*}(\|P_{*}\|,\bar{\rho})=\begin{cases}\big(\|P_{*}\|+\|P_{*}\|^{(6-\bar{\rho})/2}\big)^{\frac{(6-\bar{\rho})}{4}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8\end{cases}.$
With this adjustment, then
$R_{4}(t):=$ $4\sum_{k=1}^{t}\bigg(1+\beta_{*}(\|P_{*}\|,\bar{\rho})\bigg)\,\|P(\mathcal{C}_{k})\|\mu_{\tau(k)}\big({z}^{\top}_{k}V_{k}^{-1}{z}_{k}\big)1_{\mathcal{E}_{k}`}.$ (185)
Following the same steps in bounding $\mu_{t}$ from Lemma 21, and using the fact that
$c_{z}^{2}(\kappa_{*},\bar{\rho})\lesssim\,n\|P_{*}\|^{7-\frac{\bar{\rho}}{2}}.$ (186)
Thus,
$\mu_{\tau(k)}\leq\mu_{k}\lesssim$ $\vartheta\,\sqrt{r_{k}}\,\sqrt{\zeta_{k}(\kappa_{*},\bar{\rho})\,k}$
$\lesssim$ $\vartheta\sqrt{n(n+m)\frac{2nt}{\delta}}\sqrt{n\,k\,\|P_{*}\|^{7-\frac{\bar{\rho}}{2}}\log\frac{k}{\delta}}$
$\lesssim$ $\vartheta\sqrt{n^{2}(n+m)\|P_{*}\|^{7-\bar{\rho}/2}\,k\,\log^{2}\frac{2nk}{\delta}}.$
As in Lemma 21, we also use
$\sum_{t=1}^{t-1}z_{t}^{\top}V_{t}^{-1}z_{t}$ $\leq 2\log\frac{\det(V_{t})}{\det(\lambda I)}\lesssim(n+m)\log\frac{nt}{\delta}.$
Substituting all bounds into (185) yields:
$R_{4}(t)\lesssim\,\vartheta\,\beta_{*}(\|P_{*}\|,\bar{\rho})\,\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{9-\bar{\rho}/2}\,t\,\log^{4}\frac{nt}{\delta}}$
which completes the proof.
Lemma 30
The term $R_{5}(t)$ admits the following upper bound:
$R_{5}(t)\lesssim\alpha_{1}(n^{4}(n+m))^{\frac{1}{4}}\sqrt{\vartheta\|P_{*}\|^{(50-\bar{\rho})/4}\log^{3}\frac{nt}{\delta}}\,t^{1/4}.$
with probability at least $1-\delta$.
Proof The proof follows the same steps as in Lemma 22. Similar to (144), we may write
$R_{5}\leq 4\sqrt{2\bar{p}_{t}\log\frac{4}{\delta}}\alpha_{1}\sigma_{\omega}\kappa_{*}^{2}X_{t}(\kappa_{*},\bar{\rho})(t+\bar{c})^{1/4}.$ (187)
Using the definition of $\bar{p}_{t}$ in (170) and the fact that $r_{t}\lesssim n(n+m)\log\tfrac{nt}{\delta}$, together with the bound in (186), we obtain
$\bar{p}_{t}\lesssim\vartheta\sqrt{n^{2}(n+m)\|P_{*}\|^{(9+\bar{\rho}/2)}\,\log^{2}\frac{nt}{\delta}}.$ (188)
Moreover, we already derived the bound on $X_{t}^{2}(\kappa_{*},\bar{\rho})$ in (184):
$X^{2}_{t}(\kappa_{*},\bar{\rho})\lesssim\|P_{*}\|^{6-\frac{\bar{\rho}}{2}}n\log\frac{t}{\delta}.$ (189)
Substituting these expressions into (187) completes the proof and yields
$R_{5}(t)\lesssim\alpha_{1}(n^{4}(n+m))^{1/4}\sqrt{\vartheta\|P_{*}\|^{(50-\bar{\rho})/4}\log^{3}\frac{nt}{\delta}}t^{1/4}.$
Lemma 31
The term $R_{6}(t)$ has the following upper bound.
$R_{6}(t)\lesssim\alpha_{1}\vartheta\sqrt{n^{2}(n+m)m^{2}\|P_{*}\|^{(11+\bar{\rho}/2)}t\log^{4}\frac{nt}{\delta}}$
with probability at least $1-\delta$.
Proof The following inequality holds
$R_{6}(t)=$ $\sum_{k=1}^{t}\eta^{\top}_{k}\big(R+4\mu_{\tau(k)}\|P(\mathcal{C}_{k})\|V^{-1}_{\tau(k)}\big)\eta_{k}1_{\mathcal{E}_{k}}$
$\leq$ $\sum_{k=1}^{t}\eta^{\top}_{k}\big(R+\alpha_{0}I\big)\eta_{k}1_{\mathcal{E}_{k}}$
$\leq$ $2\alpha_{1}\sum_{k=1}^{t}\eta^{\top}_{k}\eta_{k}1_{\mathcal{E}_{k}}.$
The second inequality follows from (118), namely
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I\quad\forall t$
which is ensured by appropriate tuning of $\lambda$, $\bar{p}_{t}$, and $\bar{\epsilon}$, as followed by the proof of Theorem 3.
The proof follows the same steps of teh counter part lemma for ARSLO algorithm, i.e., Lemma 23. By applying Hanson–Wright inequality (Lemma 6), it yields
$\sum_{k=1}^{t}z_{k}^{\top}\Gamma_{k}z_{k}$ $\leq 5\sum_{k=1}^{t}\|\Gamma_{k}\|_{*}\log\frac{k}{\delta}\leq 10m\|\Gamma_{t}\|\sigma_{\omega}^{2}\log\frac{t}{\delta}\kappa_{*}^{2}\sum_{k=1}^{t}\frac{1}{\sqrt{k+\bar{c}}}$
$\leq 20\,m\,\bar{p}_{t}\,\sigma_{\omega}^{2}\kappa_{*}^{2}\,\log\frac{t}{\delta}\big(\sqrt{t+\bar{c}}-\sqrt{\bar{c}}\big)$
in which we applied the fact that $\|\Gamma_{k}\|_{*}\leq m\|\Gamma_{k}\|$.
Finally by upper-bound of $\bar{p}_{t}$ given by (188) we conclude that
$R_{6}(t)\lesssim\alpha_{1}\vartheta\sqrt{n^{2}(n+m)m^{2}\|P_{*}\|^{(11+\bar{\rho}/2)}t\log^{4}\frac{nt}{\delta}}.$
To complete the proof, it suffices to observe that the terms $R_{4}(t)$ and $R_{6}(t)$ are dominant, and depending on the value of $\bar{\rho}$, either term may dominate individually. Therefore, for any $\bar{\rho}\in[0,\,8)$, the regret upper bound can be expressed as
$R_{\text{ARSLO}^{+}(\bar{\rho})}(t)\leq$ $\{\mathcal{O}\Big(\sqrt{\max\bigg\{n^{2}(n+m)^{3}\|P_{*}\|^{9-\frac{\bar{\rho}}{2}}\,\beta_{*}^{2}\big(\|P_{*}\|,\bar{\rho}\big)\,\;,n^{2}(n+m)m^{2}\|P_{*}\|^{(11+\frac{\bar{\rho}}{2})}\bigg\}\,t\;\log^{4}\frac{t}{\delta}}\,\Big)$
where
$\beta_{*}\big(\|P_{*}\|,\bar{\rho}\big)=\begin{cases}\big(\|P_{*}\|+\|P_{*}\|^{(6-\bar{\rho})/2}\big)^{\frac{(6-\bar{\rho})}{4}}&\text{if }\bar{\rho}<6,\\ 2&6\leq\bar{\rho}<8.\end{cases}$
To optimize the regret bound, it suffices to choose $\bar{\rho}$ to balance the dominant terms in their dependence on $\|P_{*}\|$. Setting $\bar{\rho}=\bar{\rho}_{*}=2$ achieves this balance. In terms of dimensional dependence, the dominant term is $n^{2}(n+m)^{3}$, which yields
$R_{\text{ARSLO}^{+}(\bar{\rho}_{*})}(t)\leq$ $\mathcal{O}\Big(\sqrt{\,n^{2}(n+m)^{3}\;\|P_{*}\|^{12}\;t\;\log^{4}\frac{t}{\delta}}\,\Big).$
that completes the proof.
Appendix L Supporting analysis of warm-up phase
L.1 Proof of Theorem 6
Proof Recall that $\Theta_{*}$ belongs to the confidence ellipsoid $\mathcal{C}_{t}^{w}(\delta)$ with probability $1-\delta$, which implies
$\operatorname{Tr}\Big((\Theta_{*}-\hat{\Theta}^{w}_{t})^{\top}\bar{V}_{t}(\Theta_{*}-\hat{\Theta}^{w}_{t})\Big)\leq r^{w}_{t}.$
It then follows that
$\|\Theta_{*}-\hat{\Theta}^{w}_{t}\|_{F}^{2}\leq\frac{r_{t}^{w}}{\lambda_{\min}(\bar{V}_{t})}.$
Hence, a sufficient condition to guarantee
$\|\Theta_{*}-\hat{\Theta}^{w}_{t}\|_{F}\leq\varepsilon$
where $\varepsilon$ can be either $\bar{\epsilon}(\bar{\kappa}_{1})$ or $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ is
$\sqrt{\frac{r_{t}^{w}}{\lambda_{\min}(\bar{V}_{t})}}\leq\varepsilon,$ (190)
which completes the proof.
Before providing the proof of Theorem 7, we need to specify an upper bound on the state norm during warm-up phase.
Lemma 32
Consider the linear system $(A_{*},B_{*})$ with initial state $x_{0}=0$, and let $K_{0}$ be a $(\kappa_{0},\gamma_{0})$-stabilizing policy. Then, during the warm-up phase, the state satisfies
$\|x_{t}\|\;\leq\;\sigma_{\omega}\,\sqrt{10\Bigl(n+m\kappa_{0}^{2}\bar{\vartheta}_{B_{*}}^{2}\Bigr)\log\frac{t}{\delta}}$ (191)
and
$\|z_{t}\|\leq\frac{12\sigma_{\omega}\kappa_{0}^{2}}{\gamma_{0}}\sqrt{\Bigl(n+m\kappa_{0}^{2}\bar{\vartheta}_{B_{*}}^{2}\Bigr)\log\frac{4t_{w}}{\delta}}.$ (192)
with probability at least $1-\delta$, for any $\delta\in(0,1/e)$.
Proof The closed-loop dynamics at any time $s$ can be written as
$x_{s+1}=(A_{*}+B_{*}K_{0})x_{s}+B_{*}\nu_{s}+\omega_{s}.$
Unrolling this recursion to time $t$ gives
$x_{t}=M_{1}x_{1}+\sum_{s=1}^{t-1}M_{s}(B_{*}\nu_{s}+\omega_{s+1})$
where
$M_{t}=I,\;\;M_{s}=M_{s+1}(A_{*}+B_{*}K_{0})=\prod_{j=s}^{t-1}(A_{*}+B_{*}K_{0}),\;\forall 1\leq s\leq t-1.$
From strong $(\kappa_{0},\gamma_{0})$- stability of $K_{0}$ we have decomposition $A_{*}+B_{*}K_{0}=HLH^{-1}$ with $\|L\|\leq(1-\gamma_{0})$ and $\|H\|\|H^{-1}\|\leq\kappa_{0}$. This implies
$\|M_{s}\|\leq$ $\|H\|\bigg(\prod_{j=s}^{t-1}\|L\|\;\bigg)\|H^{-1}\|\leq{\kappa_{0}}(1-\gamma_{0})^{t-s}.$
Therefore,
$\|x_{t}\|$ $\leq\kappa_{0}(1-\gamma_{0})^{t}x_{0}+\max_{0\leq s\leq t-1}\|(B_{*}\nu_{t-1-s}+\omega_{t-1-s})\|\|\sum_{s=0}^{t-1}M_{s}\|$
$\leq\kappa_{0}(1-\gamma_{0})^{t}x_{0}+\max_{0\leq s\leq t-1}\|(B_{*}\nu_{t-1-s}+\omega_{t-1-s})\|\sum_{s=0}^{\infty}\|M_{s}\|$
$\leq\kappa_{0}(1-\gamma_{0})^{t}\|x_{0}\|+\frac{\kappa_{0}}{\gamma_{0}}\max_{0\leq s\leq t-1}\|(B_{*}\nu_{t-1-s}+\omega_{t-1-s})\|$
$\leq\kappa_{0}e^{-\gamma_{0}t}\|x_{0}\|+\frac{\kappa_{0}}{\gamma_{0}}\max_{0\leq s\leq t-1}\|(B_{*}\nu_{s}+\omega_{s})\|$
Let $\zeta_{t}:=B_{*}\nu_{s}+\omega_{s}$. Then $\zeta_{t}\sim\mathcal{N}(0,(\sigma_{\omega}^{2}+2\kappa_{0}^{2}\sigma_{\omega}^{2}B_{*}B_{*}^{\top})I)$, i.e., $\zeta_{t}$ is Gaussian. Applying the Hanson-Wright inequality (Lemma 6) we obtain
$\|B_{*}\nu_{t}+\omega_{t}\|^{2}$ $\leq 5\operatorname{Tr}((\sigma_{\omega}^{2}+2\kappa_{0}^{2}\sigma_{\omega}^{2}B_{*}B_{*}^{\top})I)\log\frac{T_{0}}{\delta}$
$\leq 5\sigma_{\omega}^{2}(n+2\kappa_{0}^{2}\operatorname{Tr}(B_{*}B_{*}^{\top}))\log\frac{T_{0}}{\delta}$
$\leq 10\sigma_{\omega}^{2}(n+m\kappa_{0}^{2}\vartheta^{2})\log\frac{T_{0}}{\delta}.$
Setting $x_{0}=0$ concludes the proof of the first claim.
By definition, we also have
$z_{k}=\begin{pmatrix}x_{k}\\ K_{0}x_{k}+\nu_{k}\end{pmatrix}$,
which implies
$\|z_{k}\|$ $\leq\|x_{k}\|+\|K_{0}x_{k}\|+\|\nu_{k}\|$
$\leq 2\kappa_{0}\|x_{k}\|+\|\nu_{k}\|.$
Applying (191), with probability at least $1-\delta/2$, we have
$\|x_{t}\|\leq\sigma_{\omega}\sqrt{10\Bigl(n+m\kappa_{0}^{2}\bar{\vartheta}_{B_{*}}^{2}\Bigr)\log\frac{4t}{\delta}}.$
Moreover, by the Hanson-Wright inequality (Lemma 6),
$\max_{1\leq k\leq t}\|\nu_{k}\|^{2}\leq 10\sigma_{\omega}^{2}m\kappa_{0}^{2}\log\frac{4t}{\delta}.$
Combining these results gives
$\|z_{t}\|\leq\frac{12\sigma_{\omega}\kappa_{0}^{2}}{\gamma_{0}}\sqrt{\Bigl(n+m\kappa_{0}^{2}\bar{\vartheta}_{B_{*}}^{2}\Bigr)\log\frac{4t}{\delta}}$
that completes the proof.
L.2 Proof of Theorem 7
Proof Let $t_{w}$ denote the duration of the warm-up phase which is determined by termination condition of the warm-up phase (190):
$t_{w}=\min\Bigg\{t\;\Big|\;\frac{r_{t}^{w}}{\lambda_{\min}(\bar{V}_{t})}\leq\varepsilon\Bigg\}.$
A lower bound on the minimum eigenvalue of the covariance matrix $\bar{V}_{t}$ follows from Theorem 20 in cohen2019learning:
$\bar{V}_{t}\succeq\frac{t\sigma_{\omega}^{2}}{80}I,\quad t\geq 400(n+m+\log\tfrac{1}{\delta}).$
Therefore, the condition on $t_{w}$ becomes
$t_{w}\leq\min\Bigg\{t\;\Big|\;t\geq\frac{80r_{t}^{w}}{\sigma_{\omega}^{2}\varepsilon^{2}}\Bigg\}.$ (193)
Since $r_{t}^{w}=\mathcal{O}(\log t)$, it follows from (193) that such a $t_{w}$ exists.
Using the definition of $r_{t}^{w}$ and the bound on $\|z_{t}\|$, we obtain
$r_{t}^{w}$ $\leq 4\sigma_{\omega}^{2}n(n+m)\log\frac{n}{\delta}$
$\quad+4\sigma_{\omega}^{2}n(n+m)\log\Bigl(1+t\cdot\frac{144\sigma_{\omega}^{2}\kappa_{0}^{4}}{\gamma_{0}^{2}}(n+m\kappa_{0}^{2})\log\frac{4t}{\delta}\Bigr)$
$\quad+2\sigma_{\omega}^{2}m$
$\leq 10\sigma_{\omega}^{2}n(n+m)\log\frac{1200(n+m)\sigma_{\omega}^{2}\kappa_{0}^{6}}{\gamma_{0}^{2}}\frac{t}{\delta}.$
Then (193) can be simplified applying the inequality $x\geq a\log(bx)$ for $a\geq 3$, which holds if $x\geq 3a\log(ab)$:
$t_{w}\leq\min\Bigg\{t\geq 800n(n+m)\frac{1}{\varepsilon^{2}}\log\frac{1200(n+m)\sigma_{\omega}^{2}\kappa_{0}^{6}}{\gamma_{0}^{2}\delta}\bigg\}.$
For the ARSLO algorithm, using the definition $\varepsilon=\bar{\epsilon}(\bar{\kappa}_{1})$, we obtain
$\frac{1}{\bar{\epsilon}^{2}(\bar{\kappa}_{1})}$ $\lesssim n(n+m)\|P_{*}\|^{14}\bar{\vartheta}_{B_{*}}^{4}\vartheta^{2}\log\frac{\|P_{*}\|n(n+m)\bar{\vartheta}_{B_{*}}\vartheta}{\delta}$
$\lesssim(n+m)^{2}\|P_{*}\|^{14}\bar{\vartheta}_{B_{*}}^{6}\log\frac{\|P_{*}\|(n+m)\bar{\vartheta}_{B_{*}}}{\delta}.$
Consequently, the duration of the warm-up phase satisfies
$t_{w}\lesssim(n+m)^{4}\|P_{*}\|^{14}\bar{\vartheta}_{B_{*}}^{6}\log\frac{\|P_{*}\|(n+m)\bar{\vartheta}_{B_{*}}}{\delta}\log\frac{(n+m)\kappa_{0}}{\gamma_{0}\delta}.$
[…]
$\frac{1}{\underline{\epsilon}^{2}(\bar{\kappa}_{1},\bar{\rho})}$ $\lesssim n(n+m)^{2}\|P_{*}\|^{h}\bar{\vartheta}_{B_{*}}^{4}\vartheta^{2}\log^{2}\frac{\|P_{*}\|n(n+m)\bar{\vartheta}_{B_{*}}\vartheta}{\delta}$
$\lesssim(n+m)^{3}\|P_{*}\|^{h}\bar{\vartheta}_{B_{*}}^{6}\log^{2}\frac{\|P_{*}\|(n+m)\bar{\vartheta}_{B_{*}}}{\delta},$
[…]
$t_{w}\lesssim(n+m)^{5}\|P_{*}\|^{h}\bar{\vartheta}_{B_{*}}^{6}\log^{2}\frac{\|P_{*}\|(n+m)\bar{\vartheta}_{B_{*}}}{\delta}\log\frac{(n+m)\kappa_{0}}{\gamma_{0}\delta}.$
Additionally, from (192), we have
$\|z_{t}\|^{2}\lesssim\frac{\kappa_{0}^{6}}{\gamma_{0}^{2}}(n+m)\bar{\vartheta}_{B_{*}}^{2}\log\frac{t}{\delta}.$
The regret accumulated during the warm-up phase can therefore be written as
$R_{\text{warm-up}}(t_{w})$ $=\sum_{k=0}^{t_{w}-1}z_{k}^{\top}\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}z_{k}$
$\leq\alpha_{1}\sum_{k=0}^{t_{w}-1}\|z_{k}\|^{2}\lesssim\alpha_{1}t_{w}\|z_{t_{w}}\|^{2}.$
As a result, for the ARSLO algorithm, we obtain
$R_{\text{warm-up}}(t_{w})\lesssim\alpha_{1}(n+m)^{5}\|P_{*}\|^{14}\bar{\vartheta}_{B_{*}}^{8}\frac{\kappa_{0}^{6}}{\gamma_{0}^{2}}\log^{2}\frac{\|P_{*}\|(n+m)\bar{\vartheta}_{B_{*}}}{\delta}\log\frac{(n+m)\kappa_{0}}{\gamma_{0}\delta},$
which implies
$R_{\text{warm-up}}(t_{w})\leq\mathcal{O}\!\left(\frac{\kappa_{0}^{6}\,\alpha_{1}\,\bar{\vartheta}_{B_{*}}^{8}}{\gamma_{0}^{2}}(n+m)^{5}\|P_{*}\|^{14}\right).$
[…]
$R_{\text{warm-up}}(t_{w})\lesssim\alpha_{1}(n+m)^{6}\|P_{*}\|^{h}\bar{\vartheta}_{B_{*}}^{8}\frac{\kappa_{0}^{6}}{\gamma_{0}^{2}}\log^{3}\frac{\|P_{*}\|(n+m)\bar{\vartheta}_{B_{*}}}{\delta}\log\frac{(n+m)\kappa_{0}}{\gamma_{0}\delta},$
and consequently,
$R_{\text{warm-up}}(t_{w})\leq\mathcal{O}\!\left(\frac{\kappa_{0}^{6}\,\alpha_{1}\,\bar{\vartheta}_{B_{*}}^{8}}{\gamma_{0}^{2}}(n+m)^{6}\|P_{*}\|^{h}\right).$
This completes the proof.
Appendix M Proof of Proposition 2
Proof We analyze two cases for the perturbation noise $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, corresponding to deviations of $\|\Gamma_{t}\|$ from the nominal order $\mathcal{O}(t^{-1/2})$.
Case (i): $\|\Gamma_{t}\|=\mathcal{O}(t^{-1/2+\bar{\alpha}})$ for some $\bar{\alpha}>0$.
Applying the procedure for lower bounding the minimum eigenvalue of the covariance matrix $V_{t}$ (see Appendix F), we obtain
$\lambda_{\min}(V_{t})=\mathcal{O}(t^{1/2-\bar{\alpha}}).$
[…]
Case (ii): $\|\Gamma_{t}\|=\mathcal{O}(t^{-1/2-\bar{\alpha}})$ for some $\bar{\alpha}>0$.
In this case, the same argument (see Appendix F) yields
$\lambda_{\min}(V_{t})=\mathcal{O}(t^{1/2+\bar{\alpha}}),$
which is sufficient to satisfy the stability conditions in (30) and (31). However, under this choice of perturbation noise, an upper bound on the regret term $R_{6}(t)$ shows that its order increases from $\mathcal{O}(\sqrt{t})$ to $\mathcal{O}(t^{1/2+\bar{\alpha}})$, thereby degrading the overall regret bound.
Combining the two cases completes the proof.